### v1

# Nonreciprocal tripartite entanglement and asymmetric Einstein-Podolsky-Rosen steering via directional quantum squeezing

Abstract
The generation and manipulation of multipartite entanglement and EPR steering in macroscopic systems not only play a fundamental role in exploring the nature of quantum mechanics, but are also at the core of current developments of various nascent quantum technologies. Here we report a theoretical method using directional injection of quantum squeezing to produce nonreciprocal multipartite entanglement and EPR steering in a three-mode optomechanical system with closed-loop coupling. We show that by directionally applying a two-photon parametric driving field with a phase-matched squeezed vacuum reservoir to an optomechanical resonator, a squeezed optical mode can be introduced for one of its input directions, thereby yielding an asymmetric enhancement of optomechanical interaction and the time-reversal symmetry breaking of the system. Based on this feature, it is found that bipartite and tripartite entanglement and the associated EPR steering of the subsystems can only be generated when the coherent driving field input from the squeezing injection direction, namely, achieving nonreciprocity in such quantum correlations. More excitingly, it is also found that by properly adjusting the squeezing parameter, the overall asymmetry of EPR steering can be stepwise driven from no-way regime, one-way regime to two-way regime. These findings, holding promise for preparing rich types of entangled quantum resources with nonreciprocal correlations, may have potential applications in the area of quantum information processing such as quantum secure direct communication and one-way quantum computing.
I Introduction
Entanglement, allowing perfectly correlated positions and momenta for two spatially separated particles, has long been intriguing in quantum physics and enables numerous advanced quantum information protocols spanning from quantum networking to quantum sensing Horodecki et al. (2009). The concept of entanglement was originally addressed by Schrödinger Schrödinger (1935) in his response to the issue of the “spooky action-at-a-distance” predicted by Einstein, Podolsky, and Rosen (EPR) in their famous paradox Einstein et al. (1935), where he also coined a term that came to be known as the EPR steering Reid (1989); Ou et al. (1992). From the perspective of violations of local-hidden-state models, Wiseman et al. has formalized an operational benchmark for EPR steering Wiseman et al. (2007), by which they further proved that under the hierarchy of quantum nonlocality, EPR steering is a strict subset of entanglement and a strict superset of Bell nonlocality. An appealing feature of EPR steering is that it describes how local measurements on one part of the system can steer (alter) the state of the other part at a different location. This defining characteristics, not held by the other two types of quantum nonlocality, reveals the intrinsic asymmetry of EPR steering and offers an insight into directional nonlocality He et al. (2015a), which plays an indispensable role in enabling quantum techniques using untrusted devices Xiang et al. (2022), such as secure quantum key distribution Branciard et al. (2012); Walk et al. (2016), randomness certification Li et al. (2024), and no-cloning quantum teleportation He et al. (2015b); Chiu et al. (2016). In the past few decades, after a series of rigorous mathematical characterizations of quantum nonlocality Simon (2000); Adesso et al. (2004); Adesso and Illuminati (2007); Kogias et al. (2015), a great deal of progresses have been made experimentally to prepare entangled or steerable states of microscopic and macroscopic particles, involving platforms based on photons Wang et al. (2016a), ions Stute et al. (2012), atoms van Leent et al. (2022), superconducting circuits Kurpiers et al. (2018), and cavity optomechanical (COM) devices Palomaki et al. (2013); Kotler et al. (2021); Mercier de Lépinay et al. (2021). However, in terms of generation and manipulation of macroscopic entanglement, it is still challenging to avoid the decoherence effect induced by device imperfection. Very recently, to overcome this obstacle and achieve entangled state with high fidelity, a large number of theoretical proposals have been raised, which relies on synthetic gauge field Lai et al. (2022a); Liu et al. (2023), reservoir engineering Wang and Clerk (2013); Yang et al. (2015); Jiao et al. , dark-mode or feedback control Lai et al. (2022b); Huang et al. (2022a); Li et al. (2017), photon counting Ho et al. (2018), dynamical modulation Wang et al. (2016b); Yang et al. (2024), and optical nonreciprocity Jiao et al. (2020, 2022); Huang et al. (2024).
In parallel, nonreciprocal optical devices, such as optical diodes and circulators, have attracted intense interests due to their potential applications in the next-generation photonic information processing Sounas and Alù (2017). The conventional approach for achieving nonreciprocal transmission of light mainly relies on magneto-optical effect Shoji and Mizumoto (2014), which requires magnetic-based materials that are bulk and quite lossy at optical frequency. In recent years, by breaking the time-reversal symmetry Caloz et al. (2018) via spatial-temporal modulation, optical nonlinearity or Sagnac effect, several magnetic-free scheme have been theoretically proposed and experimentally demonstrated based on atomic ensembles Zhang et al. (2018); Yang et al. (2019); Liang et al. (2020), optical Dong et al. (2015); Xia et al. (2018); Rosario Hamann et al. (2018) or COM Manipatruni et al. (2009); Shen et al. (2016); Bernier et al. (2017) devices, synthetic structures Peng et al. (2014); Chang et al. (2014), moving medium Wang et al. (2013); Lü et al. (2017); Maayani et al. (2018), and squeezing injectionTang et al. (2022). With these advances, the peculiar feature of nonreciprocal optical devices, allowing the flow of light from one side but block it from the other, has recently been employed to realize the unidirectional manipulation of light-matter interaction, which enables a variety of classical nonreciprocal phenomenon, such as one-way optical chaos Zhang et al. (2021) or solitons Li et al. (2021), unidirectional phonon Jiang et al. (2018) or magnon Huang et al. (2022b) lasing, and nonreciprocal enhancement optomechanical sidebands Li et al. (2020), to name a few. Very recently, this functionality has also been extended into the quantum regime to achieve nonreciprocal control of diverse nonclassical effect, such as photon Huang et al. (2018); Yang et al. (2023); Wang et al. (2023) or magnon Wang et al. (2022) blockade, microscopic Orr et al. (2023); Shen et al. (2023); Bin et al. (2024) or macroscopic Jiao et al. (2020, 2022); Huang et al. (2024) entanglement, quantum phase transition Fruchart et al. (2021); Zhu et al. (2024), and EPR steering Guan et al. (2024); Wu et al. (2024); Zhong et al. (2023).
Inspired by these studies, we investigate how to achieve nonreciprocal entanglement and asymmetric EPR steering in a hybrid COM system based on whispering-gallery-mode (WGM) resonator. We show that by applying a two-photon parametric driving field from one input direction while not applying it from the other, the two degenerate counterpropagating optical modes of the WGM resonator can be unidirectionally squeezed, leading to the time-reversal symmetry breaking of the system. Based on this chiral squeezing characteristics, we show that various types of bipartite entanglement of the subsystems and the tripartite entanglement of the whole system can be generated only when coherent light input from the squeezing injection direction, which indicates the achievement of nonreciprocal entanglement. More interestingly, by properly adjusting the squeezing parameters, we further show that the directionality of EPR steering can also be stepwise driven from no-way regime, one-way regime to two-way regime in an asymmetric way. These results, opening up a promising way to prepare entangled quantum resources with nonreciprocal and asymmetric features, are useful for a variety of secure quantum information protocols Branciard et al. (2012); Walk et al. (2016); Li et al. (2024); He et al. (2015b); Chiu et al. (2016).
This paper is structured as follows. In Sec. II, we introduce the theoretical model of the proposed COM system and obtain the effective Hamiltonian and master equation of this system, by which we calculate the system dynamics and evaluate the quantitative measures for entanglement and EPR steering. In Sec. III, based on the numerical simulations, we analyze the method and mechanism for achieving nonreciprocal generation and manipulation of various types of entanglement and EPR steering. In Sec. IV, a brief summary is given. The detailed derivation process of the effective Hamiltonian, the associated master equation, and the system dynamics is given in the Appendix.
Figure 1: Schematic of the proposed COM setup consisting of a bottle WGM resonator, an optical waveguide, and a tapered glass-fiber nanospike. The WGM resonator supports two degenerate counterpropagating optical modes, i.e., CW and CCW modes. The optical driving field is input from the waveguide, which can be coupled into and out of the WGM resonator via evanescent coupling. When placing a nanospike close to the WGM resonator, it excites two mechanical modes due to the COM interaction induced by the WGM evanescent field, whose vibrating directions are parallel and orthogonal to the resonator surface, respectively. To achieve nonreciprocal manipulation of quantum entanglement and EPR steering, we consider two different scenarios for the the CW and CCW input cases. For the CW input case, only a coherent driving field is applied, while for the CCW input case, an additional two-photon parametric driving field with a phase-matched squeezed vacuum reservoir is applied.
II Theoretical model and system dynamics
The proposed setup consists of a bottle WGM resonator, an optical waveguide, and a tapered glass-fiber nanospike, which is depicted in Fig. 1. Due to the spatial symmetry of the bottle WGM resonator, it can support two degenerate counterpropagating optical modes with resonance frequency $\omega_{c}$, i.e., the clockwise (CW) and counterclockwise (CCW) modes, respectively. The nanospike is usually fabricated by scanning an oxybutane flame along the length of a glass fiber while gently pulling it. As demonstrated in the recent experiments Xie et al. (2016, 2019); Pennetta et al. (2020), by mounting such nanospike on a stainless-steel holder and placing it close to the bottle WGM resonator, it offers two mechanical modes due to the WGM evanescent field induced COM interaction. These two mechanical modes with frequency $\omega_{m,j}$ $(j=1,2)$ are vibrating in the directions that are orthogonal and parallel to the bottle surface, and they are coupled linearly due to the optically mediated hybridization effect Shkarin et al. (2014); Lin et al. (2010). The WGM resonator can be driven from both sides, where the optical driving fields from the waveguide are coupled into and out of the WGM resonator via evanescent coupling.
In addition, we consider two different scenarios for the driving fields applied from the CW and CCW input directions, respectively. For the CCW input case, a coherent driving field with frequency $\omega_{d}/2$ and a two-photon parametric driving field with frequency $\omega_{d}$ are simultaneously applied to the waveguide. In this situation, in a rotating reference frame with respect to the driving frequency $\omega_{d}/2$, the Hamiltonian of the whole system can be expressed as (setting $\hbar=1$)
$\hat{H}_{\circlearrowleft}=$ $\hat{H}_{\textrm{sys}}+\hat{H}_{\textrm{dr},\circlearrowleft},$
$\hat{H}_{\textrm{sys}}=$ $\Delta_{c}\hat{a}_{\circlearrowleft}^{\dagger}\hat{a}_{\circlearrowleft}+\sum_{j=1,2}\left[\dfrac{\omega_{m,j}}{2}(\hat{p}_{j}^{2}+\hat{q}_{j}^{2})-g_{j}\hat{a}_{\circlearrowleft}^{\dagger}\hat{a}_{\circlearrowleft}\hat{q}_{j}\right]$
$+\chi\hat{q}_{1}\hat{q}_{2},$
$\hat{H}_{\textrm{dr},\circlearrowleft}=$ $\Xi_{d}(e^{-i\theta_{d}}\hat{a}_{\circlearrowleft}^{\dagger 2}+e^{i\theta_{d}}\hat{a}_{\circlearrowleft}^{2})+i\varepsilon_{d,\circlearrowleft}(\hat{a}_{\circlearrowleft}^{\dagger}-\hat{a}_{\circlearrowleft}),$ (1)
where $\hat{a}_{\circlearrowleft}$ ($\hat{a}_{\circlearrowleft}^{\dagger}$) is the annihilation (creation) operator of the CCW mode, while $\hat{q}_{j}$ and $\hat{p}_{j}$ are the dimensionless displacement and momentum operators of the $j$th mechanical mode, respectively. $\Delta_{c}=\omega_{c}-\omega_{d}/2$ denotes the optical detuning, and $\omega_{c}$ is the resonance frequency of the CCW (CW) mode. $g_{j}=(\omega_{c}/R)\sqrt{\hbar/m_{j}\omega_{m,j}}$ is the single-photon COM coupling strength, with $R$ the bottle radius and $m_{j}$ the effective mass of the $j$th mechanical mode. $\chi$ is the mechanical coupling strength. $\Xi_{d}$ and $\theta_{d}$ are the amplitude and phase of the two-photon parametric driving field. The amplitude of the coherent driving field is given by $|\varepsilon_{d,\circlearrowleft}|=\sqrt{2\kappa P_{d}/\hbar\omega_{d}}$, where $P_{d}$ is the input laser power and $\kappa$ is the optical decay rate.
After performing the Bogoliubov transformation with a unitary operator, $S(\eta_{d})=\exp[(-\eta_{d}\hat{a}_{\circlearrowleft}^{\dagger 2}+\eta_{d}^{\ast}\hat{a}_{\circlearrowleft}^{2})/2]$, a squeezed optical mode $\hat{a}_{s,\circlearrowleft}$ can be introduced, i.e.,
$S^{\dagger}(\eta_{d})\hat{a}_{\circlearrowleft}S(\eta_{d})=\cosh(r_{d})\hat{a}_{s,\circlearrowleft}-e^{-i\theta_{d}}\sinh(r_{d})\hat{a}_{s,\circlearrowleft}^{\dagger},$ (2)
where $\eta_{d}=r_{d}e^{-i\theta_{d}}$ is the complex squeezing parameter, with a squeezing strength $r_{d}$ and a squeezing reference angle $\theta_{d}$. Hence, by dropping the constant terms, the effective Hamiltonian of the total system in the squeezing picture is derived as [see Appendix A for more details]
$\hat{\tilde{H}}_{\circlearrowleft}=$ $\omega_{s}\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{a}_{s,\circlearrowleft}+\sum_{j=1,2}\left[\dfrac{\omega_{m,j}}{2}(\hat{p}_{j}^{2}+\hat{q}_{j}^{2})-\zeta_{s,j}\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{a}_{s,\circlearrowleft}\hat{q}_{j}\right.$
$\left.+\dfrac{\zeta_{p,j}}{2}(e^{-i\theta_{d}}\hat{a}_{s,\circlearrowleft}^{\dagger 2}+e^{i\theta_{d}}\hat{a}_{s,\circlearrowleft}^{2})\hat{q}_{j}-F_{j}\hat{q}_{j}\right]+\chi\hat{q}_{1}\hat{q}_{2}$
$+i\varepsilon_{d,\circlearrowleft}\sinh(r_{d})(e^{-i\theta_{d}}\hat{a}_{s,\circlearrowleft}^{\dagger}-e^{i\theta_{d}}\hat{a}_{s,\circlearrowleft})$
$+i\varepsilon_{d,\circlearrowleft}\cosh(r_{d})(\hat{a}_{s,\circlearrowleft}^{\dagger}-\hat{a}_{s,\circlearrowleft}),$ (3)
where $\omega_{s}=(\Delta_{c}-2\Xi_{d})\exp(2r_{d})$ is the effective resonance frequency of the squeezed optical mode. $\zeta_{s,j}=g_{j}\cosh(2r_{d})$ and $\zeta_{p,j}=g_{j}\sinh(2r_{d})$ are the effective COM coupling strength for the $j$th mechanical mode induced by the radiation pressure and the parametric amplification, respectively. $F_{j}=g_{j}\sinh^{2}(r_{d})$ denotes the strength of the constant mechanical driving force exerted on the $j$th mechanical mode, which is caused by the parametric amplification.
As discussed in the preceding studies Jiao et al. ; Tang et al. (2022); Guan et al. (2024); Wu et al. (2024); Lü et al. (2015), by adjusting the squeezing parameters with respect to the two-photon parametric driving field, the effective light-matter interaction can be manipulated and enhanced in a controllable way. However, it is worth emphasizing that when the squeezed optical mode is coupled to a normal vacuum reservoir, the optical dissipation will be inevitably amplified in the meantime, thus leading to its quantum decoherence. To eliminate such optical dissipation induced by the intracavity squeezing, one can apply a broadband squeezed-vacuum reservoir that is phase-matched with the squeezed optical mode, i.e., the squeezing strength $r_{e}$ and phase $\theta_{e}$ of this reservoir satisfying the relationships $r_{e}-r_{d}=0$ and $\theta_{e}-\theta_{d}=\pm n\pi~{}(n=1,3,5,\ldots)$. Note that the injection of this required broadband squeezed-vacuum reservoir has been reported through using an optical parametric amplifier Kashiwazaki et al. (2020). Additionally, given that the two mechanical modes are coupled with two independent thermal reservoirs at same bath temperature $T$, the Born-Markovian master equation describing the dynamics of the whole system in the squeezing picture can be derived as [see Appendix B for more details]
$\dot{\tilde{\rho}}_{\circlearrowleft}\!=$ $i\!\left[\tilde{\rho}_{\circlearrowleft},\hat{\tilde{H}}_{\circlearrowleft}\right]\!+\!\frac{\kappa}{2}\mathcal{D}[\hat{a}_{s,\circlearrowleft}]\tilde{\rho}_{\circlearrowleft}\!-\!\sum_{j=1,2}\left(i\dfrac{\gamma_{m,j}}{2}[\hat{q}_{j},\{\hat{p}_{j},\tilde{\rho}_{\circlearrowleft}\}]\right.$
$\left.+\gamma_{m,j}\bar{n}_{m,j}[\hat{q}_{j},[\hat{q}_{j},\tilde{\rho}_{\circlearrowleft}]]\right),$ (4)
where
$\mathcal{D}[\hat{a}_{s,\circlearrowleft}]\tilde{\rho}_{\circlearrowleft}=2\hat{a}_{s,\circlearrowleft}\tilde{\rho}_{\circlearrowleft}\hat{a}_{s,\circlearrowleft}^{\dagger}-(\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{a}_{s,\circlearrowleft}\tilde{\rho}_{\circlearrowleft}+\tilde{\rho}_{\circlearrowleft}\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{a}_{s,\circlearrowleft})$ (5)
[…]
For the CW input case, there is merely a coherent driving field with frequency $\omega_{d}/2$ applied to the waveguide. In this situation, the Hamiltonian of the whole system reads
$\hat{H}_{\circlearrowright}=$ $\hat{H}_{\textrm{sys}}+\hat{H}_{\textrm{dr},\circlearrowright},$
$\hat{H}_{\textrm{sys}}=$ $\Delta_{c}\hat{a}_{\circlearrowright}^{\dagger}\hat{a}_{\circlearrowright}+\sum_{j=1,2}\left[\dfrac{\omega_{m,j}}{2}(\hat{p}_{j}^{2}+\hat{q}_{j}^{2})-g_{j}\hat{a}_{\circlearrowright}^{\dagger}\hat{a}_{\circlearrowright}\hat{q}_{j}\right]$
$+\chi\hat{q}_{1}\hat{q}_{2},$
$\hat{H}_{\textrm{dr},\circlearrowright}=$ $i\varepsilon_{d,\circlearrowright}(\hat{a}_{\circlearrowright}^{\dagger}-\hat{a}_{\circlearrowright}),$ (6)
and the associated Born-Markovian master equation is given by Gardiner and Zoller (2000)
$\dot{\rho}_{\circlearrowright}\!=$ $i\!\left[\rho_{\circlearrowright},\hat{H}_{\circlearrowright}\right]\!+\!\frac{\kappa}{2}\mathcal{D}[\hat{a}_{\circlearrowright}]\rho_{\circlearrowright}\!-\!\sum_{j=1,2}\left(i\dfrac{\gamma_{m,j}}{2}[\hat{q}_{j},\{\hat{p}_{j},\rho_{\circlearrowright}\}]\right.$
$\left.+\gamma_{m,j}\bar{n}_{m,j}[\hat{q}_{j},[\hat{q}_{j},\rho_{\circlearrowright}]]\right),$ (7)
where $\hat{a}_{\circlearrowright}$ ($\hat{a}_{\circlearrowright}^{\dagger}$) is the annihilation (creation) operator of the CW mode, and $\rho_{\circlearrowright}$ is the density operator with respect to the CW input case. Note that in the absence of the two-photon parametric driving field, the Bogoliubov transformation has no influence on the system dynamics. Therefore, we apply a normal vacuum reservoir to the CW mode and analyze the dynamical evolution of this system in the original picture.
[…]
For the CCW input case, due to the nonlinear COM interactions in Hamiltonian (3), the system dynamics is difficult to be directly solved. In order to deal with this problem, one can linearize the Hamiltonian by expanding each operator as a sum of its steady-state mean value and a small quantum fluctuation around it under the condition of strong coherent optical driving, i.e., $\hat{a}_{s,\circlearrowleft}=\bar{a}_{s,\circlearrowleft}+\delta\hat{a}_{s,\circlearrowleft}$, $\hat{q}_{j}=\bar{q}_{j}+\delta\hat{q}_{j}$, $\hat{p}_{j}=\bar{p}_{j}+\delta\hat{p}_{j}$. The steady-state mean values of the optical and mechanical modes can be derived by using the master equation (4), i.e.,
$\bar{a}_{s,\circlearrowright}=$ $-\dfrac{(i\Delta_{s}-\dfrac{\kappa}{2})A_{1}+iA_{2}\beta_{p}}{(\Delta_{s}^{2}+\dfrac{\kappa^{2}}{4})-\beta_{p}^{2}}\varepsilon_{d,\circlearrowleft},$
$\bar{q}_{1}=$ $\dfrac{\omega_{m,2}B_{1}-\chi B_{2}}{\omega_{m,1}\omega_{m,2}-\chi^{2}},~{}~{}\bar{p}_{1}=0,$
$\bar{q}_{2}=$ $\dfrac{\omega_{m,1}B_{2}-\chi B_{1}}{\omega_{m,1}\omega_{m,2}-\chi^{2}},~{}~{}\bar{p}_{2}=0,$ (8)
with
$\alpha_{s}$ $=e^{-i\theta_{d}}\bar{a}_{s,\circlearrowleft}^{\ast 2}+e^{i\theta_{d}}\bar{a}_{s,\circlearrowleft}^{2},$
$\beta_{s}$ $=\zeta_{s,1}\bar{q}_{1}+\zeta_{s,2}\bar{q}_{2},$
$\beta_{p}$ $=\zeta_{p,1}\bar{q}_{1}+\zeta_{p,2}\bar{q}_{2},$
$\Delta_{s}$ $=\omega_{s}-\beta_{s},$
$A_{1}$ $=\cosh(r_{d})+\sinh(r_{d})e^{-i\theta_{d}},$
$A_{2}$ $=\cosh(r_{d})e^{-i\theta_{d}}+\sinh(r_{d}),$
$B_{1}$ $=\zeta_{s,1}|\bar{a}_{s,\circlearrowleft}|^{2}-\dfrac{\zeta_{p,1}}{2}\alpha_{s}+F_{1},$
$B_{2}$ $=\zeta_{s,2}|\bar{a}_{s,\circlearrowleft}|^{2}-\dfrac{\zeta_{p,2}}{2}\alpha_{s}+F_{2}.$ (9)
Equation (8) indicates that the field amplitude $\bar{a}_{s,\circlearrowleft}$ is not only dependent on the strength of the coherent driving field, but also relies on the squeezing strength and reference angle of the squeezed optical mode. This allows us to regulate the COM coupling strength and break the optical reciprocity of the system by adjusting the squeezing parameters. Moreover, it is also seen that the mean values of the mechanical displacement $\bar{q}_{j}$ are coupled to each other by a factor of $\chi$, which is due to the phonon hopping process between the two mechanical modes.
Then, by substituting the expansions of the quantum operators into the Hamiltonian (3), one can directly derive the linearized Hamiltonian as
$\hat{\tilde{H}}_{lin,\circlearrowleft}=$ $\Delta_{s}\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{a}_{s,\circlearrowleft}+\sum_{j=1,2}\left[\dfrac{\omega_{m,j}}{2}(\hat{p}_{j}^{2}+\hat{q}_{j}^{2})\right.$
$\left.-(\Lambda_{j}\hat{a}_{s,\circlearrowleft}^{\dagger}+\Lambda_{j}^{\ast}\hat{a}_{s,\circlearrowleft})\hat{q}_{j}\right]+\chi\hat{q}_{1}\hat{q}_{2},$ (10)
and the associated master equation as
$\dot{\tilde{\rho}}_{\circlearrowleft}\!=$ $i\!\left[\tilde{\rho}_{\circlearrowleft},\tilde{H}_{lin,\circlearrowleft}\right]\!+\!\frac{\kappa}{2}\mathcal{D}[\hat{a}_{s,\circlearrowleft}]\tilde{\rho}_{\circlearrowleft}\!-\!\sum_{j=1,2}\!\left(i\dfrac{\gamma_{m,j}}{2}[\hat{q}_{j},\{\hat{p}_{j},\tilde{\rho}_{\circlearrowleft}\}]\right.$
$\left.+\gamma_{m,j}\bar{n}_{m,j}[\hat{q}_{j},[\hat{q}_{j},\tilde{\rho}_{\circlearrowleft}]]\right),$ (11)
where
$\Lambda_{j}=G_{j}\cosh(2r_{d})-G_{j}^{\ast}\sinh(2r_{d})e^{-i\theta_{d}}$ (12)
is the effective COM coupling rate, with $G_{j}=g_{j}\bar{a}_{s,\circlearrowleft}$. For notational convenience, we have neglected the symbol ``$\delta"$ in the expression of quantum fluctuation operators in Eqs. (10) and (11). We also emphasize that in the weak COM coupling regime, the steady-state mean value of the optical mode is much larger than those of the mechanical modes, i.e., $|\bar{a}_{s,\circlearrowleft}|\gg|\bar{q}_{j}|$. Under this condition, we have ignored the terms $\hat{a}_{s,\circlearrowleft}^{2}$ and $\hat{a}_{s,\circlearrowleft}^{\dagger 2}$ in Hamiltonian (10), whose coefficient $\beta_{p}$ is dominated by $\bar{q}_{j}$. In the following discussions, for ensuring the validity of Hamiltonian (10), the system parameters have been strictly restricted to satisfy the condition of weak COM coupling.
Figure 2: Squeezing-induced nonreciprocal bipartite and tripartite entanglement. The logarithmic negativity $E_{\mathcal{N}}^{\textit{a}|\textit{m}_{\textit{1}}}$ (a), $E_{\mathcal{N}}^{\textit{a}|\textit{m}_{\textit{2}}}$ (b), $E_{\mathcal{N}}^{\textit{m}_{\textit{1}}|\textit{m}_{\textit{2}}}$ (c), and the minimum residual contangle $\mathcal{R}^{\textrm{min}}_{\tau}$ (d) for opposite optical input directions are plotted as functions of the scaled optical detuning $\Delta_{c}/\omega_{m}$. The left panels of these figures show the bipartite and tripartite entanglement with respect to two specific squeezing reference angles of $\theta_{d}=0.75\pi$ and $\pi$, respectively, with $r_{d}=0.3$. The right panels show in detail the dependence of such entanglement on squeezing reference angle $\theta_{d}$.
Since the system dynamics is linearized now and the input noises for the optical and mechanical modes are Gaussian, the steady state of the system, independently of any initial state, could eventually evolve into a tripartite zero-mean Gaussian state, whose statistic is fully characterized by a $6\times 6$ covariance matrix (CM) $V$ with its matrix element
$V_{kl}=\langle\psi_{k}\psi_{l}+\psi_{l}\psi_{k}\rangle/2,~{}(k,l=1,2,\ldots,6).$ (13)
Here $\psi=(\hat{X},\hat{Y},\hat{q}_{1},\hat{p}_{1},\hat{q}_{2},\hat{p}_{2})^{T}$ is the vector of optical and mechanical quadrature operators, with its components defined by
$\hat{X}=\dfrac{1}{\sqrt{2}}\left(\hat{a}_{s,\circlearrowleft}^{\dagger}+\hat{a}_{s,\circlearrowleft}\right),~{}~{}~{}\hat{Y}=\dfrac{i}{\sqrt{2}}\left(\hat{a}_{s,\circlearrowleft}^{\dagger}-\hat{a}_{s,\circlearrowleft}\right).$ (14)
Employing the master equation (11), one can obtain the dynamics of arbitrary quantum correlation between the optical and mechanical quadrature operators in the CM $V$. In terms of optical bosonic operator, we define the following quantum correlation functions
$x_{1}=\langle\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{a}_{s,\circlearrowleft}\rangle,~{}x_{2}=\langle\hat{a}_{s,\circlearrowleft}\hat{a}_{s,\circlearrowleft}^{\dagger}\rangle,~{}x_{3}=\langle\hat{q}_{1}\hat{q}_{1}\rangle,$
$x_{4}=\langle\hat{p}_{1}\hat{p}_{1}\rangle,~{}x_{5}=\langle\hat{q}_{2}\hat{q}_{2}\rangle,~{}x_{6}=\langle\hat{p}_{2}\hat{p}_{2}\rangle,$
$x_{7}=\langle\hat{a}_{s,\circlearrowleft}\hat{a}_{s,\circlearrowleft}\rangle,~{}x_{8}=\langle\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{a}_{s,\circlearrowleft}^{\dagger}\rangle,~{}x_{9}=\langle\hat{q}_{1}\hat{p}_{1}\rangle,$
$x_{10}=\langle\hat{p}_{1}\hat{q}_{1}\rangle,~{}x_{11}=\langle\hat{q}_{2}\hat{p}_{2}\rangle,~{}x_{12}=\langle\hat{p}_{2}\hat{q}_{2}\rangle,$
$x_{13}=\langle\hat{a}_{s,\circlearrowleft}\hat{q}_{1}\rangle,~{}x_{14}=\langle\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{q}_{1}\rangle,~{}x_{15}=\langle\hat{a}_{s,\circlearrowleft}\hat{p}_{1}\rangle,$
$x_{16}=\langle\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{p}_{1}\rangle,~{}x_{17}=\langle\hat{a}_{s,\circlearrowleft}\hat{q}_{2}\rangle,~{}x_{18}=\langle\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{q}_{2}\rangle,$
$x_{19}=\langle\hat{a}_{s,\circlearrowleft}\hat{p}_{2}\rangle,~{}x_{20}=\langle\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{p}_{2}\rangle,~{}x_{21}=\langle\hat{q}_{1}\hat{q}_{2}\rangle,$
$x_{22}=\langle\hat{p}_{1}\hat{p}_{2}\rangle,~{}x_{23}=\langle\hat{q}_{1}\hat{p}_{2}\rangle,~{}x_{24}=\langle\hat{q}_{2}\hat{p}_{1}\rangle.$ (15)
By grouping together them in a vector, $X=(x_{1},x_{2},\ldots,x_{24})^{T}$, one can obtain its time evolution equation as
$\dfrac{d}{dt}X=M\cdot X+N,$ (16)
where
$N=$ $\left[0,\kappa,0,2\gamma_{m,1}\bar{n}_{m,1},0,2\gamma_{m,2}\bar{n}_{m,2},0,0,\right.$
$\left.0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0\right]^{T},$ (17)
is a vector involving the correlation functions for the input quantum noises. Here the exact expression for the coefficient matrix $M$ and the detailed derivation process of the time evolution equation (16) are too cumbersome, and, for convenience, we have reported them in Appendix C. Notably, by numerically solving Eq. (16) and using the relations between bosonic operators and quadrature operators, one can directly obtain the steady-state CM $V$.
Regarding the verification of bipartite and tripartite entanglement, we adopt the logarithmic negativity $E_{\mathcal{N}}$ and the minimum residual contangle $\mathcal{R}^{\textrm{min}}_{\tau}$ as the quantitative entanglement measures, respectively, which are defined based on specifying the positivity of the partial transpose of the CM $V$. For continuous-variable (CV) bipartite Gaussian state, the logarithmic negativity $E_{\mathcal{N}}^{\mu|\nu}$ is defined as Adesso et al. (2004)
$E_{\mathcal{N}}^{\mu|\nu}=$ $\max\,\big{[}0,-\ln(2\eta_{0}^{-})\big{]},$ (18)
where $\eta_{0}^{-}\!\equiv\!2^{-1/2}\{\Sigma(V_{\mu\nu})-[\Sigma(V_{\mu\nu})^{2}-4\det\!V_{\mu\nu}]^{1/2}\}^{1/2}$ , with $\Sigma(V_{\mu\nu})\!\equiv\!\det\mathcal{A}_{\mu}+\det\mathcal{B}_{\nu}-2\det\mathcal{C}_{\mu\nu}$, is the minimum symplectic eigenvalue of the partial transpose of a reduced $4\times 4$ CM $V_{\mu\nu}$. The reduced CM $V_{\mu\nu}$ contains the entries of $V$ associated with the selected bipartition $\mu$ and $\nu$, and it can be obtained by removing the rows and columns of the unwanted modes in $V$, whose $2\times 2$ block form is given by
$V_{\mu\nu}=\left(\begin{matrix}\mathcal{A}_{\mu}&\mathcal{C}_{\mu\nu}\\ \mathcal{C}_{\mu\nu}^{\textit{T}}&\mathcal{B}_{\nu}\end{matrix}\right).$ (19)
Equation (18) quantifies how much the positivity of the partial transpose condition for separability is violated for the considered Gaussian states, and it is equivalent to Simon’s necessary and sufficient entanglement nonpositive partial transpose criterion (or the related Peres-Horodecki criterion) Simon (2000). The selected bipartition $\mu$ and $\nu$ gets entangled if and only if $\eta_{0}^{-}<1/2$, where $E_{\mathcal{N}}$ has a nonzero value. To further confirm the presence and directionality of Gaussian steering between the two entangled bipartition $\mu$ and $\nu$, we introduce an intuitive and computable quantification of quantum steerability Kogias et al. (2015), which is defined as
$\mathcal{G}^{\mu\rightarrow\nu}=$ $\max\left[0,\dfrac{1}{2}\ln\dfrac{\det\mathcal{A}_{\mu}}{4\det V_{\mu\nu}}\right],$
$\mathcal{G}^{\nu\rightarrow\mu}=$ $\max\left[0,\dfrac{1}{2}\ln\dfrac{\det\mathcal{B}_{\nu}}{4\det V_{\mu\nu}}\right].$ (20)
$\mathcal{G}^{\mu\rightarrow\nu}>0$ ($\mathcal{G}^{\nu\rightarrow\mu}>0$) implies that the bipartite Gaussian state characterized by CM $V_{\mu\nu}$ is steerable from mode $\mu$ ($\nu$) to mode $\nu$ ($\mu$) through Gaussian measurements on mode $\mu$ ($\nu$), where a higher value of $\mathcal{G}$ represents the stronger Gaussian steerability.
The minimum residual contangle, $\mathcal{R}^{\textrm{min}}_{\tau}$, which provides a bona fide quantification of CV tripartite entanglement, is defined as Adesso and Illuminati (2007)
$\mathcal{R}^{\textrm{min}}_{\tau}=$ $\min_{(r,s,t)}\,\big{[}E^{r|st}_{\tau}-E^{r|s}_{\tau}-E^{r|t}_{\tau}\big{]},$ (21)
where $(r,s,t)$ denotes all the possible permutations of the three-mode indexes. $E_{\tau}^{\mu|\nu}$ is the contangle of subsystems of $\mu$ ($\mu$ contains one mode) and $\nu$ ($\nu$ contains one or two modes), which can be defined by a proper entanglement monotone, e.g., the squared logarithmic negativity. Based on Eq. (18), the one-mode-vs-one-mode contangle $E^{r|s}_{\tau}$ and $E^{r|t}_{\tau}$ can be directly obtained by employing its definition, namely, $E_{\tau}^{\mu|\nu}\equiv[E_{\mathcal{N}}^{\mu|\nu}]^{2}$. However, when calculating the one-mode-vs-two-modes contangle $E^{r|st}_{\tau}$, one must alter the basic definition of Eq. (18) by rewriting the definition of $\eta_{0}^{-}$ as given by $\eta_{0}^{-}\equiv\min\,[\textrm{eig}|i\Omega_{3}\tilde{V}_{r|st}|]$, where $\eta_{0}^{-}$ becomes the minimum symplectic eigenvalue of the partial transpose of a $6\times 6$ CM $V$, with $\Omega_{3}=\oplus_{k=1}^{3}i\sigma_{y}$ and $\sigma_{y}$ the y-Pauli matrix. $\tilde{V}_{r|st}$ corresponds to the partial transpose of $V$, which connects to $V$ with the relation of $\tilde{V}_{r|st}=P_{r|st}VP_{r|st}$, and $P_{r|st}=\textrm{diag}(1,-1,1,1,1,1)$ is the partial transposition matrix. In terms of $E^{s|rt}_{\tau}$ and $E^{t|sr}_{\tau}$, the corresponding partial transpose matrices are given by $P_{s|rt}=\textrm{diag}(1,1,1,-1,1,1)$ and $P_{t|sr}=\textrm{diag}(1,1,1,1,1,-1)$, respectively. In addition, according to the Coffman-Kundu-Wootters monogamy inequality for quantum entanglement, we also note that the residual contangle is required to satisfy the following monogamy condition, i.e., $E^{r|st}_{\tau}-E^{r|s}_{\tau}-E^{r|t}_{\tau}\geq 0$, which means that the bipartite entanglement between the partition $r$ and the remaining two partitions $st$ is never smaller than the sum of the $r|s$ and $r|t$ bipartite entanglements in the reduced states. As such, if there are nonzero values of the minimum residual contangle, i.e., $\mathcal{R}_{\tau}^{\textrm{min}}>0$, one can verify that the full tripartite entanglement is present for the CV system.
III Nonreciprocal manipulation of quantum entanglement and EPR steering
Now we start to investigate the behaviors of quantum entanglement and EPR steering in regards to the coherent driving fields input from the CW and CCW direction in our proposed scheme. For this purpose, we evaluate the entanglement measures by numerically solving the dynamics of the steady-state CM $V$ (16) within the following parameters: $\kappa/2\pi=\textrm{1750\,Hz}$, $\omega_{m}/2\pi=\textrm{1945\,Hz}$, $\gamma_{m}/2\pi=\textrm{195\,Hz}$, $g_{1}/2\pi=\textrm{214\,Hz}$, $g_{2}/2\pi=\textrm{321\,Hz}$, $\chi/2\pi=\textrm{292\,Hz}$, $\bar{n}_{m}=\textrm{0.9}$, $\theta_{e}=\textrm{0}$, which is partially chosen from a recent experiment Pennetta et al. (2020). Note that, for simplicity and without loss of generality, we have supposed that the two mechanical modes are identical, i.e, $\omega_{m,1}=\omega_{m,2}=\omega_{m}$, $\gamma_{m,1}=\gamma_{m,2}=\gamma_{m}$, $\bar{n}_{m,1}=\bar{n}_{m,2}=\bar{n}_{m}$, while their COM coupling strength are asymmetric, i.e., $g_{2}/g_{1}=\textrm{1.5}$. This condition is convenient for producing quantum entanglement and EPR steering.
Figure 3: Squeezing-induced nonreciprocal EPR steering. The quantum steerability $\mathcal{G}^{\textit{a}\rightarrow\textit{m}_{\textit{1}}}$ (a), $\mathcal{G}^{\textit{m}_{\textit{1}}\rightarrow\textit{a}}$ (b), $\mathcal{G}^{\textit{m}_{\textit{1}}\rightarrow\textit{m}_{\textit{2}}}$ (c), and $\mathcal{G}^{\textit{m}_{\textit{2}}\rightarrow\textit{m}_{\textit{1}}}$ (d) for opposite optical input directions are plotted as as functions of the scaled optical detuning $\Delta_{c}/\omega_{m}$ under different squeezing parameter $r_{d}$. With the increase of $r_{d}$, EPR steering can be stepwise driven from no-way regime, one-way regime to two-way regime for the CCW input case but not for the CW input case, indicating the achievement of nonreciprocal EPR steering.
We first present how to achieve nonreciprocal bipartite and tripartite entanglement by leveraging a directional two-photon parametric driving field. Figure 2 shows the logarithmic negativity $E_{\mathcal{N}}^{\textit{a}|\textit{m}_{\textit{1}}}$, $E_{\mathcal{N}}^{\textit{a}|\textit{m}_{\textit{2}}}$, $E_{\mathcal{N}}^{\textit{m}_{\textit{1}}|\textit{m}_{\textit{2}}}$ and the minimum residual contangle $\mathcal{R}^{\textrm{min}}_{\tau}$ for opposite optical input directions as functions of the scaled optical detuning $\Delta_{c}/\omega_{m}$. Here $E_{\mathcal{N}}^{\textit{a}|\textit{m}_{\textit{j}}}$ corresponds to the logarithmic negativity of the bipartition of the optical mode and the $j$th mechanical mode, while $E_{\mathcal{N}}^{\textit{m}_{\textit{1}}|\textit{m}_{\textit{2}}}$ corresponds to that of the bipartition of the two mechanical modes. In terms of the CW input case, as shown in the left panels of Figs. 2(a)-2(d), it is seen that in the absence of a two-photon parametric driving field, i.e., $r_{d}=r_{e}=0$, all of the logarithmic negativity $E_{\mathcal{N}}^{\textit{a}|\textit{m}_{\textit{1}}}$, $E_{\mathcal{N}}^{\textit{a}|\textit{m}_{\textit{2}}}$, $E_{\mathcal{N}}^{\textit{m}_{\textit{1}}|\textit{m}_{\textit{2}}}$ and the minimum residual contangle $\mathcal{R}^{\textrm{min}}_{\tau}$ are null within the considered optical detuning (yellow dashed curve), indicating that there is no bipartite and tripartite entanglement in this system when only applying a coherent driving field. In contrast, for the CCW input case, it is seen that in the presence of a two-photon parametric driving field (with, e.g., squeezing strength $r_{d}=r_{e}=0.3$), the profiles of $E_{\mathcal{N}}^{\textit{a}|\textit{m}_{\textit{1}}}$, $E_{\mathcal{N}}^{\textit{a}|\textit{m}_{\textit{2}}}$, $E_{\mathcal{N}}^{\textit{m}_{\textit{1}}|\textit{m}_{\textit{2}}}$ and $\mathcal{R}^{\textrm{min}}_{\tau}$ are all characterized by a sharp peak around the optical detuning at COM resonance $\Delta_{c}/\omega_{m}\sim\textrm{1}$. Meanwhile, as shown in the right panels of Figs. 2(a)-2(d), the bipartite and tripartite entanglement are also sensitive to the variation of the squeezing reference angle $\theta_{d}$, which all reach their maximum values at an optimal phase angle of $\theta_{d}=\pi$. This result is consistent with that of the previous study Jiao et al. , in which it reveals that the maximum COM interaction and the minimum optical noise are both obtained under the phase-matched condition with $r_{e}-r_{d}=0$ and $\theta_{e}-\theta_{d}=\pm n\pi~{}(n=1,3,5,\ldots)$, resulting in an enhanced COM entanglement. By comparing these results of the CW and CCW input cases, one can intuitively find that by applying a directional two-photon parametric driving field with a proper squeezing phase, the nonreciprocal generation and manipulation of bipartite and tripartite entanglement can be achieved, namely, such entanglement can emerge only for one specific optical input direction when reversing the coherent driving field. Physically, the mechanism behind these results is that the directional input of a two-photon parametric driving field allows one to regulate the effective COM coupling strength for the CW and CCW modes in an asymmetric way, by which one can break the time-reversal symmetry of this system. This feature of the system provides a flexible tool to tailor the parameter condition for entanglement generation based on its optical input direction. Consequently, when preparing the system under a directional entanglement generation condition, nonreciprocal bipartite and tripartite entanglement will be achieved.
After having established nonreciprocal entanglement, we further show how to manipulate the directionality of EPR steering. In Fig. 3, we plot the quantum steerability $\mathcal{G}^{\textit{a}\rightarrow\textit{m}_{\textit{1}}}$, $\mathcal{G}^{\textit{m}_{\textit{1}}\rightarrow\textit{a}}$, $\mathcal{G}^{\textit{m}_{\textit{1}}\rightarrow\textit{m}_{\textit{2}}}$ and $\mathcal{G}^{\textit{m}_{\textit{2}}\rightarrow\textit{m}_{\textit{1}}}$ as a function of the scaled optical detuning $\Delta_{c}/\omega_{m}$ under different squeezing parameter $r_{d}$. Figures 3(a) and 3(b) show the EPR steering between the optical mode and the mechanical mode for opposite optical input directions. For the CW input case ($r_{d}=r_{e}=0$), since the optomechanical entanglement is absent, EPR steering, as a stronger quantum correlaiton, can also not be produced, where $\mathcal{G}^{\textit{a}\rightarrow\textit{m}_{\textit{1}}}$ and $\mathcal{G}^{\textit{m}_{\textit{1}}\rightarrow\textit{a}}$ always keep zero with the variation of the controlling parameters [see scattered dotted curve in Figs. 3(a) and 3(b)]. For the CCW input case, due to the introduction of the squeezed optical mode and the emergence of the optomechanical entanglement, one can create an asymmetric and directional EPR steering by adjusting the squeezing parameters. For instance, it is seen that although when choosing $\theta_{d}=\pi$ and some small values of $r_{d}$, such as $r_{d}=0.15$, the EPR steering still does not exist for both $\textit{a}\rightarrow\textit{m}_{\textit{1}}$ and $\textit{m}_{\textit{1}}\rightarrow\textit{a}$, i.e., $\mathcal{G}^{\textit{a}\rightarrow\textit{m}_{\textit{1}}}=\mathcal{G}^{\textit{m}_{\textit{1}}\rightarrow\textit{a}}=0$ [solid curve in the top panels of Figs. 3(a) and 3(b)]. However, when the squeezing strength $r_{d}$ is continued to be raised, the value of $\mathcal{G}^{\textit{a}\rightarrow\textit{m}_{\textit{1}}}$ will become nonzero and monotonically increase, resulting in an asymmetric one-way EPR steering with $\mathcal{G}^{\textit{a}\rightarrow\textit{m}_{\textit{1}}}>0$ and $\mathcal{G}^{\textit{m}_{\textit{1}}\rightarrow\textit{a}}=0$. Moreover, we also show the EPR steering between the two mechanical modes for opposite optical input directions in Figs. 3(c) and 3(d). For the CW input case, there is no EPR steering for both $\textit{m}_{\textit{1}}\rightarrow\textit{m}_{\textit{2}}$ and $\textit{m}_{\textit{2}}\rightarrow\textit{m}_{\textit{1}}$, i.e., $\mathcal{G}^{\textit{m}_{\textit{1}}\rightarrow\textit{m}_{\textit{2}}}=\mathcal{G}^{\textit{m}_{\textit{2}}\rightarrow\textit{m}_{\textit{1}}}=0$ [see scattered dotted curve in Figs. 3(c) and 3(d)], which is similar to the case discussed above for the optomechanical bipartition. For the CCW input case, accompanied by the occurrence of entanglement, we can get rich properties of EPR steering, where the overall state’s asymmetry is stepwise driven through the no-way regime ($\mathcal{G}^{\textit{m}_{\textit{1}}\rightarrow\textit{m}_{\textit{2}}}=\mathcal{G}^{\textit{m}_{\textit{2}}\rightarrow\textit{m}_{\textit{1}}}=0$), one-way regime ($\mathcal{G}^{\textit{m}_{\textit{1}}\rightarrow\textit{m}_{\textit{2}}}=0$ but $\mathcal{G}^{\textit{m}_{\textit{2}}\rightarrow\textit{m}_{\textit{1}}}>0$), and finally two-way regime ($\mathcal{G}^{\textit{m}_{\textit{1}}\rightarrow\textit{m}_{\textit{2}}}>0$ and $\mathcal{G}^{\textit{m}_{\textit{2}}\rightarrow\textit{m}_{\textit{1}}}>0$) with the increase of the squeezing strength $r_{d}$. From the above discussions, it is found that EPR steering can be generated and manipulated only for the CCW input case but not for the CW input case, which implies the achievement of nonreciprocal EPR steering. These results can be understood as follows: The proposed three-mode COM system has a bilinear cyclic coupling among the optical mode and the two mechanical modes. For an arbitrary bipartition $\mu$ and $\nu$, apart from their direct interaction, there is also an indirect interaction induced by their coupling to the third intermediate mode. Both direct and indirect interaction paths can generate EPR steering between $\mu$ and $\nu$, whose superposition may induce quantum interference effect that depends on the relative phase and strength of the two interaction paths. As discussed previously, by applying a directional two two-photon parametric driving field, one can manipulate the COM coupling strength in an asymmetric way. In this case, the relative strength of the two interaction paths for any bipartition of this system is influenced, based on which the manipulation of the directionality of EPR steering can be achieved by tuning the squeezing strength.
IV Conclusion
In summary, we have presented how to generate and manipulate squeezing-induced nonreciprocal quantum entanglement and asymmetric EPR steering in a three-mode COM system consisting of a bottle WGM resonator, an optical waveguide, and a tapered glass-fiber nanospike. The WGM resonator supports two counterpropagating optical modes, which are uncoupled and degenerate. The coherent driving field from the waveguide is coupled into and out of the WGM resonator via evanescent coupling. Moreover, when placing the nanospike close to the WGM resonator, it excites two mechanical modes due to the COM interaction induced by the WGM evanescent field, whose vibrating directions are orthogonal or parallel to the resonator surface. First, we show that when applying an additional two-photon parametric driving field for one specific input direction of the WGM resonator with a phase-matched squeezed vacuum reservoir, one can introduce a squeezed optical mode and an enhanced COM interaction in this input direction. This directional injection of quantum squeezing breaks the time-reversal symmetry of the system, and it allows one to tailor the parameter condition for entanglement generation with respect to its optical input direction, which is essential for achieving quantum nonreciprocity. Second, based on this unique feature of the system, we find that when reversing the input direction of the coherent driving field, quantum entanglement and the associated EPR steering is regulated in an asymmetric way, whereby a controllable nonreciprocal generation and manipulation of quantum entanglement and EPR steering is achieved. More interestingly, it is also found that by properly tuning the squeezing parameters, the directionality of EPR steering can be well controlled. According to this, we further show that with the increase of the squeezing strength, EPR steering can be stepwise driven from no-way regime, one-way regime to two-way regime. These results might open up new perspectives for the experimental generation and application of nonreciprocal bipartite and tripartite entanglement and asymmetric EPR steering as a precious resource for a variety of nascent quantum technologies ranging from one-sided device-independent quantum key distribution to no-cloning quantum teleportation Branciard et al. (2012); Walk et al. (2016); Li et al. (2024); He et al. (2015b); Chiu et al. (2016).
V Acknowledgment
H.J. is supported by the National Natural Science Foundation of China (NSFC, Grant No. 11935006, 12421005), the National Key R&\&&D Program of China (Grants No. 2024YFE0102400), the Hunan Provincial Major Sci-Tech Program (Grant No. 2023ZJ1010), and the Science and Technology Innovation Program of Hunan Province (Grant No. 2020RC4047). L.-M.K. is supported by the NSFC (Grants No. 12247105, 11935006, 12175060 and 12421005), the Hunan Provincial Major Sci-Tech Program (Grant No. 2023ZJ1010), and the Henan Science and Technology Major Project (Grant No. 241100210400). Y.-F.J. is supported by the NSFC (Grant No. 12405029). D.-Y.W. is supported by the NSFC (Grant No. 12204424) and the China Postdoctoral Science Foundation (Grant No. 2022M722889). L.T. is supported by the NSFC (Grant No. 12305023) and the Sichuan Science and Technology Program (Grant No. 2024NSFSC1353). Y.W. is supported by the NSFC (Grant No. 12205256) and the Henan Provincial Science and Technology Research Project (Grant No. 232102221001). W.-S.B. is supported by the Henan Science and Technology Major Project of the Department of Science & Technology of Henan Province (Grant No. 241100210400).
Appendix A Derivation of the effective Hamiltonian
To diagonalize the optical mode of Hamiltonian (1), one can perform a unitary Bogoliubov transformation with $S(\eta_{d})=\exp[(-\eta_{d}\hat{a}_{\circlearrowleft}^{\dagger 2}+\eta_{d}^{\ast}\hat{a}_{\circlearrowleft}^{2})/2]$, where $\eta_{d}=r_{d}e^{-i\theta_{d}}$ is a complex squeezing parameter with $r_{d}$ the squeezing strength and $\theta_{d}$ the squeezing reference angle. By implementing this Bogoliubov transformation, a squeezed optical mode is introduced and the system Hamiltonian (1) becomes
$\hat{\tilde{H}}_{\circlearrowleft}=$ $S^{\dagger}(\eta_{d})\hat{H}_{\circlearrowleft}S(\eta_{d})$
$=$ $\dfrac{\omega_{m,1}}{2}(\hat{p}_{1}^{2}+\hat{q}_{1}^{2})+\dfrac{\omega_{m,2}}{2}(\hat{p}_{2}^{2}+\hat{q}_{2}^{2})+\chi\hat{q}_{1}\hat{q}_{2}$
$+\left[\Delta_{c}\cosh(2r_{d})-2\Xi_{d}\sinh(2r_{d})\right]\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{a}_{s,\circlearrowleft}$
$+\left[\Xi_{d}\cosh(2r_{d})-\dfrac{\Delta_{c}}{2}\sinh(2r_{d})\right]e^{-i\theta_{d}}\hat{a}_{s,\circlearrowleft}^{\dagger 2}$
$+\left[\Xi_{d}\cosh(2r_{d})-\dfrac{\Delta_{c}}{2}\sinh(2r_{d})\right]e^{i\theta_{d}}\hat{a}_{s,\circlearrowleft}^{2}$
$-g_{1}\cosh(2r_{d})\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{a}_{s,\circlearrowleft}\hat{q}_{1}-g_{2}\cosh(2r_{d})\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{a}_{s,\circlearrowleft}\hat{q}_{2}$
$+\dfrac{1}{2}g_{1}\sinh(2r_{d})(e^{-i\theta_{d}}\hat{a}_{s,\circlearrowleft}^{\dagger 2}+e^{i\theta_{d}}\hat{a}_{s,\circlearrowleft}^{2})\hat{q}_{1}$
$+\dfrac{1}{2}g_{2}\sinh(2r_{d})(e^{-i\theta_{d}}\hat{a}_{s,\circlearrowleft}^{\dagger 2}+e^{i\theta_{d}}\hat{a}_{s,\circlearrowleft}^{2})\hat{q}_{2}$
$-g_{1}\sinh^{2}(r_{d})\hat{q}_{1}-g_{2}\sinh^{2}(r_{d})\hat{q}_{2}$
$+i\varepsilon_{d,\circlearrowleft}\cosh(r_{d})(\hat{a}_{s,\circlearrowleft}^{\dagger}-\hat{a}_{s,\circlearrowleft})$
$+i\varepsilon_{d,\circlearrowleft}\sinh(r_{d})(e^{-i\theta_{d}}\hat{a}_{s,\circlearrowleft}^{\dagger}-e^{i\theta_{d}}\hat{a}_{s,\circlearrowleft})$
$+\Delta_{c}\sinh^{2}(r_{d})-\Xi_{d}\sinh(2r_{d}).$ (22)
By setting the coefficients of the quadratic terms $\hat{a}_{s,\circlearrowleft}^{\dagger 2}$ and $\hat{a}_{s,\circlearrowleft}^{2}$ to be zero, i.e.,
$\Delta_{c}\cosh(2r_{d})-2\Xi_{d}\sinh(2r_{d})=0,$ (23)
we have $r_{d}=(1/4)\ln[(\Delta_{c}+2\Xi_{d})/(\Delta_{c}-2\Xi_{d})]$. Correspondingly, the effective Hamiltonian of the system is obtained as
$\hat{\tilde{H}}_{\circlearrowleft}=$ $\omega_{s}\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{a}_{s,\circlearrowleft}+\dfrac{\omega_{m,1}}{2}(\hat{p}_{1}^{2}+\hat{q}_{1}^{2})+\dfrac{\omega_{m,2}}{2}(\hat{p}_{2}^{2}+\hat{q}_{2}^{2})$
$-\zeta_{s,1}\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{a}_{s,\circlearrowleft}\hat{q}_{1}+\dfrac{\zeta_{p,1}}{2}(e^{-i\theta_{d}}\hat{a}_{s,\circlearrowleft}^{\dagger 2}+e^{i\theta_{d}}\hat{a}_{s,\circlearrowleft}^{2})\hat{q}_{1}$
$-\zeta_{s,2}\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{a}_{s,\circlearrowleft}\hat{q}_{2}+\dfrac{\zeta_{p,2}}{2}(e^{-i\theta_{d}}\hat{a}_{s,\circlearrowleft}^{\dagger 2}+e^{i\theta_{d}}\hat{a}_{s,\circlearrowleft}^{2})\hat{q}_{2}$
$\!-\!F_{1}\hat{q}_{1}\!-\!F_{2}\hat{q}_{2}\!+\!\chi\hat{q}_{1}\hat{q}_{2}\!+\!i\varepsilon_{d,\circlearrowleft}\!\cosh(r_{d})(\hat{a}_{s,\circlearrowleft}^{\dagger}\!-\!\hat{a}_{s,\circlearrowleft})$
$+i\varepsilon_{d,\circlearrowleft}\sinh(r_{d})(e^{-i\theta_{d}}\hat{a}_{s,\circlearrowleft}^{\dagger}-e^{i\theta_{d}}\hat{a}_{s,\circlearrowleft})+C,$ (24)
where
$\omega_{s}\!$ $=\!\Delta_{c}\cosh(2r_{d})\!-\!2\Xi_{d}\sinh(2r_{d})\!=\!(\Delta_{c}\!-\!2\Xi_{d})\exp(2r_{d}),$
$\zeta_{s,j}$ $=\dfrac{g_{j}\Delta_{c}}{\sqrt{\Delta_{c}^{2}-4\Xi_{d}^{2}}}=g_{j}\cosh(2r_{d}),$
$\zeta_{p,j}$ $=\dfrac{2g_{j}\Xi_{d}}{\sqrt{\Delta_{c}^{2}-4\Xi_{d}^{2}}}=g_{j}\sinh(2r_{d}),$
$F_{j}$ $=g_{j}\sinh^{2}(r_{d}),~{}C=\Delta_{c}\sinh^{2}(r_{d})-\Xi_{d}\sinh(2r_{d}).$ (25)
Appendix B Derivation of the effective master equation under phase-matched condition
For the CCW input case, to eliminate the optical dissipation induced by the intracavity squeezing, we consider the injection of a broadband squeezed optical field in this direction, which can act as a squeezed-vacuum reservoir. We assume the squeezed optical field, with squeezing strength $r_{e}$ and reference phase angle $\theta_{e}$, is around the central frequency $\omega_{c}$. Besides, the two mechanical modes are assumed to be coupled with two independent thermal reservoirs at the same bath temperature $T$. Then, including the dissipations of the optical and mechanical modes, the dynamics of the total system in the original picture is governed by the Born-Markovian master equation Gardiner and Zoller (2000)
$\dot{\rho}_{\circlearrowleft}=$ $i\big{[}\rho_{\circlearrowleft},\hat{H}_{\circlearrowleft}\big{]}+\frac{\kappa}{2}(N_{e}+1)\mathcal{D}[\hat{a}_{\circlearrowleft}]\rho_{\circlearrowleft}+\frac{\kappa}{2}N_{e}\mathcal{D}[\hat{a}_{\circlearrowleft}^{\dagger}]\rho_{\circlearrowleft}$
$-\frac{\kappa}{2}M_{e}\mathcal{G}[\hat{a}_{\circlearrowleft}]\rho_{\circlearrowleft}-\frac{\kappa}{2}M_{e}^{\ast}\mathcal{G}[\hat{a}_{\circlearrowleft}^{\dagger}]\rho_{\circlearrowleft}$
$-\sum_{j=1,2}\left(i\dfrac{\gamma_{m,j}}{2}[\hat{q}_{j},\{\hat{p}_{j},\rho_{\circlearrowleft}\}]+\gamma_{m,j}\bar{n}_{m,j}[\hat{q}_{j},[\hat{q}_{j},\rho_{\circlearrowleft}]]\right),$ (26)
where
$\mathcal{D}[\hat{a}_{\circlearrowleft}]\rho_{\circlearrowleft}$ $=2\hat{a}_{\circlearrowleft}\rho_{\circlearrowleft}\hat{a}_{\circlearrowleft}^{\dagger}-(\hat{a}_{\circlearrowleft}^{\dagger}\hat{a}_{\circlearrowleft}\rho_{\circlearrowleft}+\rho_{\circlearrowleft}\hat{a}_{\circlearrowleft}^{\dagger}\hat{a}_{\circlearrowleft}),$
$\mathcal{G}[\hat{a}_{\circlearrowleft}]\rho_{\circlearrowleft}$ $=2\hat{a}_{\circlearrowleft}\rho_{\circlearrowleft}\hat{a}_{\circlearrowleft}-(\hat{a}_{\circlearrowleft}\hat{a}_{\circlearrowleft}\rho_{\circlearrowleft}+\rho_{\circlearrowleft}\hat{a}_{\circlearrowleft}\hat{a}_{\circlearrowleft}),$ (27)
and $\kappa~{}(\gamma_{m,j})$ is the optical (mechanical) decay rate. $\bar{n}_{m,j}=1/[\exp(\omega_{m,j}/k_{B}T)-1]$ is the thermal phonon number, while $N_{e}=\sinh^{2}(r_{e})$ and $M_{e}=\cosh(r_{e})\sinh(r_{e})e^{i\theta_{e}}$ describe the dissipation and the two-photon correlation of the cavity field caused by the squeezed-vacuum reservoir, respectively.
By introducing the Bogoliubov transformation in Eq. (2), the Lindblad operators becomes
$S^{\dagger}(\eta_{d})\mathcal{D}[\hat{a}_{\circlearrowleft}]\rho_{\circlearrowleft}S(\eta_{d})$
$=$ $2\cosh^{2}(r_{d})\hat{a}_{s,\circlearrowleft}\tilde{\rho}_{\circlearrowleft}\hat{a}_{s,\circlearrowleft}^{\dagger}+2\sinh^{2}(r_{d})\hat{a}_{s,\circlearrowleft}^{\dagger}\tilde{\rho}_{\circlearrowleft}\hat{a}_{s,\circlearrowleft}$
$-\sinh(2r_{d})(e^{-i\theta_{d}}\hat{a}_{s,\circlearrowleft}^{\dagger}\tilde{\rho}_{\circlearrowleft}\hat{a}_{s,\circlearrowleft}^{\dagger}+e^{i\theta_{d}}\hat{a}_{s,\circlearrowleft}\tilde{\rho}_{\circlearrowleft}\hat{a}_{s,\circlearrowleft})$
$-\left[\cosh^{2}(r_{d})\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{a}_{s,\circlearrowleft}\tilde{\rho}_{\circlearrowleft}+\sinh^{2}(r_{d})\hat{a}_{s,\circlearrowleft}\hat{a}_{s,\circlearrowleft}^{\dagger}\tilde{\rho}_{\circlearrowleft}\right.$
$\left.-\sinh(r_{d})\cosh(r_{d})(e^{-i\theta_{d}}\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{a}_{s,\circlearrowleft}^{\dagger}\tilde{\rho}_{\circlearrowleft}+e^{i\theta_{d}}\hat{a}_{s,\circlearrowleft}\hat{a}_{s,\circlearrowleft}\tilde{\rho}_{\circlearrowleft})\right]$
$-\left[\cosh^{2}(r_{d})\tilde{\rho}_{\circlearrowleft}\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{a}_{s,\circlearrowleft}+\sinh^{2}(r_{d})\tilde{\rho}_{\circlearrowleft}\hat{a}_{s,\circlearrowleft}\hat{a}_{s,\circlearrowleft}^{\dagger}\right.$
$\left.-\sinh(r_{d})\cosh(r_{d})(e^{-i\theta_{d}}\tilde{\rho}_{\circlearrowleft}\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{a}_{s,\circlearrowleft}^{\dagger}+e^{i\theta_{d}}\tilde{\rho}_{\circlearrowleft}\hat{a}_{s,\circlearrowleft}\hat{a}_{s,\circlearrowleft})\right],$
$S^{\dagger}(\eta_{d})\mathcal{G}[\hat{a}_{\circlearrowleft}]\rho_{\circlearrowleft}S(\eta_{d})$
$=$ $2e^{-2i\theta_{d}}\sinh^{2}(r_{d})\hat{a}_{s,\circlearrowleft}^{\dagger}\tilde{\rho}_{\circlearrowleft}\hat{a}_{s,\circlearrowleft}^{\dagger}+2\cosh^{2}(r_{d})\hat{a}_{s,\circlearrowleft}\tilde{\rho}_{\circlearrowleft}\hat{a}_{s,\circlearrowleft}$
$-e^{-i\theta_{d}}\sinh(2r_{d})(\hat{a}_{s,\circlearrowleft}^{\dagger}\tilde{\rho}_{\circlearrowleft}\hat{a}_{s,\circlearrowleft}+\hat{a}_{s,\circlearrowleft}\tilde{\rho}_{\circlearrowleft}\hat{a}_{s,\circlearrowleft}^{\dagger})$
$-\left[e^{-2i\theta_{d}}\sinh^{2}(r_{d})\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{a}_{s,\circlearrowleft}^{\dagger}\tilde{\rho}_{\circlearrowleft}+\cosh^{2}(r_{d})\hat{a}_{s,\circlearrowleft}\hat{a}_{s,\circlearrowleft}\tilde{\rho}_{\circlearrowleft}\right.$
$\left.-e^{-i\theta_{d}}\sinh(r_{d})\cosh(r_{d})(\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{a}_{s,\circlearrowleft}\tilde{\rho}_{\circlearrowleft}+\hat{a}_{s,\circlearrowleft}\hat{a}_{s,\circlearrowleft}^{\dagger}\tilde{\rho}_{\circlearrowleft})\right]$
$-\left[e^{-2i\theta_{d}}\sinh^{2}(r_{d})\tilde{\rho}_{\circlearrowleft}\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{a}_{s,\circlearrowleft}^{\dagger}+\cosh^{2}(r_{d})\tilde{\rho}_{\circlearrowleft}\hat{a}_{s,\circlearrowleft}\hat{a}_{s,\circlearrowleft}\right.$
$\left.-e^{-i\theta_{d}}\sinh(r_{d})\cosh(r_{d})(\tilde{\rho}_{\circlearrowleft}\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{a}_{s,\circlearrowleft}+\tilde{\rho}_{\circlearrowleft}\hat{a}_{s,\circlearrowleft}\hat{a}_{s,\circlearrowleft}^{\dagger})\right].$ (28)
In this case, the effective Born-Markovian master equation could be rewritten as
$\dot{\tilde{\rho}}_{\circlearrowleft}=$ $S^{\dagger}(\eta_{d})\dot{\rho}_{\circlearrowleft}S(\eta_{d})$
$=$ $i\big{[}\tilde{\rho}_{\circlearrowleft},\hat{\tilde{H}}_{\circlearrowleft}\big{]}+\frac{\kappa}{2}(N_{s}+1)\mathcal{D}[\hat{a}_{s,\circlearrowleft}]\tilde{\rho}_{\circlearrowleft}+\frac{\kappa}{2}N_{s}\mathcal{D}[\hat{a}_{s,\circlearrowleft}^{\dagger}]\tilde{\rho}_{\circlearrowleft}$
$-\frac{\kappa}{2}M_{s}\mathcal{G}[\hat{a}_{s,\circlearrowleft}]\tilde{\rho}_{\circlearrowleft}-\frac{\kappa}{2}M_{s}^{\ast}\mathcal{G}[\hat{a}_{s,\circlearrowleft}]\tilde{\rho}_{\circlearrowleft}$
$-\sum_{j=1,2}\left(i\dfrac{\gamma_{m,j}}{2}[\hat{q}_{j},\{\hat{p}_{j},\tilde{\rho}_{\circlearrowleft}\}]+\gamma_{m,j}\bar{n}_{m,j}[\hat{q}_{j},[\hat{q}_{j},\tilde{\rho}_{\circlearrowleft}]]\right),$ (29)
where
$N_{s}=$ $\sinh^{2}(r_{d})\cosh^{2}(r_{e})+\cosh^{2}(r_{d})\sinh^{2}(r_{e})$
$+\dfrac{1}{2}\cos(\theta_{e}-\theta_{d})\sinh(2r_{d})\sinh(2r_{e}),$
$M_{s}=$ $e^{i\theta_{d}}[\cosh(r_{d})\cosh(r_{e})+e^{-i(\theta_{e}-\theta_{d})}\sinh(r_{d})\sinh(r_{e})]$
×$\times[\sinh(r_{d})\cosh(r_{e})+e^{i(\theta_{e}-\theta_{d})}\cosh(r_{d})\sinh(r_{e})].$ (30)
Here $N_{s}$ and $M_{s}$ denote the effective thermal noise and two-photon-correlation strength, respectively, which can be simplified in case of $r_{d}=r_{e}=r$, i.e.,
$N_{s}=$ $\dfrac{1}{2}\sinh^{2}(2r)[1+\cos(\theta_{e}-\theta_{d})],$
$M_{s}=$ $\dfrac{1}{2}e^{i\theta_{d}}\sinh(2r)[1+e^{i(\theta_{e}-\theta_{d})}]$
×$\times[\cosh^{2}(r)+e^{-i(\theta_{e}-\theta_{d})}\sinh^{2}(r)].$ (31)
Obviously, the thermal noise $N_{s}$ and the two-photon-correlation strength $M_{s}$ can be completely eliminated by choosing $r_{p}=r_{e}$ and $\theta_{e}-\theta_{d}=\pm n\pi(n=1,3,5...)$, i.e., $N_{s}=M_{s}=0$. As discussed in Ref. , the reservoir of the original optical mode is squeezed along the axis with angle $\dfrac{\theta_{e}}{2}$. In the basis of the squeezed optical mode $\hat{a}_{s,\circlearrowleft}$, this effect can be canceled by the squeezing (along axis $\dfrac{\theta_{d}}{2}$) induced by the parametric amplification of $\hat{a}_{\circlearrowleft}$, when $r_{p}=r_{e}$ and $\theta_{e}+\theta_{d}=\pm n\pi(n=1,3,5...)$, yielding the phase-matched condition. This means the squeezed-vacuum reservoir of $\hat{a}_{\circlearrowleft}$ corresponds to an effective vacuum reservoir of $\hat{a}_{s,\circlearrowleft}$ under the phase-matched condition. In this case, the effective Born-Markovian master equation is derived as
$\dot{\tilde{\rho}}_{\circlearrowleft}\!=$ $i\!\left[\tilde{\rho}_{\circlearrowleft},\hat{\tilde{H}}_{\circlearrowleft}\right]\!+\!\frac{\kappa}{2}\mathcal{D}[\hat{a}_{s,\circlearrowleft}]\tilde{\rho}_{\circlearrowleft}\!-\!\sum_{j=1,2}\left(i\dfrac{\gamma_{m,j}}{2}[\hat{q}_{j},\{\hat{p}_{j},\tilde{\rho}_{\circlearrowleft}\}]\right.$
$\left.+\gamma_{m,j}\bar{n}_{m,j}[\hat{q}_{j},[\hat{q}_{j},\tilde{\rho}_{\circlearrowleft}]]\right),$ (32)
where
$\mathcal{D}[\hat{a}_{s,\circlearrowleft}]\tilde{\rho}_{\circlearrowleft}=2\hat{a}_{s,\circlearrowleft}\tilde{\rho}_{\circlearrowleft}\hat{a}_{s,\circlearrowleft}^{\dagger}-(\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{a}_{s,\circlearrowleft}\tilde{\rho}_{\circlearrowleft}+\tilde{\rho}_{\circlearrowleft}\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{a}_{s,\circlearrowleft})$ (33)
[…]
Appendix C Derivation of the dynamics of the quantum correlation function
By employing the linearized effective master equation (11), the dynamics of the quantum correlation function in the vector X can be derived as
$\dot{x}_{1}=-\kappa x_{1}+i\Lambda_{1}x_{14}-i\Lambda_{1}^{\ast}x_{13}+i\Lambda_{2}x_{18}-i\Lambda_{2}^{\ast}x_{17},$ (34)
$\dot{x}_{2}=$ $-\kappa x_{2}+i\Lambda_{1}x_{14}-i\Lambda_{1}^{\ast}x_{13}+i\Lambda_{2}x_{18}-i\Lambda_{2}^{\ast}x_{17}+\kappa,$ (35)
$\dot{x}_{3}=2\omega_{m,1}x_{9}+2\omega_{m,1}x_{10},$ (36)
$\dot{x}_{4}=$ $-2\gamma_{m,1}x_{4}-2\omega_{m,1}x_{9}-2\omega_{m,1}x_{10}+2\Lambda_{1}x_{16}$
$+2\Lambda_{1}^{\ast}x_{15}-2\chi x_{24}+2\gamma_{m,1}\bar{n}_{m,1},$ (37)
$\dot{x}_{5}=2\omega_{m,2}x_{11}+2\omega_{m,2}x_{12},$ (38)
$\dot{x}_{6}=$ $-2\gamma_{m,2}x_{6}-2\omega_{m,2}x_{11}-2\omega_{m,2}x_{12}+2\Lambda_{2}x_{20}$
$+2\Lambda_{2}^{\ast}x_{19}-2\chi x_{23}+2\gamma_{m,2}\bar{n}_{m,2},$ (39)
$\dot{x}_{7}=-(2i\Delta_{s}+\kappa)x_{7}+2i\Lambda_{1}x_{13}+2i\Lambda_{2}x_{17},$ (40)
$\dot{x}_{8}=(2i\Delta_{s}-\kappa)x_{8}-2i\Lambda_{1}^{\ast}x_{14}-2i\Lambda_{2}^{\ast}x_{18},$ (41)
$\dot{x}_{9}=$ $-\dfrac{\gamma_{m,1}}{2}x_{9}-\dfrac{\gamma_{m,1}}{2}x_{10}+\omega_{m,1}x_{4}-\omega_{m,1}x_{3}$
$-\chi x_{21}+\Lambda_{1}x_{14}+\Lambda_{1}^{\ast}x_{13},$ (42)
$\dot{x}_{10}=$ $-\dfrac{\gamma_{m,1}}{2}x_{10}-\dfrac{\gamma_{m,1}}{2}x_{9}+\omega_{m,1}x_{4}-\omega_{m,1}x_{3}$
$-\chi x_{21}+\Lambda_{1}x_{14}+\Lambda_{1}^{\ast}x_{13},$ (43)
$\dot{x}_{11}=$ $-\dfrac{\gamma_{m,2}}{2}x_{11}-\dfrac{\gamma_{m,2}}{2}x_{12}+\omega_{m,2}x_{6}-\omega_{m,2}x_{5}$
$-\chi x_{21}+\Lambda_{2}x_{18}+\Lambda_{2}^{\ast}x_{17},$ (44)
$\dot{x}_{12}=$ $-\dfrac{\gamma_{m,2}}{2}x_{12}-\dfrac{\gamma_{m,2}}{2}x_{11}+\omega_{m,2}x_{6}-\omega_{m,2}x_{5}$
$-\chi x_{21}+\Lambda_{2}x_{18}+\Lambda_{2}^{\ast}x_{17},$ (45)
$\dot{x}_{13}=-\left(i\Delta_{s}+\dfrac{\kappa}{2}\right)x_{13}+\omega_{m,1}x_{15}+i\Lambda_{1}x_{3}+i\Lambda_{2}x_{21},$ (46)
$\dot{x}_{14}=\left(i\Delta_{s}-\dfrac{\kappa}{2}\right)x_{14}+\omega_{m,1}x_{16}-i\Lambda_{1}^{\ast}x_{3}-i\Lambda_{2}^{\ast}x_{21},$ (47)
$\dot{x}_{15}=$ $-\left(i\Delta_{s}+\gamma_{m,1}+\dfrac{\kappa}{2}\right)x_{15}-\omega_{m,1}x_{13}+i\Lambda_{1}x_{10}$
$+i\Lambda_{2}x_{24}-\chi x_{17},$ (48)
$\dot{x}_{16}=$ $\left(i\Delta_{s}-\gamma_{m,1}-\dfrac{\kappa}{2}\right)x_{16}-\omega_{m,1}x_{14}-i\Lambda_{1}^{\ast}x_{10}$
$-i\Lambda_{2}^{\ast}x_{24}-\chi x_{18},$ (49)
$\dot{x}_{17}=-\left(i\Delta_{s}+\dfrac{\kappa}{2}\right)x_{17}+\omega_{m,1}x_{19}+i\Lambda_{1}x_{21}+i\Lambda_{2}x_{5},$ (50)
$\dot{x}_{18}=\left(i\Delta_{s}-\dfrac{\kappa}{2}\right)x_{18}+\omega_{m,1}x_{20}-i\Lambda_{1}^{\ast}x_{21}-i\Lambda_{2}^{\ast}x_{5},$ (51)
$\dot{x}_{19}=$ $-\left(i\Delta_{s}+\gamma_{m,2}+\dfrac{\kappa}{2}\right)x_{19}-\omega_{m,2}x_{17}+i\Lambda_{1}x_{23}$
$+i\Lambda_{2}x_{12}-\chi x_{13},$ (52)
$\dot{x}_{20}=$ $\left(i\Delta_{s}-\gamma_{m,2}-\dfrac{\kappa}{2}\right)x_{20}-\omega_{m,2}x_{18}-i\Lambda_{1}^{\ast}x_{23}$
$-i\Lambda_{2}^{\ast}x_{12}-\chi x_{14},$ (53)
$\dot{x}_{21}=\omega_{m,1}x_{24}+\omega_{m,2}x_{23},$ (54)
$\dot{x}_{22}=$ $-(\gamma_{m,1}+\gamma_{m,2})x_{22}-\omega_{m,1}x_{23}-\omega_{m,2}x_{24}-\chi x_{10}$
$-\chi x_{11}+\Lambda_{1}x_{20}+\Lambda_{1}^{\ast}x_{19}+\Lambda_{2}x_{16}+\Lambda_{2}^{\ast}x_{15},$ (55)
$\dot{x}_{23}=$ $-\gamma_{m,2}x_{23}+\omega_{m,1}x_{22}-\omega_{m,2}x_{21}-\chi x_{3}$
$+\Lambda_{2}x_{14}+\Lambda_{2}^{\ast}x_{13},$ (56)
$\dot{x}_{24}=$ $-\gamma_{m,1}x_{24}-\omega_{m,1}x_{21}+\omega_{m,2}x_{22}-\chi x_{5}$
$+\Lambda_{1}x_{18}+\Lambda_{1}^{\ast}x_{17}.$ (57)
By rewriting the above equations in a compact form, i.e.,
$\dot{X}=M\cdot X+N,$ (58)
and introducing the following expressions
$\Omega_{1,\pm}$ $=i\Delta_{s}\pm\dfrac{\kappa}{2},$
$\Omega_{2,\pm}$ $=i\Delta_{s}\pm\gamma_{m,1}\pm\dfrac{\kappa}{2},$
$\Omega_{3,\pm}$ $=i\Delta_{s}\pm\gamma_{m,2}\pm\dfrac{\kappa}{2}.,$ (59)
we have
$N=$ $\left[0,\kappa,0,2\gamma_{m,1}\bar{n}_{m,1},0,2\gamma_{m,2}\bar{n}_{m,2},0,0,\right.$
$\left.0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0\right]^{T},$ (60)
and
$M=$
($\begin{smallmatrix}-\kappa&0&0&0&0&0&0&0&0&0&0&0&-i\Lambda_{1}^{\ast}&i\Lambda_{1}&0&0\\ -i\Lambda_{2}^{\ast}&i\Lambda_{2}&0&0&0&0&0&0\\ \\ 0&-\kappa&0&0&0&0&0&0&0&0&0&0&-i\Lambda_{1}^{\ast}&i\Lambda_{1}&0&0\\ -i\Lambda_{2}^{\ast}&i\Lambda_{2}&0&0&0&0&0&0\\ \\ 0&0&0&0&0&0&0&0&2\omega_{m,1}&2\omega_{m,1}&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0\\ \\ 0&0&0&-2\gamma_{m,1}&0&0&0&0&-2\omega_{m,1}&-2\omega_{m,1}&0&0&0&0&2\Lambda_{1}^{\ast}&2\Lambda_{1}\\ 0&0&0&0&0&0&0&-2\chi\\ \\ 0&0&0&0&0&0&0&0&0&0&2\omega_{m,2}&2\omega_{m,2}&0&0&0&0\\ 0&0&0&0&0&0&0&0\\ \\ 0&0&0&0&0&-2\gamma_{m,2}&0&0&0&0&-2\omega_{m,2}&-2\omega_{m,2}&0&0&0&0\\ 0&0&2\Lambda_{2}^{\ast}&2\Lambda_{2}&0&0&-2\chi&0\\ \\ 0&0&0&0&0&0&-2\Omega_{1,+}&0&0&0&0&0&2i\Lambda_{1}&0&0&0\\ 2i\Lambda_{2}&0&0&0&0&0&0&0\\ \\ 0&0&0&0&0&0&0&2\Omega_{1,-}&0&0&0&0&0&-2i\Lambda_{1}^{\ast}&0&0\\ 0&-2i\Lambda_{2}^{\ast}&0&0&0&0&0&0\\ \\ 0&0&-\omega_{m,1}&\omega_{m,1}&0&0&0&0&-\gamma_{m,1}/2&-\gamma_{m,1}/2&0&0&\Lambda_{1}^{\ast}&\Lambda_{1}&0&0\\ 0&0&0&0&-\chi&0&0&0\\ \\ 0&0&-\omega_{m,1}&\omega_{m,1}&0&0&0&0&-\gamma_{m,1}/2&-\gamma_{m,1}/2&0&0&\Lambda_{1}^{\ast}&\Lambda_{1}&0&0\\ 0&0&0&0&-\chi&0&0&0\\ \\ 0&0&0&0&-\omega_{m,2}&\omega_{m,2}&0&0&0&0&-\gamma_{m,2}/2&-\gamma_{m,2}/2&0&0&0&0\\ \Lambda_{2}^{\ast}&\Lambda_{2}&0&0&-\chi&0&0&0\\ \\ 0&0&0&0&-\omega_{m,2}&\omega_{m,2}&0&0&0&0&-\gamma_{m,2}/2&-\gamma_{m,2}/2&0&0&0&0\\ \Lambda_{2}^{\ast}&\Lambda_{2}&0&0&-\chi&0&0&0\\ \\ 0&0&i\Lambda_{1}&0&0&0&0&0&0&0&0&0&-\Omega_{1,+}&0&\omega_{m,1}&0\\ 0&0&0&0&i\Lambda_{2}&0&0&0\\ \\ 0&0&-i\Lambda_{1}^{\ast}&0&0&0&0&0&0&0&0&0&0&\Omega_{1,-}&0&\omega_{m,1}\\ 0&0&0&0&-i\Lambda_{2}^{\ast}&0&0&0\\ \\ 0&0&0&0&0&0&0&0&0&i\Lambda_{1}&0&0&-\omega_{m,1}&0&-\Omega_{2,+}&0\\ -\chi&0&0&0&0&0&0&i\Lambda_{2}\\ \\ 0&0&0&0&0&0&0&0&0&-i\Lambda_{1}^{\ast}&0&0&0&-\omega_{m,1}&0&\Omega_{2,-}\\ 0&-\chi&0&0&0&0&0&-i\Lambda_{2}^{\ast}\\ \\ 0&0&0&0&i\Lambda_{2}&0&0&0&0&0&0&0&0&0&0&0\\ -\Omega_{1,+}&0&\omega_{m,1}&0&i\Lambda_{1}&0&0&0\\ \\ 0&0&0&0&-i\Lambda_{2}^{\ast}&0&0&0&0&0&0&0&0&0&0&0\\ 0&\Omega_{1,-}&0&\omega_{m,1}&-i\Lambda_{1}^{\ast}&0&0&0\\ \\ 0&0&0&0&0&0&0&0&0&0&0&i\Lambda_{2}&-\chi&0&0&0\\ -\omega_{m,2}&0&-\Omega_{3,+}&0&0&0&i\Lambda_{1}&0\\ \\ 0&0&0&0&0&0&0&0&0&0&0&-i\Lambda_{2}^{\ast}&0&-\chi&0&0\\ 0&-\omega_{m,2}&0&\Omega_{3,-}&0&0&-i\Lambda_{1}^{\ast}&0\\ \\ 0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&\omega_{m,2}&\omega_{m,1}\\ 0&0&0&0&0&0&0&0&0&-\chi&-\chi&0&0&0&\Lambda_{2}^{\ast}&\Lambda_{2}\\ 0&0&\Lambda_{1}^{\ast}&\Lambda_{1}&0&-\gamma_{m,1}-\gamma_{m,2}&-\omega_{m,1}&-\omega_{m,2}\\ 0&0&-\chi&0&0&0&0&0&0&0&0&0&\Lambda_{2}^{\ast}&\Lambda_{2}&0&0\\ 0&0&0&0&-\omega_{m,2}&\omega_{m,1}&-\gamma_{m,2}&0\\ 0&0&0&0&-\chi&0&0&0&0&0&0&0&0&0&0&0\\ \Lambda_{1}^{\ast}&\Lambda_{1}&0&0&-\omega_{m,1}&\omega_{m,2}&0&-\gamma_{m,1}\\ \end{smallmatrix}\right).$

### v2

@@
-Nonreciprocal tripartite entanglement and asymmetric Einstein-Podolsky-Rosen steering via directional quantum squeezing
+Phase-selective tripartite entanglement and asymmetric Einstein-Podolsky-Rosen steering in squeezed optomechanics
 Abstract
 The generation and manipulation of multipartite entanglement and EPR steering in macroscopic systems not only play a fundamental role in exploring the nature of quantum mechanics, but are also at the core of current developments of various nascent quantum technologies.
-Here we report a theoretical method using directional injection of quantum squeezing to produce nonreciprocal multipartite entanglement and EPR steering in a three-mode optomechanical system with closed-loop coupling.
-We show that by directionally applying a two-photon parametric driving field with a phase-matched squeezed vacuum reservoir to an optomechanical resonator, a squeezed optical mode can be introduced for one of its input directions, thereby yielding an asymmetric enhancement of optomechanical interaction and the time-reversal symmetry breaking of the system.
-Based on this feature, it is found that bipartite and tripartite entanglement and the associated EPR steering of the subsystems can only be generated when the coherent driving field input from the squeezing injection direction, namely, achieving nonreciprocity in such quantum correlations.
-More excitingly, it is also found that by properly adjusting the squeezing parameter, the overall asymmetry of EPR steering can be stepwise driven from no-way regime, one-way regime to two-way regime.
-These findings, holding promise for preparing rich types of entangled quantum resources with nonreciprocal correlations, may have potential applications in the area of quantum information processing such as quantum secure direct communication and one-way quantum computing.
+Here we report a theoretical method using squeezing-phase-controlled quantum noise flows to selectively generate and manipulate quantum entanglement and asymmetric EPR steering in a nonlinear $\chi^{(2)}$ whispering-gallery-mode (WGM) optomechanical resonator.
+We show that by pumping the $\chi^{(2)}$ nonlinear medium with two-photon optical fields and broadband squeezed lights, a pair of counterpropagating squeezed optical modes could be introduced to the WGM resonator, each coupled with an independent squeezed vacuum reservoir.
+This configuration could enable squeezing-phase-controlled light-reservoir interaction for each squeezed optical mode, providing a flexible tool for tailoring asymmetric optical noise flows in the counterpropagating modes.
+Based on this unique feature, it is found that with the injection of asymmetric noise flows, the generation of various types of bipartite and tripartite entanglement become phase-dependent and thus they can be produced in an asymmetric way.
+More excitingly, it is also found that by further properly adjusting the squeezing parameters, the overall asymmetry of EPR steering can also be stepwise driven from no-way regime, one-way regime to two-way regime.
+These findings, holding promise for preparing rich types of entangled quantum resources with asymmetric features, may have potential applications in the area of secure quantum information processing such as quantum secure direct communication and one-way quantum computing.
 I Introduction
-Entanglement, allowing perfectly correlated positions and momenta for two spatially separated particles, has long been intriguing in quantum physics and enables numerous advanced quantum information protocols spanning from quantum networking to quantum sensing Horodecki et al.
-(2009).
-The concept of entanglement was originally addressed by Schrödinger Schrödinger (1935) in his response to the issue of the “spooky action-at-a-distance” predicted by Einstein, Podolsky, and Rosen (EPR) in their famous paradox Einstein et al.
-(1935), where he also coined a term that came to be known as the EPR steering Reid (1989); Ou et al.
-(1992).
+Entanglement, allowing perfectly correlated positions and momenta for two spatially separated particles, has long been intriguing in quantum physics and enables numerous advanced quantum information protocols spanning from quantum networking to quantum sensing [1].
+The concept of entanglement was originally addressed by Schrödinger [2] in his response to the issue of the “spooky action-at-a-distance” predicted by Einstein, Podolsky, and Rosen (EPR) in their famous paradox [3], where he also coined a term that came to be known as the EPR steering [4, 5].
 From the perspective of violations of local-hidden-state models, Wiseman et al.
-has formalized an operational benchmark for EPR steering Wiseman et al.
-(2007), by which they further proved that under the hierarchy of quantum nonlocality, EPR steering is a strict subset of entanglement and a strict superset of Bell nonlocality.
+has formalized an operational benchmark for EPR steering [6], by which they further proved that under the hierarchy of quantum nonlocality, EPR steering is a strict subset of entanglement and a strict superset of Bell nonlocality.
 An appealing feature of EPR steering is that it describes how local measurements on one part of the system can steer (alter) the state of the other part at a different location.
-This defining characteristics, not held by the other two types of quantum nonlocality, reveals the intrinsic asymmetry of EPR steering and offers an insight into directional nonlocality He et al.
-(2015a), which plays an indispensable role in enabling quantum techniques using untrusted devices Xiang et al.
-(2022), such as secure quantum key distribution Branciard et al.
-(2012); Walk et al.
-(2016), randomness certification Li et al.
-(2024), and no-cloning quantum teleportation He et al.
-(2015b); Chiu et al.
-(2016).
-In the past few decades, after a series of rigorous mathematical characterizations of quantum nonlocality Simon (2000); Adesso et al.
-(2004); Adesso and Illuminati (2007); Kogias et al.
-(2015), a great deal of progresses have been made experimentally to prepare entangled or steerable states of microscopic and macroscopic particles, involving platforms based on photons Wang et al.
-(2016a), ions Stute et al.
-(2012), atoms van Leent et al.
-(2022), superconducting circuits Kurpiers et al.
-(2018), and cavity optomechanical (COM) devices Palomaki et al.
-(2013); Kotler et al.
-(2021); Mercier de Lépinay et al.
-(2021).
+This defining characteristics, not held by the other two types of quantum nonlocality, reveals the intrinsic asymmetry of EPR steering and offers an insight into directional nonlocality [7], which plays an indispensable role in enabling quantum techniques using untrusted devices [8], such as secure quantum key distribution [9, 10], randomness certification [11], and no-cloning quantum teleportation [12, 13].
+In the past few decades, after a series of rigorous mathematical characterizations of quantum nonlocality [14, 15, 16, 17], a great deal of progresses have been made experimentally to prepare entangled or steerable states of microscopic and macroscopic particles, involving platforms based on photons [18], ions [19], atoms [20], superconducting circuits [21], and cavity optomechanical (COM) devices [22, 23, 24].
 However, in terms of generation and manipulation of macroscopic entanglement, it is still challenging to avoid the decoherence effect induced by device imperfection.
-Very recently, to overcome this obstacle and achieve entangled state with high fidelity, a large number of theoretical proposals have been raised, which relies on synthetic gauge field Lai et al.
-(2022a); Liu et al.
-(2023), reservoir engineering Wang and Clerk (2013); Yang et al.
-(2015); Jiao et al.
-, dark-mode or feedback control Lai et al.
-(2022b); Huang et al.
-(2022a); Li et al.
-(2017), photon counting Ho et al.
-(2018), dynamical modulation Wang et al.
-(2016b); Yang et al.
-(2024), and optical nonreciprocity Jiao et al.
-(2020, 2022); Huang et al.
-(2024).
-In parallel, nonreciprocal optical devices, such as optical diodes and circulators, have attracted intense interests due to their potential applications in the next-generation photonic information processing Sounas and Alù (2017).
-The conventional approach for achieving nonreciprocal transmission of light mainly relies on magneto-optical effect Shoji and Mizumoto (2014), which requires magnetic-based materials that are bulk and quite lossy at optical frequency.
-In recent years, by breaking the time-reversal symmetry Caloz et al.
-(2018) via spatial-temporal modulation, optical nonlinearity or Sagnac effect, several magnetic-free scheme have been theoretically proposed and experimentally demonstrated based on atomic ensembles Zhang et al.
-(2018); Yang et al.
-(2019); Liang et al.
-(2020), optical Dong et al.
-(2015); Xia et al.
-(2018); Rosario Hamann et al.
-(2018) or COM Manipatruni et al.
-(2009); Shen et al.
-(2016); Bernier et al.
-(2017) devices, synthetic structures Peng et al.
-(2014); Chang et al.
-(2014), moving medium Wang et al.
-(2013); Lü et al.
-(2017); Maayani et al.
-(2018), and squeezing injectionTang et al.
-(2022).
-With these advances, the peculiar feature of nonreciprocal optical devices, allowing the flow of light from one side but block it from the other, has recently been employed to realize the unidirectional manipulation of light-matter interaction, which enables a variety of classical nonreciprocal phenomenon, such as one-way optical chaos Zhang et al.
-(2021) or solitons Li et al.
-(2021), unidirectional phonon Jiang et al.
-(2018) or magnon Huang et al.
-(2022b) lasing, and nonreciprocal enhancement optomechanical sidebands Li et al.
-(2020), to name a few.
-Very recently, this functionality has also been extended into the quantum regime to achieve nonreciprocal control of diverse nonclassical effect, such as photon Huang et al.
-(2018); Yang et al.
-(2023); Wang et al.
-(2023) or magnon Wang et al.
-(2022) blockade, microscopic Orr et al.
-(2023); Shen et al.
-(2023); Bin et al.
-(2024) or macroscopic Jiao et al.
-(2020, 2022); Huang et al.
-(2024) entanglement, quantum phase transition Fruchart et al.
-(2021); Zhu et al.
-(2024), and EPR steering Guan et al.
-(2024); Wu et al.
-(2024); Zhong et al.
-(2023).
-Inspired by these studies, we investigate how to achieve nonreciprocal entanglement and asymmetric EPR steering in a hybrid COM system based on whispering-gallery-mode (WGM) resonator.
-We show that by applying a two-photon parametric driving field from one input direction while not applying it from the other, the two degenerate counterpropagating optical modes of the WGM resonator can be unidirectionally squeezed, leading to the time-reversal symmetry breaking of the system.
-Based on this chiral squeezing characteristics, we show that various types of bipartite entanglement of the subsystems and the tripartite entanglement of the whole system can be generated only when coherent light input from the squeezing injection direction, which indicates the achievement of nonreciprocal entanglement.
-More interestingly, by properly adjusting the squeezing parameters, we further show that the directionality of EPR steering can also be stepwise driven from no-way regime, one-way regime to two-way regime in an asymmetric way.
-These results, opening up a promising way to prepare entangled quantum resources with nonreciprocal and asymmetric features, are useful for a variety of secure quantum information protocols Branciard et al.
-(2012); Walk et al.
-(2016); Li et al.
-(2024); He et al.
-(2015b); Chiu et al.
-(2016).
+Very recently, to overcome this obstacle and achieve entangled state with high fidelity, a large number of theoretical proposals have been raised, which relies on synthetic gauge field [25, 26], reservoir engineering [27, 28], dark-mode or feedback control [29, 30, 31], photon counting [32], dynamical modulation [33, 34], injection of quantum squeezing [35, 36], and optical nonreciprocity [37, 38, 39].
+On the other hand, squeezed light, characterized by reduced quantum noise in selected quadratures, has emerged as a vital resource in modern quantum science [40].
+Squeezed light has been widely utilized to cool mechanical motion [41], enhance light-matter interaction [42, 43, 44], and advance quantum information processing [45, 46] across a broad range of quantum platforms, including optomechanical devices [47], cavity QED systems [48], and trapped ions [49].
+In particular, the peculiar feature of quadrature noise reduction make squeezed light an indispensable resource for improving the sensitivity of quantum measurements, leading to practical applications in gravitational-wave detection [50], nanoscale sensing [51], and nondemolition qubit readout [52].
+In practice, squeezed light is typically introduced to quantum systems either by external injection or by internal generation via $\chi^{(2)}$ nonlinearity.
+However, for current experimental techniques, both approaches face significant limitations: external injection suffers from inevitable transmission and coupling losses that render squeezed states fragile, while intracavity generation is fundamentally constrained by the 3 dB limit [53], which prevents quantum noise of a cavity field from being reduced below half of the zero-point fluctuations.
+Recently, Lü et al.
+proposed a novel squeezed-light scheme that combines external injection and intracavity generation [54], effectively addressing the limitations of the conventional methods and enabling single-photon quantum processes in the weak-coupling regime of an optomechanical system.
+Building on this approach, subsequent studies have further explored the potential of this hybrid squeezing strategy in various quantum applications, such as surpassing the 3 dB limit of intracavity squeezing [55], enhancing qubit readout [56], and generating long-lived cat states [57, 58].
+Inspired by these studies, we here investigate how to achieve selective generation of tripartite entanglement and asymmetric EPR steering through tuning squeezing phase in a WGM optomechanical resonator.
+The WGM resonator is made of $\chi^{(2)}$ nonlinear materials and supports two degenerate counterpropagating optical modes and two mechanical modes.
+Particularly, we show that by pumping the $\chi^{(2)}$ nonlinear medium with two-photon optical fields and broadband squeezed lights, a pair of counterpropagating squeezed optical modes could be introduced to the WGM resonator, each coupled with an independent squeezed vacuum reservoir.
+This configuration enables a squeezing-phase-controlled light-reservoir interaction for each squeezed optical mode and provides a flexible tool for manipulating the flow of optical input noises,which plays a key role in producing phase-selective entanglement.
+Particularly, it is found that when tailoring asymmetric flow of optical noises in different input directions, the generation of various types of bipartite and tripartite entanglement become phase-dependent and they can only be generated in an asymmetric way.
+More interestingly, by properly adjusting the squeezing phase-matched condition, we further show that the overall asymmetry of EPR steering can be stepwise driven from no-way regime, one-way regime to two-way regime.
+These generated asymmetric entangled states, with their rich variety, are applicable to a wide range of on-chip secure quantum information protocols [9, 10, 11, 12, 13].
+Besides, our work, providing an efficient all-optical approach to engineer quantum states in an asymmetric way, can also be extended to explore the manipulation of other types of quantum effect, such as photon blockade [59, 60, 61] and quantum phase transition [62, 63].
+As such, we believe that our proposal would not only be promising for serving in fundamental tests of quantum theories, but also can serve as key quantum resources for nascent quantum technologies like quantum secure communication [12, 13] and one-way quantum computing [64, 65].
+Figure 1: (a) Schematic of a squeezed COM setup, showing a tapered-fibers-interfaced microdisk WGM resonator dispersively coupled to a string of nanomechanical oscillators.
+The WGM resonator is made of $\chi^{(2)}$ nonlinear materials, supporting two degenerate counterpropagating modes.
+The mechanical string offers two mechanical modes, which can be excited by placing the string in the evanescent field of the WGM resonator.
+When applying two-photon pumping fields along with broadband squeezed lights from the right fiber ports, two squeezed counterpropagating modes $\hat{a}_{s,\circlearrowright}$ and $\hat{a}_{s,\circlearrowleft}$ could be introduced in the WGM resonator, which are coupled with independent squeezed vacuum reservoirs.
+The coherent driving field is input from the left fiber port and is coupled into or out of the WGM resonator via evanescent coupling.
+(b) Density plot of the enhancement factor $\Pi$ as a function of squeezing strength $r_{d}$ and squeezing reference angle $\theta_{d}$.
+The effective COM interaction can be either enhanced ($\Pi>1$) or reduced ($\Pi<1$) depending on squeezing parameters.
+(c) Plot of the effective optical thermal noise $N_{s,i}$ and the effective two-photon correlation strength $M_{s,i}$ versus the phase difference $\delta\theta_{i}$ for different values of $\delta r_{i}$.
+Specifically, for the CCW input case, we consider the squeezed optical mode $\hat{a}_{s,\circlearrowleft}$ is phase-matched with its squeezed vacuum reservoir, leading to $N_{s,\circlearrowleft}=M_{s,\circlearrowleft}=0$.
+For the CW input case, it is phase-mismatched for $\hat{a}_{s,\circlearrowright}$, leading to $N_{s,\circlearrowright}\neq 0$ and $M_{s,\circlearrowright}\neq 0$.
+(d) The steady-state CM $V$ for opposite input cases.
+The parameters used here are given in the main text.
 This paper is structured as follows.
@@
 In Sec.
-III, based on the numerical simulations, we analyze the method and mechanism for achieving nonreciprocal generation and manipulation of various types of entanglement and EPR steering.
+III, based on the numerical simulations, we analyze the method and mechanism for achieving phase-selective generation and manipulation of various types of entanglement and EPR steering.
 In Sec.
@@
 The detailed derivation process of the effective Hamiltonian, the associated master equation, and the system dynamics is given in the Appendix.
-Figure 1: Schematic of the proposed COM setup consisting of a bottle WGM resonator, an optical waveguide, and a tapered glass-fiber nanospike.
-The WGM resonator supports two degenerate counterpropagating optical modes, i.e., CW and CCW modes.
-The optical driving field is input from the waveguide, which can be coupled into and out of the WGM resonator via evanescent coupling.
-When placing a nanospike close to the WGM resonator, it excites two mechanical modes due to the COM interaction induced by the WGM evanescent field, whose vibrating directions are parallel and orthogonal to the resonator surface, respectively.
-To achieve nonreciprocal manipulation of quantum entanglement and EPR steering, we consider two different scenarios for the the CW and CCW input cases.
-For the CW input case, only a coherent driving field is applied, while for the CCW input case, an additional two-photon parametric driving field with a phase-matched squeezed vacuum reservoir is applied.
 II Theoretical model and system dynamics
-The proposed setup consists of a bottle WGM resonator, an optical waveguide, and a tapered glass-fiber nanospike, which is depicted in Fig.
-1.
-Due to the spatial symmetry of the bottle WGM resonator, it can support two degenerate counterpropagating optical modes with resonance frequency $\omega_{c}$, i.e., the clockwise (CW) and counterclockwise (CCW) modes, respectively.
-The nanospike is usually fabricated by scanning an oxybutane flame along the length of a glass fiber while gently pulling it.
-As demonstrated in the recent experiments Xie et al.
-(2016, 2019); Pennetta et al.
-(2020), by mounting such nanospike on a stainless-steel holder and placing it close to the bottle WGM resonator, it offers two mechanical modes due to the WGM evanescent field induced COM interaction.
-These two mechanical modes with frequency $\omega_{m,j}$ $(j=1,2)$ are vibrating in the directions that are orthogonal and parallel to the bottle surface, and they are coupled linearly due to the optically mediated hybridization effect Shkarin et al.
-(2014); Lin et al.
-(2010).
-The WGM resonator can be driven from both sides, where the optical driving fields from the waveguide are coupled into and out of the WGM resonator via evanescent coupling.
-In addition, we consider two different scenarios for the driving fields applied from the CW and CCW input directions, respectively.
-For the CCW input case, a coherent driving field with frequency $\omega_{d}/2$ and a two-photon parametric driving field with frequency $\omega_{d}$ are simultaneously applied to the waveguide.
-In this situation, in a rotating reference frame with respect to the driving frequency $\omega_{d}/2$, the Hamiltonian of the whole system can be expressed as (setting $\hbar=1$)
-$\hat{H}_{\circlearrowleft}=$ $\hat{H}_{\textrm{sys}}+\hat{H}_{\textrm{dr},\circlearrowleft},$
-$\hat{H}_{\textrm{sys}}=$ $\Delta_{c}\hat{a}_{\circlearrowleft}^{\dagger}\hat{a}_{\circlearrowleft}+\sum_{j=1,2}\left[\dfrac{\omega_{m,j}}{2}(\hat{p}_{j}^{2}+\hat{q}_{j}^{2})-g_{j}\hat{a}_{\circlearrowleft}^{\dagger}\hat{a}_{\circlearrowleft}\hat{q}_{j}\right]$
-$+\chi\hat{q}_{1}\hat{q}_{2},$
-$\hat{H}_{\textrm{dr},\circlearrowleft}=$ $\Xi_{d}(e^{-i\theta_{d}}\hat{a}_{\circlearrowleft}^{\dagger 2}+e^{i\theta_{d}}\hat{a}_{\circlearrowleft}^{2})+i\varepsilon_{d,\circlearrowleft}(\hat{a}_{\circlearrowleft}^{\dagger}-\hat{a}_{\circlearrowleft}),$ (1)
-where $\hat{a}_{\circlearrowleft}$ ($\hat{a}_{\circlearrowleft}^{\dagger}$) is the annihilation (creation) operator of the CCW mode, while $\hat{q}_{j}$ and $\hat{p}_{j}$ are the dimensionless displacement and momentum operators of the $j$th mechanical mode, respectively.
-$\Delta_{c}=\omega_{c}-\omega_{d}/2$ denotes the optical detuning, and $\omega_{c}$ is the resonance frequency of the CCW (CW) mode.
-$g_{j}=(\omega_{c}/R)\sqrt{\hbar/m_{j}\omega_{m,j}}$ is the single-photon COM coupling strength, with $R$ the bottle radius and $m_{j}$ the effective mass of the $j$th mechanical mode.
-$\chi$ is the mechanical coupling strength.
-$\Xi_{d}$ and $\theta_{d}$ are the amplitude and phase of the two-photon parametric driving field.
-The amplitude of the coherent driving field is given by $|\varepsilon_{d,\circlearrowleft}|=\sqrt{2\kappa P_{d}/\hbar\omega_{d}}$, where $P_{d}$ is the input laser power and $\kappa$ is the optical decay rate.
-After performing the Bogoliubov transformation with a unitary operator, $S(\eta_{d})=\exp[(-\eta_{d}\hat{a}_{\circlearrowleft}^{\dagger 2}+\eta_{d}^{\ast}\hat{a}_{\circlearrowleft}^{2})/2]$, a squeezed optical mode $\hat{a}_{s,\circlearrowleft}$ can be introduced, i.e.,
-$S^{\dagger}(\eta_{d})\hat{a}_{\circlearrowleft}S(\eta_{d})=\cosh(r_{d})\hat{a}_{s,\circlearrowleft}-e^{-i\theta_{d}}\sinh(r_{d})\hat{a}_{s,\circlearrowleft}^{\dagger},$ (2)
-where $\eta_{d}=r_{d}e^{-i\theta_{d}}$ is the complex squeezing parameter, with a squeezing strength $r_{d}$ and a squeezing reference angle $\theta_{d}$.
-Hence, by dropping the constant terms, the effective Hamiltonian of the total system in the squeezing picture is derived as [see Appendix A for more details]
+In this paper, we propose how to use phase-controlled quantum noise flows to selectively generate and manipulate quantum entanglement and asymmetric EPR steering in a three-mode COM system with two mechanical modes.
+Here, as depicted in Fig.
+1(a), we consider a near-field COM system as a possible experimental setup [66, 67], which consists of a tapered-fibers-interfaced microdisk WGM resonator dispersively coupled to a string of nanomechanical oscillators.
+The microdisk WGM resonator is assumed to be made up of $\chi^{(2)}$ nonlinear materials, e.g., lithium niobate or aluminum nitride, thereby supporting optical parametric down-conversion process.
+Due to the spatial symmetry of the microdisk WGM resonator, it supports two degenerate counterpropagating optical modes with resonance frequency $\omega_{c}$, i.e., the clockwise (CW) and counterclockwise (CCW) modes, respectively.
+For high-quality WGM resonators without backscattering losses, the CW and CCW modes could keep uncoupled.
+The two mechanical modes with fundamental frequency $\omega_{m,j}$ $(j=1,2)$ are offered by the nanomechanical oscillators in the mechanical string.
+As demonstrated in the experiment [66], by placing such mechanical string close to the tightly confined evanescent field of the microdisk WGM resonator, each oscillator can excite a mechanical vibration mode in the direction orthogonal to the microdisk surface, which is resulting from the evanescent field induced COM interaction.
+Furthermore, due to the high-finesse of optical and mechanical modes, as well as the mitigated losses and thermal effects during the coupling process, the observed COM interaction in this experiment is purely dispersive.
+In addition, given that the two mechanical modes are degenerate and interact with the same cavity mode, an effective phonon-hopping interaction could be generated between them, which is mediated by the cavity field induced hybridization effect [68].
+Note that such optomechanically induced hybridization of distinct mechanical modes has been experimentally demonstrated in nano-optomechanical structures [69], WGM resonators [70, 71], and multimode circuit systems [72].
+As shown in Fig.
+1(a), the microdisk WGM resonator can be driven from two opposite input directions.
+The coherent driving field with frequency $\omega_{d}/2$ applied to port $1$ (port $2$) could be evanescently coupled into the CCW (CW) mode, which is referred to as the CCW (CW) input case.
+For the CCW input case, the Hamiltonian of the whole system in a frame rotating with $\omega_{d}/2$ can be expressed as (setting $\hbar=1$)
+$\hat{H}_{\circlearrowleft}=$ $\hat{H}_{\textrm{c}}+\hat{H}_{\textrm{m}}+\hat{H}_{\textrm{om}}+\hat{H}_{\textrm{dr},\circlearrowleft},$
+$\hat{H}_{\textrm{c}}=$ $\Delta_{c}\hat{a}_{\circlearrowleft}^{\dagger}\hat{a}_{\circlearrowleft}+\Xi_{d}(e^{-i\theta_{d}}\hat{a}_{\circlearrowleft}^{\dagger 2}+e^{i\theta_{d}}\hat{a}_{\circlearrowleft}^{2}),$
+$\hat{H}_{\textrm{m}}=$ $\dfrac{\omega_{m,1}}{2}(\hat{p}_{1}^{2}+\hat{q}_{1}^{2})+\dfrac{\omega_{m,2}}{2}(\hat{p}_{2}^{2}+\hat{q}_{2}^{2})+\lambda\hat{q}_{1}\hat{q}_{2},$
+$\hat{H}_{\textrm{om}}=$ $-g_{1}\hat{a}_{\circlearrowleft}^{\dagger}\hat{a}_{\circlearrowleft}\hat{q}_{1}-g_{2}\hat{a}_{\circlearrowleft}^{\dagger}\hat{a}_{\circlearrowleft}\hat{q}_{2},$
+$\hat{H}_{\textrm{dr},\circlearrowleft}=$ $i\varepsilon_{d,\circlearrowleft}(\hat{a}_{\circlearrowleft}^{\dagger}-\hat{a}_{\circlearrowleft}),$ (1)
+[…]
+In our proposed scheme, the primary distinction between the CW and CCW input cases lies in the squeezing phase and strength of their injected squeezed vacuum reservoirs, which could induce asymmetric quantum noise flows in these counterpropagating modes.
+In the following, we discuss how to independently inject squeezed vacuum reservoirs to CW and CCW modes, and then derive the effective master equation for CCW input case as a specific example.
+As shown in Fig.
+1(a), we consider applying two different broadband squeezed optical fields to these counterpropagating modes, which are centred around frequency $\omega_{c}$ and characterized by squeezing strengths $r_{e,i}$ and reference phase angles $\theta_{e,i}$ ($i=\circlearrowleft,\circlearrowleft$), respectively.
+In practice, these broadband squeezed optical fields can act as two independent squeezed-vacuum reservoirs to CW and CCW modes [74].
+This is because the bandwidth of squeezed optical field can be up to gigahertz (GHz[48], while the typical linewidth of WGM resonators is of the order of megahertz (MHz).
+In this case, the broadband squeezed optical field can be well approximated as having infinite bandwidth to WGM optical modes and be safely treated as a squeezed vacuum reservoir.
+Besides, the two mechanical modes are assumed to be coupled with two independent thermal reservoirs at the same bath temperature $T$.
+Then, in terms of the CCW input case, including the dissipations of the optical and mechanical modes, the dynamics of the total system is governed by the following Born-Markovian master equation [75]
+$\dot{\rho}_{\circlearrowleft}=$ $i\big{[}\rho_{\circlearrowleft},\hat{H}_{\circlearrowleft}\big{]}+\frac{\kappa}{2}(N_{e,\circlearrowleft}+1)\mathcal{D}[\hat{a}_{\circlearrowleft}]\rho_{\circlearrowleft}+\frac{\kappa}{2}N_{e,\circlearrowleft}\mathcal{D}[\hat{a}_{\circlearrowleft}^{\dagger}]\rho_{\circlearrowleft}$
+$-\frac{\kappa}{2}M_{e,\circlearrowleft}\mathcal{G}[\hat{a}_{\circlearrowleft}]\rho_{\circlearrowleft}-\frac{\kappa}{2}M_{e,\circlearrowleft}^{\ast}\mathcal{G}[\hat{a}_{\circlearrowleft}^{\dagger}]\rho_{\circlearrowleft}$
+$-\sum_{j=1,2}\left(i\dfrac{\gamma_{m,j}}{2}[\hat{q}_{j},\{\hat{p}_{j},\rho_{\circlearrowleft}\}]+\gamma_{m,j}\bar{n}_{m,j}[\hat{q}_{j},[\hat{q}_{j},\rho_{\circlearrowleft}]]\right),$ (2)
+where
+$\mathcal{D}[\hat{a}_{\circlearrowleft}]\rho_{\circlearrowleft}$ $=2\hat{a}_{\circlearrowleft}\rho_{\circlearrowleft}\hat{a}_{\circlearrowleft}^{\dagger}-(\hat{a}_{\circlearrowleft}^{\dagger}\hat{a}_{\circlearrowleft}\rho_{\circlearrowleft}+\rho_{\circlearrowleft}\hat{a}_{\circlearrowleft}^{\dagger}\hat{a}_{\circlearrowleft}),$
+$\mathcal{G}[\hat{a}_{\circlearrowleft}]\rho_{\circlearrowleft}$ $=2\hat{a}_{\circlearrowleft}\rho_{\circlearrowleft}\hat{a}_{\circlearrowleft}-(\hat{a}_{\circlearrowleft}\hat{a}_{\circlearrowleft}\rho_{\circlearrowleft}+\rho_{\circlearrowleft}\hat{a}_{\circlearrowleft}\hat{a}_{\circlearrowleft})$ (3)
+[…]
+After performing the Bogoliubov transformation with a unitary operator, $U_{\circlearrowleft}(\eta_{d})=\exp[(-\eta_{d}\hat{a}_{\circlearrowleft}^{\dagger 2}+\eta_{d}^{\ast}\hat{a}_{\circlearrowleft}^{2})/2]$, a squeezed optical mode $\hat{a}_{s,\circlearrowleft}$ can be introduced, i.e.,
+$U_{\circlearrowleft}^{\dagger}(\eta_{d})\hat{a}_{\circlearrowleft}U_{\circlearrowleft}(\eta_{d})=\cosh(r_{d})\hat{a}_{s,\circlearrowleft}-e^{-i\theta_{d}}\sinh(r_{d})\hat{a}_{s,\circlearrowleft}^{\dagger},$ (4)
+where $\eta_{d}=r_{d}e^{-i\theta_{d}}$ is the complex squeezing parameter, with a squeezing strength $r_{d}=(1/4)\ln[(\Delta_{c}+2\Xi_{d})/(\Delta_{c}-2\Xi_{d})]$ and a squeezing reference angle $\theta_{d}$.
+Hence, by dropping the constant terms, the effective Hamiltonian with respect to the CCW input case in the squeezing picture is derived as [see Appendix A for more details]
 $\hat{\tilde{H}}_{\circlearrowleft}=$ $\omega_{s}\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{a}_{s,\circlearrowleft}+\sum_{j=1,2}\left[\dfrac{\omega_{m,j}}{2}(\hat{p}_{j}^{2}+\hat{q}_{j}^{2})-\zeta_{s,j}\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{a}_{s,\circlearrowleft}\hat{q}_{j}\right.$
-$\left.+\dfrac{\zeta_{p,j}}{2}(e^{-i\theta_{d}}\hat{a}_{s,\circlearrowleft}^{\dagger 2}+e^{i\theta_{d}}\hat{a}_{s,\circlearrowleft}^{2})\hat{q}_{j}-F_{j}\hat{q}_{j}\right]+\chi\hat{q}_{1}\hat{q}_{2}$
+$\left.+\dfrac{\zeta_{p,j}}{2}(e^{-i\theta_{d}}\hat{a}_{s,\circlearrowleft}^{\dagger 2}+e^{i\theta_{d}}\hat{a}_{s,\circlearrowleft}^{2})\hat{q}_{j}-F_{j}\hat{q}_{j}\right]+\lambda\hat{q}_{1}\hat{q}_{2}$
 $+i\varepsilon_{d,\circlearrowleft}\sinh(r_{d})(e^{-i\theta_{d}}\hat{a}_{s,\circlearrowleft}^{\dagger}-e^{i\theta_{d}}\hat{a}_{s,\circlearrowleft})$
@@
 $F_{j}=g_{j}\sinh^{2}(r_{d})$ denotes the strength of the constant mechanical driving force exerted on the $j$th mechanical mode, which is caused by the parametric amplification.
-As discussed in the preceding studies Jiao et al.
-; Tang et al.
-(2022); Guan et al.
-(2024); Wu et al.
-(2024); Lü et al.
-(2015), by adjusting the squeezing parameters with respect to the two-photon parametric driving field, the effective light-matter interaction can be manipulated and enhanced in a controllable way.
-However, it is worth emphasizing that when the squeezed optical mode is coupled to a normal vacuum reservoir, the optical dissipation will be inevitably amplified in the meantime, thus leading to its quantum decoherence.
-To eliminate such optical dissipation induced by the intracavity squeezing, one can apply a broadband squeezed-vacuum reservoir that is phase-matched with the squeezed optical mode, i.e., the squeezing strength $r_{e}$ and phase $\theta_{e}$ of this reservoir satisfying the relationships $r_{e}-r_{d}=0$ and $\theta_{e}-\theta_{d}=\pm n\pi~{}(n=1,3,5,\ldots)$.
-Note that the injection of this required broadband squeezed-vacuum reservoir has been reported through using an optical parametric amplifier Kashiwazaki et al.
-(2020).
-Additionally, given that the two mechanical modes are coupled with two independent thermal reservoirs at same bath temperature $T$, the Born-Markovian master equation describing the dynamics of the whole system in the squeezing picture can be derived as [see Appendix B for more details]
-$\dot{\tilde{\rho}}_{\circlearrowleft}\!=$ $i\!\left[\tilde{\rho}_{\circlearrowleft},\hat{\tilde{H}}_{\circlearrowleft}\right]\!+\!\frac{\kappa}{2}\mathcal{D}[\hat{a}_{s,\circlearrowleft}]\tilde{\rho}_{\circlearrowleft}\!-\!\sum_{j=1,2}\left(i\dfrac{\gamma_{m,j}}{2}[\hat{q}_{j},\{\hat{p}_{j},\tilde{\rho}_{\circlearrowleft}\}]\right.$
-$\left.+\gamma_{m,j}\bar{n}_{m,j}[\hat{q}_{j},[\hat{q}_{j},\tilde{\rho}_{\circlearrowleft}]]\right),$ (4)
+Correspondingly, the effective master equation in the squeezing picture is derived as [see Appendix B for more details]
+$\dot{\tilde{\rho}}_{\circlearrowleft}=$ $U^{\dagger}(\eta_{d})\dot{\rho}_{\circlearrowleft}U(\eta_{d})$
+$=$ $i\big{[}\tilde{\rho}_{\circlearrowleft},\hat{\tilde{H}}_{\circlearrowleft}\big{]}\!+\!\frac{\kappa}{2}(N_{s,\circlearrowleft}+1)\mathcal{D}[\hat{a}_{s,\circlearrowleft}]\tilde{\rho}_{\circlearrowleft}\!+\!\frac{\kappa}{2}N_{s,\circlearrowleft}\mathcal{D}[\hat{a}_{s,\circlearrowleft}^{\dagger}]\tilde{\rho}_{\circlearrowleft}$
+$-\frac{\kappa}{2}M_{s,\circlearrowleft}\mathcal{G}[\hat{a}_{s,\circlearrowleft}]\tilde{\rho}_{\circlearrowleft}-\frac{\kappa}{2}M_{s,\circlearrowleft}^{\ast}\mathcal{G}[\hat{a}_{s,\circlearrowleft}]\tilde{\rho}_{\circlearrowleft}$
+$-\sum_{j=1,2}\left(i\dfrac{\gamma_{m,j}}{2}[\hat{q}_{j},\{\hat{p}_{j},\tilde{\rho}_{\circlearrowleft}\}]+\gamma_{m,j}\bar{n}_{m,j}[\hat{q}_{j},[\hat{q}_{j},\tilde{\rho}_{\circlearrowleft}]]\right),$ (6)
 where
-$\mathcal{D}[\hat{a}_{s,\circlearrowleft}]\tilde{\rho}_{\circlearrowleft}=2\hat{a}_{s,\circlearrowleft}\tilde{\rho}_{\circlearrowleft}\hat{a}_{s,\circlearrowleft}^{\dagger}-(\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{a}_{s,\circlearrowleft}\tilde{\rho}_{\circlearrowleft}+\tilde{\rho}_{\circlearrowleft}\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{a}_{s,\circlearrowleft})$ (5)
+$N_{s,\circlearrowleft}=$ $\sinh^{2}(r_{d})\cosh^{2}(r_{e,\circlearrowleft})+\cosh^{2}(r_{d})\sinh^{2}(r_{e,\circlearrowleft})$
+$+\dfrac{1}{2}\cos\delta\theta\sinh(2r_{d})\sinh(2r_{e,\circlearrowleft}),$
+$M_{s,\circlearrowleft}\!=$ $e^{i\theta_{d}}\![\sinh(r_{d})\!\cosh(r_{e,\circlearrowleft})\!+\!e^{i\delta\theta_{\circlearrowleft}}\!\cosh(r_{d})\!\sinh(r_{e,\circlearrowleft})\!]$
+×$\times\![\cosh(r_{d})\!\cosh(r_{e,\circlearrowleft})\!+\!e^{-i\delta\theta_{\circlearrowleft}}\!\sinh(r_{d})\!\sinh(r_{e,\circlearrowleft})],$ (7)
 […]
-For the CW input case, there is merely a coherent driving field with frequency $\omega_{d}/2$ applied to the waveguide. In this situation, the Hamiltonian of the whole system reads
-$\hat{H}_{\circlearrowright}=$ $\hat{H}_{\textrm{sys}}+\hat{H}_{\textrm{dr},\circlearrowright},$
-$\hat{H}_{\textrm{sys}}=$ $\Delta_{c}\hat{a}_{\circlearrowright}^{\dagger}\hat{a}_{\circlearrowright}+\sum_{j=1,2}\left[\dfrac{\omega_{m,j}}{2}(\hat{p}_{j}^{2}+\hat{q}_{j}^{2})-g_{j}\hat{a}_{\circlearrowright}^{\dagger}\hat{a}_{\circlearrowright}\hat{q}_{j}\right]$
-$+\chi\hat{q}_{1}\hat{q}_{2},$
-$\hat{H}_{\textrm{dr},\circlearrowright}=$ $i\varepsilon_{d,\circlearrowright}(\hat{a}_{\circlearrowright}^{\dagger}-\hat{a}_{\circlearrowright}),$ (6)
-and the associated Born-Markovian master equation is given by Gardiner and Zoller (2000)
-$\dot{\rho}_{\circlearrowright}\!=$ $i\!\left[\rho_{\circlearrowright},\hat{H}_{\circlearrowright}\right]\!+\!\frac{\kappa}{2}\mathcal{D}[\hat{a}_{\circlearrowright}]\rho_{\circlearrowright}\!-\!\sum_{j=1,2}\left(i\dfrac{\gamma_{m,j}}{2}[\hat{q}_{j},\{\hat{p}_{j},\rho_{\circlearrowright}\}]\right.$
-$\left.+\gamma_{m,j}\bar{n}_{m,j}[\hat{q}_{j},[\hat{q}_{j},\rho_{\circlearrowright}]]\right),$ (7)
-where $\hat{a}_{\circlearrowright}$ ($\hat{a}_{\circlearrowright}^{\dagger}$) is the annihilation (creation) operator of the CW mode, and $\rho_{\circlearrowright}$ is the density operator with respect to the CW input case.
-Note that in the absence of the two-photon parametric driving field, the Bogoliubov transformation has no influence on the system dynamics.
-Therefore, we apply a normal vacuum reservoir to the CW mode and analyze the dynamical evolution of this system in the original picture.
-[…]
-For the CCW input case, due to the nonlinear COM interactions in Hamiltonian (3), the system dynamics is difficult to be directly solved.
-In order to deal with this problem, one can linearize the Hamiltonian by expanding each operator as a sum of its steady-state mean value and a small quantum fluctuation around it under the condition of strong coherent optical driving, i.e., $\hat{a}_{s,\circlearrowleft}=\bar{a}_{s,\circlearrowleft}+\delta\hat{a}_{s,\circlearrowleft}$, $\hat{q}_{j}=\bar{q}_{j}+\delta\hat{q}_{j}$, $\hat{p}_{j}=\bar{p}_{j}+\delta\hat{p}_{j}$.
-The steady-state mean values of the optical and mechanical modes can be derived by using the master equation (4), i.e.,
-$\bar{a}_{s,\circlearrowright}=$ $-\dfrac{(i\Delta_{s}-\dfrac{\kappa}{2})A_{1}+iA_{2}\beta_{p}}{(\Delta_{s}^{2}+\dfrac{\kappa^{2}}{4})-\beta_{p}^{2}}\varepsilon_{d,\circlearrowleft},$
-$\bar{q}_{1}=$ $\dfrac{\omega_{m,2}B_{1}-\chi B_{2}}{\omega_{m,1}\omega_{m,2}-\chi^{2}},~{}~{}\bar{p}_{1}=0,$
-$\bar{q}_{2}=$ $\dfrac{\omega_{m,1}B_{2}-\chi B_{1}}{\omega_{m,1}\omega_{m,2}-\chi^{2}},~{}~{}\bar{p}_{2}=0,$ (8)
+To quantitatively investigate the behavior of quantum entanglement and EPR steering, one should solve the system dynamics by employing the effective mater equation (6).
+However, due to the nonlinear COM interactions in Hamiltonian (5), the system dynamics is difficult to be directly solved.
+In order to deal with this problem, one can linearize the Hamiltonian by expanding each operator as a sum of its steady-state mean value and a small quantum fluctuation around it under the condition of strong coherent optical driving, i.e., $\hat{a}_{s,i}=\bar{a}_{s,i}+\delta\hat{a}_{s,i}$, $\hat{q}_{j}=\bar{q}_{j}+\delta\hat{q}_{j}$, $\hat{p}_{j}=\bar{p}_{j}+\delta\hat{p}_{j}$.
+Since the linearization procedure is identical for the CW and CCW input cases, we here only present the linearization of the effective mater equation (6) regarding the CCW input case.
+By using the master equation (6), the steady-state mean values of the optical and mechanical modes with respect to the CCW input case are derived as
+$\bar{a}_{s,\circlearrowleft}=$ $-\dfrac{(i\Delta_{s}-\dfrac{\kappa}{2})A_{1}+iA_{2}\beta_{p}}{(\Delta_{s}^{2}+\dfrac{\kappa^{2}}{4})-\beta_{p}^{2}}\varepsilon_{d,\circlearrowleft},$
+$\bar{q}_{1}=$ $\dfrac{\omega_{m,2}B_{1,\circlearrowleft}-\lambda B_{2,\circlearrowleft}}{\omega_{m,1}\omega_{m,2}-\lambda^{2}},~{}~{}\bar{p}_{1}=0,$
+$\bar{q}_{2}=$ $\dfrac{\omega_{m,1}B_{2,\circlearrowleft}-\lambda B_{1,\circlearrowleft}}{\omega_{m,1}\omega_{m,2}-\lambda^{2}},~{}~{}\bar{p}_{2}=0,$ (8)
 with
-$\alpha_{s}$ $=e^{-i\theta_{d}}\bar{a}_{s,\circlearrowleft}^{\ast 2}+e^{i\theta_{d}}\bar{a}_{s,\circlearrowleft}^{2},$
+$\Delta_{s}$ $=\omega_{s}-\beta_{s},$
+$\alpha_{s,\circlearrowleft}$ $=e^{-i\theta_{d}}\bar{a}_{s,\circlearrowleft}^{\ast 2}+e^{i\theta_{d}}\bar{a}_{s,\circlearrowleft}^{2},$
 $\beta_{s}$ $=\zeta_{s,1}\bar{q}_{1}+\zeta_{s,2}\bar{q}_{2},$
 $\beta_{p}$ $=\zeta_{p,1}\bar{q}_{1}+\zeta_{p,2}\bar{q}_{2},$
-$\Delta_{s}$ $=\omega_{s}-\beta_{s},$
 $A_{1}$ $=\cosh(r_{d})+\sinh(r_{d})e^{-i\theta_{d}},$
 $A_{2}$ $=\cosh(r_{d})e^{-i\theta_{d}}+\sinh(r_{d}),$
-$B_{1}$ $=\zeta_{s,1}|\bar{a}_{s,\circlearrowleft}|^{2}-\dfrac{\zeta_{p,1}}{2}\alpha_{s}+F_{1},$
-$B_{2}$ $=\zeta_{s,2}|\bar{a}_{s,\circlearrowleft}|^{2}-\dfrac{\zeta_{p,2}}{2}\alpha_{s}+F_{2}.$ (9)
+$B_{1,\circlearrowleft}$ $=\zeta_{s,1}|\bar{a}_{s,\circlearrowleft}|^{2}-\dfrac{\zeta_{p,1}}{2}\alpha_{s,\circlearrowleft}+F_{1},$
+$B_{2,\circlearrowleft}$ $=\zeta_{s,2}|\bar{a}_{s,\circlearrowleft}|^{2}-\dfrac{\zeta_{p,2}}{2}\alpha_{s,\circlearrowleft}+F_{2}.$ (9)
 Equation (8) indicates that the field amplitude $\bar{a}_{s,\circlearrowleft}$ is not only dependent on the strength of the coherent driving field, but also relies on the squeezing strength and reference angle of the squeezed optical mode.
-This allows us to regulate the COM coupling strength and break the optical reciprocity of the system by adjusting the squeezing parameters.
-Moreover, it is also seen that the mean values of the mechanical displacement $\bar{q}_{j}$ are coupled to each other by a factor of $\chi$, which is due to the phonon hopping process between the two mechanical modes.
+This allows us to regulate the COM coupling strength in a controllable way by adjusting the squeezing parameters.
+Moreover, it is also seen that the mean values of the mechanical displacement $\bar{q}_{j}$ are coupled to each other by a factor of $\lambda$, which is due to the phonon hopping process between the two mechanical modes.
 Then, by substituting the expansions of the quantum operators into the Hamiltonian (5), one can directly derive the linearized Hamiltonian as
 $\hat{\tilde{H}}_{lin,\circlearrowleft}=$ $\Delta_{s}\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{a}_{s,\circlearrowleft}+\sum_{j=1,2}\left[\dfrac{\omega_{m,j}}{2}(\hat{p}_{j}^{2}+\hat{q}_{j}^{2})\right.$
-$\left.-(\Lambda_{j}\hat{a}_{s,\circlearrowleft}^{\dagger}+\Lambda_{j}^{\ast}\hat{a}_{s,\circlearrowleft})\hat{q}_{j}\right]+\chi\hat{q}_{1}\hat{q}_{2},$ (10)
+$\left.-(\Lambda_{j}\hat{a}_{s,\circlearrowleft}^{\dagger}+\Lambda_{j}^{\ast}\hat{a}_{s,\circlearrowleft})\hat{q}_{j}\right]+\lambda\hat{q}_{1}\hat{q}_{2},$ (10)
 and the associated master equation as
-$\dot{\tilde{\rho}}_{\circlearrowleft}\!=$ $i\!\left[\tilde{\rho}_{\circlearrowleft},\tilde{H}_{lin,\circlearrowleft}\right]\!+\!\frac{\kappa}{2}\mathcal{D}[\hat{a}_{s,\circlearrowleft}]\tilde{\rho}_{\circlearrowleft}\!-\!\sum_{j=1,2}\!\left(i\dfrac{\gamma_{m,j}}{2}[\hat{q}_{j},\{\hat{p}_{j},\tilde{\rho}_{\circlearrowleft}\}]\right.$
-$\left.+\gamma_{m,j}\bar{n}_{m,j}[\hat{q}_{j},[\hat{q}_{j},\tilde{\rho}_{\circlearrowleft}]]\right),$ (11)
+$\dot{\tilde{\rho}}_{\circlearrowleft}\!=$ $i\!\left[\tilde{\rho}_{\circlearrowleft},\tilde{H}_{lin,\circlearrowleft}\right]$
+$+\!\frac{\kappa}{2}(N_{s,\circlearrowleft}+1)\mathcal{D}[\hat{a}_{s,\circlearrowleft}]\tilde{\rho}_{\circlearrowleft}\!+\!\frac{\kappa}{2}N_{s,\circlearrowleft}\mathcal{D}[\hat{a}_{s,\circlearrowleft}^{\dagger}]\tilde{\rho}_{\circlearrowleft}$
+$-\frac{\kappa}{2}M_{s,\circlearrowleft}\mathcal{G}[\hat{a}_{s,\circlearrowleft}]\tilde{\rho}_{\circlearrowleft}-\frac{\kappa}{2}M_{s,\circlearrowleft}^{\ast}\mathcal{G}[\hat{a}_{s,\circlearrowleft}]\tilde{\rho}_{\circlearrowleft}$
+$-\sum_{j=1,2}\left(i\dfrac{\gamma_{m,j}}{2}[\hat{q}_{j},\{\hat{p}_{j},\tilde{\rho}_{\circlearrowleft}\}]+\gamma_{m,j}\bar{n}_{m,j}[\hat{q}_{j},[\hat{q}_{j},\tilde{\rho}_{\circlearrowleft}]]\right),$ (11)
 where
@@
 is the effective COM coupling rate, with $G_{j}=g_{j}\bar{a}_{s,\circlearrowleft}$.
-For notational convenience, we have neglected the symbol ``$\delta"$ in the expression of quantum fluctuation operators in Eqs.
+By defining an enhancement factor $\Pi_{j}=|\Lambda_{j}/G_{j}|$, it is seen that with proper squeezing parameters, the COM coupling rate can be effectively enhanced [see Fig.
+1(b)].
+Note that for notational convenience, we have neglected the symbol ``$\delta"$ in the expression of quantum fluctuation operators in Eqs.
 (10) and (11).
@@
 In the following discussions, for ensuring the validity of Hamiltonian (10), the system parameters have been strictly restricted to satisfy the condition of weak COM coupling.
-Figure 2: Squeezing-induced nonreciprocal bipartite and tripartite entanglement.
-The logarithmic negativity $E_{\mathcal{N}}^{\textit{a}|\textit{m}_{\textit{1}}}$ (a), $E_{\mathcal{N}}^{\textit{a}|\textit{m}_{\textit{2}}}$ (b), $E_{\mathcal{N}}^{\textit{m}_{\textit{1}}|\textit{m}_{\textit{2}}}$ (c), and the minimum residual contangle $\mathcal{R}^{\textrm{min}}_{\tau}$ (d) for opposite optical input directions are plotted as functions of the scaled optical detuning $\Delta_{c}/\omega_{m}$.
-The left panels of these figures show the bipartite and tripartite entanglement with respect to two specific squeezing reference angles of $\theta_{d}=0.75\pi$ and $\pi$, respectively, with $r_{d}=0.3$.
-The right panels show in detail the dependence of such entanglement on squeezing reference angle $\theta_{d}$.
+Figure 2: Phase-selective generation of optomechanical entanglement.
+(a),(c) Plot of the logarithmic negativity $E_{\mathcal{N}}^{\hat{\textit{a}}_{\textit{s},\textit{i}}|\hat{\textit{q}}_{\textit{1}}}$ and $E_{\mathcal{N}}^{\hat{\textit{a}}_{\textit{s},\textit{i}}|\hat{\textit{q}}_{\textit{2}}}$ versus the scaled optical detuning $\Delta_{c}/\omega_{m}$ for opposite optical input directions with $r_{d}=0.1$ and $\theta_{d}=\pi$.
+(b),(d) Density plot of $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ as functions of $\delta r_{\circlearrowright}$ and $\delta\theta_{\circlearrowright}$, with $\delta r_{\circlearrowleft}=0$ and $\delta\theta_{\circlearrowleft}=\pi$.
+(e),(f) Plot of the logarithmic negativity $E_{\mathcal{N}}^{\hat{\textit{a}}_{\textit{s},\textit{i}}|\hat{\textit{q}}_{\textit{1}}}$ and $E_{\mathcal{N}}^{\hat{\textit{a}}_{\textit{s},\textit{i}}|\hat{\textit{q}}_{\textit{2}}}$ versus the mechanical environment temperature $T$ for opposite optical input directions with $r_{d}=0.1$, $\theta_{d}=\pi$, and $\Delta_{c}=1$.
 Since the system dynamics is linearized now and the input noises for the optical and mechanical modes are Gaussian, the steady state of the system, independently of any initial state, could eventually evolve into a tripartite zero-mean Gaussian state, whose statistic is fully characterized by a $6\times 6$ covariance matrix (CM) $V$ with its matrix element
 $V_{kl}=\langle\psi_{k}\psi_{l}+\psi_{l}\psi_{k}\rangle/2,~{}(k,l=1,2,\ldots,6).$ (13)
-Here $\psi=(\hat{X},\hat{Y},\hat{q}_{1},\hat{p}_{1},\hat{q}_{2},\hat{p}_{2})^{T}$ is the vector of optical and mechanical quadrature operators, with its components defined by
-$\hat{X}=\dfrac{1}{\sqrt{2}}\left(\hat{a}_{s,\circlearrowleft}^{\dagger}+\hat{a}_{s,\circlearrowleft}\right),~{}~{}~{}\hat{Y}=\dfrac{i}{\sqrt{2}}\left(\hat{a}_{s,\circlearrowleft}^{\dagger}-\hat{a}_{s,\circlearrowleft}\right).$ (14)
-Employing the master equation (11), one can obtain the dynamics of arbitrary quantum correlation between the optical and mechanical quadrature operators in the CM $V$.
+Here $\psi=(\hat{X}_{s,\circlearrowleft},\hat{Y}_{s,\circlearrowleft},\hat{q}_{1},\hat{p}_{1},\hat{q}_{2},\hat{p}_{2})^{T}$ is the vector of optical and mechanical quadrature operators, with its components defined by
+$\hat{X}_{s,\circlearrowleft}=\dfrac{1}{\sqrt{2}}\left(\hat{a}_{s,\circlearrowleft}^{\dagger}+\hat{a}_{s,\circlearrowleft}\right),~{}\hat{Y}_{s,\circlearrowleft}=\dfrac{i}{\sqrt{2}}\left(\hat{a}_{s,\circlearrowleft}^{\dagger}-\hat{a}_{s,\circlearrowleft}\right).$ (14)
+Employing the linearized master equation (11), one can obtain the dynamics of arbitrary quantum correlation between the optical and mechanical quadrature operators in the CM $V$.
 In terms of optical bosonic operator, we define the following quantum correlation functions
-$x_{1}=\langle\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{a}_{s,\circlearrowleft}\rangle,~{}x_{2}=\langle\hat{a}_{s,\circlearrowleft}\hat{a}_{s,\circlearrowleft}^{\dagger}\rangle,~{}x_{3}=\langle\hat{q}_{1}\hat{q}_{1}\rangle,$
-$x_{4}=\langle\hat{p}_{1}\hat{p}_{1}\rangle,~{}x_{5}=\langle\hat{q}_{2}\hat{q}_{2}\rangle,~{}x_{6}=\langle\hat{p}_{2}\hat{p}_{2}\rangle,$
-$x_{7}=\langle\hat{a}_{s,\circlearrowleft}\hat{a}_{s,\circlearrowleft}\rangle,~{}x_{8}=\langle\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{a}_{s,\circlearrowleft}^{\dagger}\rangle,~{}x_{9}=\langle\hat{q}_{1}\hat{p}_{1}\rangle,$
-$x_{10}=\langle\hat{p}_{1}\hat{q}_{1}\rangle,~{}x_{11}=\langle\hat{q}_{2}\hat{p}_{2}\rangle,~{}x_{12}=\langle\hat{p}_{2}\hat{q}_{2}\rangle,$
-$x_{13}=\langle\hat{a}_{s,\circlearrowleft}\hat{q}_{1}\rangle,~{}x_{14}=\langle\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{q}_{1}\rangle,~{}x_{15}=\langle\hat{a}_{s,\circlearrowleft}\hat{p}_{1}\rangle,$
-$x_{16}=\langle\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{p}_{1}\rangle,~{}x_{17}=\langle\hat{a}_{s,\circlearrowleft}\hat{q}_{2}\rangle,~{}x_{18}=\langle\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{q}_{2}\rangle,$
-$x_{19}=\langle\hat{a}_{s,\circlearrowleft}\hat{p}_{2}\rangle,~{}x_{20}=\langle\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{p}_{2}\rangle,~{}x_{21}=\langle\hat{q}_{1}\hat{q}_{2}\rangle,$
-$x_{22}=\langle\hat{p}_{1}\hat{p}_{2}\rangle,~{}x_{23}=\langle\hat{q}_{1}\hat{p}_{2}\rangle,~{}x_{24}=\langle\hat{q}_{2}\hat{p}_{1}\rangle.$ (15)
-By grouping together them in a vector, $X=(x_{1},x_{2},\ldots,x_{24})^{T}$, one can obtain its time evolution equation as
-$\dfrac{d}{dt}X=M\cdot X+N,$ (16)
+$x_{1}=\langle\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{a}_{s,\circlearrowleft}\rangle,~{}~{}x_{2}=\langle\hat{a}_{s,\circlearrowleft}\hat{a}_{s,\circlearrowleft}^{\dagger}\rangle,~{}~{}x_{3}=\langle\hat{q}_{1}\hat{q}_{1}\rangle,$
+$x_{4}=\langle\hat{p}_{1}\hat{p}_{1}\rangle,~{}~{}x_{5}=\langle\hat{q}_{2}\hat{q}_{2}\rangle,~{}~{}x_{6}=\langle\hat{p}_{2}\hat{p}_{2}\rangle,$
+$x_{7}=\langle\hat{a}_{s,\circlearrowleft}\hat{a}_{s,\circlearrowleft}\rangle,~{}~{}x_{8}=\langle\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{a}_{s,\circlearrowleft}^{\dagger}\rangle,~{}~{}x_{9}=\langle\hat{q}_{1}\hat{p}_{1}\rangle,$
+$x_{10}=\langle\hat{p}_{1}\hat{q}_{1}\rangle,~{}~{}x_{11}=\langle\hat{q}_{2}\hat{p}_{2}\rangle,~{}~{}x_{12}=\langle\hat{p}_{2}\hat{q}_{2}\rangle,$
+$x_{13}=\langle\hat{a}_{s,\circlearrowleft}\hat{q}_{1}\rangle,~{}~{}x_{14}=\langle\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{q}_{1}\rangle,~{}~{}x_{15}=\langle\hat{a}_{s,\circlearrowleft}\hat{p}_{1}\rangle,$
+$x_{16}=\langle\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{p}_{1}\rangle,~{}~{}x_{17}=\langle\hat{a}_{s,\circlearrowleft}\hat{q}_{2}\rangle,~{}~{}x_{18}=\langle\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{q}_{2}\rangle,$
+$x_{19}=\langle\hat{a}_{s,\circlearrowleft}\hat{p}_{2}\rangle,~{}~{}x_{20}=\langle\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{p}_{2}\rangle,~{}~{}x_{21}=\langle\hat{q}_{1}\hat{q}_{2}\rangle,$
+$x_{22}=\langle\hat{p}_{1}\hat{p}_{2}\rangle,~{}~{}x_{23}=\langle\hat{q}_{1}\hat{p}_{2}\rangle,~{}~{}x_{24}=\langle\hat{q}_{2}\hat{p}_{1}\rangle.$ (15)
+By grouping together these quantum correlations in a vector $X=(x_{1},x_{2},\ldots,x_{24})^{T}$ and employing the linearized effective master equation (11), one can obtain its time evolution equation as
+$\dfrac{d}{dt}X=A\cdot X+b,$ (16)
 where
-$N=$ $\left[0,\kappa,0,2\gamma_{m,1}\bar{n}_{m,1},0,2\gamma_{m,2}\bar{n}_{m,2},0,0,\right.$
-$\left.0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0\right]^{T},$ (17)
+$b=$ $\left[\kappa N_{s},\kappa(N_{s}+1),0,2\gamma_{m,1}\bar{n}_{m,1},0,2\gamma_{m,2}\bar{n}_{m,2},\kappa M_{s}^{\ast},\right.$
+$\left.\kappa M_{s},0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0\right]^{T},$ (17)
 is a vector involving the correlation functions for the input quantum noises.
-Here the exact expression for the coefficient matrix $M$ and the detailed derivation process of the time evolution equation (16) are too cumbersome, and, for convenience, we have reported them in Appendix C.
+Here the exact expression for the coefficient matrix $A$ and the detailed derivation process of the time evolution equation (16) are too cumbersome, and, for convenience, we have reported them in Appendix C.
 Notably, by numerically solving Eq.
@@
 Regarding the verification of bipartite and tripartite entanglement, we adopt the logarithmic negativity $E_{\mathcal{N}}$ and the minimum residual contangle $\mathcal{R}^{\textrm{min}}_{\tau}$ as the quantitative entanglement measures, respectively, which are defined based on specifying the positivity of the partial transpose of the CM $V$.
-For continuous-variable (CV) bipartite Gaussian state, the logarithmic negativity $E_{\mathcal{N}}^{\mu|\nu}$ is defined as Adesso et al.
-(2004)
+For continuous-variable (CV) bipartite Gaussian state, the logarithmic negativity $E_{\mathcal{N}}^{\mu|\nu}$ is defined as [15]
 $E_{\mathcal{N}}^{\mu|\nu}=$ $\max\,\big{[}0,-\ln(2\eta_{0}^{-})\big{]},$ (18)
-where $\eta_{0}^{-}\!\equiv\!2^{-1/2}\{\Sigma(V_{\mu\nu})-[\Sigma(V_{\mu\nu})^{2}-4\det\!V_{\mu\nu}]^{1/2}\}^{1/2}$ , with $\Sigma(V_{\mu\nu})\!\equiv\!\det\mathcal{A}_{\mu}+\det\mathcal{B}_{\nu}-2\det\mathcal{C}_{\mu\nu}$, is the minimum symplectic eigenvalue of the partial transpose of a reduced $4\times 4$ CM $V_{\mu\nu}$.
-The reduced CM $V_{\mu\nu}$ contains the entries of $V$ associated with the selected bipartition $\mu$ and $\nu$, and it can be obtained by removing the rows and columns of the unwanted modes in $V$, whose $2\times 2$ block form is given by
+where $\eta_{0}^{-}\!\equiv\!2^{-1/2}\{\Sigma(V_{\mu\nu})-[\Sigma(V_{\mu\nu})^{2}-4\det\!V_{\mu\nu}]^{1/2}\}^{1/2}$, with $\Sigma(V_{\mu\nu})\!\equiv\!\det\mathcal{A}_{\mu}+\det\mathcal{B}_{\nu}-2\det\mathcal{C}_{\mu\nu}$, is the minimum symplectic eigenvalue of the partial transpose of a reduced $4\times 4$ CM $V_{\mu\nu}$, with $\mu$ and $\nu$ describing the selected modes under consideration ($\mu,\,\nu=\hat{a}_{s,\circlearrowright},\hat{a}_{s,\circlearrowleft},\hat{q}_{1},\hat{q}_{2}$).
+The reduced CM $V_{\mu\nu}$ contains the entries of $V$, and it can be obtained by selecting the rows and columns associated with modes $\mu$ and $\nu$ from $V$, whose $2\times 2$ block form is given by
 $V_{\mu\nu}=\left(\begin{matrix}\mathcal{A}_{\mu}&\mathcal{C}_{\mu\nu}\\ \mathcal{C}_{\mu\nu}^{\textit{T}}&\mathcal{B}_{\nu}\end{matrix}\right).$ (19)
-Equation (18) quantifies how much the positivity of the partial transpose condition for separability is violated for the considered Gaussian states, and it is equivalent to Simon’s necessary and sufficient entanglement nonpositive partial transpose criterion (or the related Peres-Horodecki criterion) Simon (2000).
-The selected bipartition $\mu$ and $\nu$ gets entangled if and only if $\eta_{0}^{-}<1/2$, where $E_{\mathcal{N}}$ has a nonzero value.
-To further confirm the presence and directionality of Gaussian steering between the two entangled bipartition $\mu$ and $\nu$, we introduce an intuitive and computable quantification of quantum steerability Kogias et al.
-(2015), which is defined as
-$\mathcal{G}^{\mu\rightarrow\nu}=$ $\max\left[0,\dfrac{1}{2}\ln\dfrac{\det\mathcal{A}_{\mu}}{4\det V_{\mu\nu}}\right],$
-$\mathcal{G}^{\nu\rightarrow\mu}=$ $\max\left[0,\dfrac{1}{2}\ln\dfrac{\det\mathcal{B}_{\nu}}{4\det V_{\mu\nu}}\right].$ (20)
-$\mathcal{G}^{\mu\rightarrow\nu}>0$ ($\mathcal{G}^{\nu\rightarrow\mu}>0$) implies that the bipartite Gaussian state characterized by CM $V_{\mu\nu}$ is steerable from mode $\mu$ ($\nu$) to mode $\nu$ ($\mu$) through Gaussian measurements on mode $\mu$ ($\nu$), where a higher value of $\mathcal{G}$ represents the stronger Gaussian steerability.
-The minimum residual contangle, $\mathcal{R}^{\textrm{min}}_{\tau}$, which provides a bona fide quantification of CV tripartite entanglement, is defined as Adesso and Illuminati (2007)
+Equation (18) quantifies how much the positivity of the partial transpose condition for separability is violated for the considered Gaussian states, which is equivalent to Simon’s necessary and sufficient entanglement nonpositive partial transpose criterion (or the related Peres-Horodecki criterion) [14].
+Note that the selected modes $\mu$ and $\nu$ gets entangled if and only if $\eta_{0}^{-}<1/2$, where $E_{\mathcal{N}}$ has a nonzero value.
+Figure 3: Phase-selective generation of asymmetric optomechanical EPR steering.
+The quantum steerability $\mathcal{S}^{\hat{a}_{s,i}\rightarrow\hat{q}_{1}}$ (a) and $\mathcal{S}^{\hat{q}_{1}\rightarrow\hat{a}_{s,i}}$ (b) for opposite optical input directions are plotted as functions of the scaled optical detuning $\Delta_{c}/\omega_{m}$ under different squeezing parameter $r_{d}$, with $\delta r_{\circlearrowleft}=\textrm{0}$, $\delta r_{\circlearrowright}=\textrm{0.4}$, and $\delta\theta_{\circlearrowleft}=\delta\theta_{\circlearrowright}=\pi$.
+With the increase of $r_{d}$, EPR steering can be driven from no-way regime to one-way regime for the CCW input case but not for the CW input case, implying that the asymmetry of EPR steering are effectively controlled by the phase-matched condition.
+The minimum residual contangle, $\mathcal{R}^{\textrm{min}}_{\tau}$, which provides a bona fide quantification of CV tripartite entanglement, is defined as [16]
 $\mathcal{R}^{\textrm{min}}_{\tau}=$ $\min_{(r,s,t)}\,\big{[}E^{r|st}_{\tau}-E^{r|s}_{\tau}-E^{r|t}_{\tau}\big{]},$ (20)
 where $(r,s,t)$ denotes all the possible permutations of the three-mode indexes.
-$E_{\tau}^{\mu|\nu}$ is the contangle of subsystems of $\mu$ ($\mu$ contains one mode) and $\nu$ ($\nu$ contains one or two modes), which can be defined by a proper entanglement monotone, e.g., the squared logarithmic negativity.
+$E_{\tau}$ is the one-mode-versus-(one- or two-mode) contangle, which can be defined by a proper entanglement monotone, e.g., the squared logarithmic negativity.
 Based on Eq.
-(18), the one-mode-vs-one-mode contangle $E^{r|s}_{\tau}$ and $E^{r|t}_{\tau}$ can be directly obtained by employing its definition, namely, $E_{\tau}^{\mu|\nu}\equiv[E_{\mathcal{N}}^{\mu|\nu}]^{2}$.
-However, when calculating the one-mode-vs-two-modes contangle $E^{r|st}_{\tau}$, one must alter the basic definition of Eq.
-(18) by rewriting the definition of $\eta_{0}^{-}$ as given by $\eta_{0}^{-}\equiv\min\,[\textrm{eig}|i\Omega_{3}\tilde{V}_{r|st}|]$, where $\eta_{0}^{-}$ becomes the minimum symplectic eigenvalue of the partial transpose of a $6\times 6$ CM $V$, with $\Omega_{3}=\oplus_{k=1}^{3}i\sigma_{y}$ and $\sigma_{y}$ the y-Pauli matrix.
-$\tilde{V}_{r|st}$ corresponds to the partial transpose of $V$, which connects to $V$ with the relation of $\tilde{V}_{r|st}=P_{r|st}VP_{r|st}$, and $P_{r|st}=\textrm{diag}(1,-1,1,1,1,1)$ is the partial transposition matrix.
+(18), the one-mode-versus-one-mode contangle $E^{r|s}_{\tau}$ and $E^{r|t}_{\tau}$ can be directly obtained by employing its definition, namely, $E_{\tau}^{r|s}\equiv[E_{\mathcal{N}}^{r|s}]^{2}$ and $E_{\tau}^{r|t}\equiv[E_{\mathcal{N}}^{r|t}]^{2}$.
+However, when calculating the one-mode-versus-two-modes contangle $E^{r|st}_{\tau}$, one must alter the basic definition of Eq.
+(18) by rewriting the definition of $\eta_{0}^{-}$ as given by $\eta_{0}^{-}\equiv\min\,[\textrm{eig}|i\Omega_{3}\tilde{V}_{r|st}|]$, where $\eta_{0}^{-}$ becomes the minimum symplectic eigenvalue of the partial transpose of a $6\times 6$ CM $V_{r|st}$, with $\Omega_{3}=\oplus_{k=1}^{3}i\sigma_{y}$ and $\sigma_{y}$ the y-Pauli matrix.
+$\tilde{V}_{r|st}$ corresponds to the partial transpose of $V_{r|st}$, which connects to $V_{r|st}$ with the relation of $\tilde{V}_{r|st}=P_{r|st}VP_{r|st}$, and $P_{r|st}=\textrm{diag}(1,-1,1,1,1,1)$ is the partial transposition matrix.
 In terms of $E^{s|rt}_{\tau}$ and $E^{t|sr}_{\tau}$, the corresponding partial transpose matrices are given by $P_{s|rt}=\textrm{diag}(1,1,1,-1,1,1)$ and $P_{t|sr}=\textrm{diag}(1,1,1,1,1,-1)$, respectively.
@@
 As such, if there are nonzero values of the minimum residual contangle, i.e., $\mathcal{R}_{\tau}^{\textrm{min}}>0$, one can verify that the full tripartite entanglement is present for the CV system.
-III Nonreciprocal manipulation of quantum entanglement and EPR steering
-Now we start to investigate the behaviors of quantum entanglement and EPR steering in regards to the coherent driving fields input from the CW and CCW direction in our proposed scheme.
-For this purpose, we evaluate the entanglement measures by numerically solving the dynamics of the steady-state CM $V$ (16) within the following parameters: $\kappa/2\pi=\textrm{1750\,Hz}$, $\omega_{m}/2\pi=\textrm{1945\,Hz}$, $\gamma_{m}/2\pi=\textrm{195\,Hz}$, $g_{1}/2\pi=\textrm{214\,Hz}$, $g_{2}/2\pi=\textrm{321\,Hz}$, $\chi/2\pi=\textrm{292\,Hz}$, $\bar{n}_{m}=\textrm{0.9}$, $\theta_{e}=\textrm{0}$, which is partially chosen from a recent experiment Pennetta et al.
-(2020).
-Note that, for simplicity and without loss of generality, we have supposed that the two mechanical modes are identical, i.e, $\omega_{m,1}=\omega_{m,2}=\omega_{m}$, $\gamma_{m,1}=\gamma_{m,2}=\gamma_{m}$, $\bar{n}_{m,1}=\bar{n}_{m,2}=\bar{n}_{m}$, while their COM coupling strength are asymmetric, i.e., $g_{2}/g_{1}=\textrm{1.5}$.
-This condition is convenient for producing quantum entanglement and EPR steering.
-Figure 3: Squeezing-induced nonreciprocal EPR steering.
-The quantum steerability $\mathcal{G}^{\textit{a}\rightarrow\textit{m}_{\textit{1}}}$ (a), $\mathcal{G}^{\textit{m}_{\textit{1}}\rightarrow\textit{a}}$ (b), $\mathcal{G}^{\textit{m}_{\textit{1}}\rightarrow\textit{m}_{\textit{2}}}$ (c), and $\mathcal{G}^{\textit{m}_{\textit{2}}\rightarrow\textit{m}_{\textit{1}}}$ (d) for opposite optical input directions are plotted as as functions of the scaled optical detuning $\Delta_{c}/\omega_{m}$ under different squeezing parameter $r_{d}$.
-With the increase of $r_{d}$, EPR steering can be stepwise driven from no-way regime, one-way regime to two-way regime for the CCW input case but not for the CW input case, indicating the achievement of nonreciprocal EPR steering.
-We first present how to achieve nonreciprocal bipartite and tripartite entanglement by leveraging a directional two-photon parametric driving field.
-Figure 2 shows the logarithmic negativity $E_{\mathcal{N}}^{\textit{a}|\textit{m}_{\textit{1}}}$, $E_{\mathcal{N}}^{\textit{a}|\textit{m}_{\textit{2}}}$, $E_{\mathcal{N}}^{\textit{m}_{\textit{1}}|\textit{m}_{\textit{2}}}$ and the minimum residual contangle $\mathcal{R}^{\textrm{min}}_{\tau}$ for opposite optical input directions as functions of the scaled optical detuning $\Delta_{c}/\omega_{m}$.
-Here $E_{\mathcal{N}}^{\textit{a}|\textit{m}_{\textit{j}}}$ corresponds to the logarithmic negativity of the bipartition of the optical mode and the $j$th mechanical mode, while $E_{\mathcal{N}}^{\textit{m}_{\textit{1}}|\textit{m}_{\textit{2}}}$ corresponds to that of the bipartition of the two mechanical modes.
-In terms of the CW input case, as shown in the left panels of Figs.
-2(a)-2(d), it is seen that in the absence of a two-photon parametric driving field, i.e., $r_{d}=r_{e}=0$, all of the logarithmic negativity $E_{\mathcal{N}}^{\textit{a}|\textit{m}_{\textit{1}}}$, $E_{\mathcal{N}}^{\textit{a}|\textit{m}_{\textit{2}}}$, $E_{\mathcal{N}}^{\textit{m}_{\textit{1}}|\textit{m}_{\textit{2}}}$ and the minimum residual contangle $\mathcal{R}^{\textrm{min}}_{\tau}$ are null within the considered optical detuning (yellow dashed curve), indicating that there is no bipartite and tripartite entanglement in this system when only applying a coherent driving field.
-In contrast, for the CCW input case, it is seen that in the presence of a two-photon parametric driving field (with, e.g., squeezing strength $r_{d}=r_{e}=0.3$), the profiles of $E_{\mathcal{N}}^{\textit{a}|\textit{m}_{\textit{1}}}$, $E_{\mathcal{N}}^{\textit{a}|\textit{m}_{\textit{2}}}$, $E_{\mathcal{N}}^{\textit{m}_{\textit{1}}|\textit{m}_{\textit{2}}}$ and $\mathcal{R}^{\textrm{min}}_{\tau}$ are all characterized by a sharp peak around the optical detuning at COM resonance $\Delta_{c}/\omega_{m}\sim\textrm{1}$.
-Meanwhile, as shown in the right panels of Figs.
-2(a)-2(d), the bipartite and tripartite entanglement are also sensitive to the variation of the squeezing reference angle $\theta_{d}$, which all reach their maximum values at an optimal phase angle of $\theta_{d}=\pi$.
-This result is consistent with that of the previous study Jiao et al.
-, in which it reveals that the maximum COM interaction and the minimum optical noise are both obtained under the phase-matched condition with $r_{e}-r_{d}=0$ and $\theta_{e}-\theta_{d}=\pm n\pi~{}(n=1,3,5,\ldots)$, resulting in an enhanced COM entanglement.
-By comparing these results of the CW and CCW input cases, one can intuitively find that by applying a directional two-photon parametric driving field with a proper squeezing phase, the nonreciprocal generation and manipulation of bipartite and tripartite entanglement can be achieved, namely, such entanglement can emerge only for one specific optical input direction when reversing the coherent driving field.
-Physically, the mechanism behind these results is that the directional input of a two-photon parametric driving field allows one to regulate the effective COM coupling strength for the CW and CCW modes in an asymmetric way, by which one can break the time-reversal symmetry of this system.
-This feature of the system provides a flexible tool to tailor the parameter condition for entanglement generation based on its optical input direction.
-Consequently, when preparing the system under a directional entanglement generation condition, nonreciprocal bipartite and tripartite entanglement will be achieved.
-After having established nonreciprocal entanglement, we further show how to manipulate the directionality of EPR steering.
+Besides, EPR steering is an asymmetric form of quantum entanglement, where a bipartite Gaussian state characterized by CM $V_{\mu\nu}$ may be steerable from mode $\mu$ to mode $\nu$, but not vice versa.
+To study the quantum steerability from mode $\mu$ to mode $\nu$ or from mode $\nu$ to mode $\mu$, we adopt an intuitive and computable quantification of EPR steering [17], which are defined as
+$\mathcal{S}^{\mu\rightarrow\nu}=$ $\max\left[0,\dfrac{1}{2}\ln\dfrac{\det\mathcal{A}_{\mu}}{4\det V_{\mu\nu}}\right],$
+$\mathcal{S}^{\nu\rightarrow\mu}=$ $\max\left[0,\dfrac{1}{2}\ln\dfrac{\det\mathcal{B}_{\nu}}{4\det V_{\mu\nu}}\right].$ (21)
+The quantity $\mathcal{S}$ is a monotone under Gaussian local operations and classical communication (LOCC) and the larger value of $\mathcal{S}$ implies the stronger Gaussian steerability.
+The steerability between mode $\mu$ and mode $\nu$ can be classified into three cases.
+For $\mathcal{S}^{\mu\rightarrow\nu}=\mathcal{S}^{\nu\rightarrow\mu}=0$, the state is non-steerable in any direction, which corresponds to the case of no-way steering.
+For $\mathcal{S}^{\mu\rightarrow\nu}>0$ and $\mathcal{S}^{\nu\rightarrow\mu}>0$, the state is steerable in both directions $\mu\rightarrow\nu$ and $\nu\rightarrow\mu$, which corresponds to the case of two-way steering.
+Finally, for $\mathcal{S}^{\mu\rightarrow\nu}>0$ and $\mathcal{S}^{\nu\rightarrow\mu}=0$ or $\mathcal{S}^{\nu\rightarrow\mu}>0$ and $\mathcal{S}^{\mu\rightarrow\nu}=0$, the state is only steerable in one certain direction $\mu\rightarrow\nu$ or $\nu\rightarrow\mu$, which corresponds to the case of one-way steering and reflects the asymmetric nature of entangled state.
+III Phase-selective manipulation of quantum entanglement and EPR steering
+First, we investigate the phase-dependent behaviors of optomechanical entanglement and the associated asymmetric EPR steering in Figs.
+2 and 3.
+To evaluate the entanglement measures, one can numerically calculate the steady-state CM $V$ by solving the dynamical equation (16) with the following parameters: $\kappa/2\pi=4.9\,\textrm{MHz}$, $\omega_{m}/2\pi=16\,\textrm{MHz}$, $Q_{m}\equiv\omega_{m}/\gamma_{m}=10^{5}$, $G_{1}/2\pi=0.16\,\textrm{MHz}$, $G_{2}/2\pi=0.21\,\textrm{MHz}$, $\lambda/2\pi=0.32\,\textrm{MHz}$, and $\bar{n}_{m}=\textrm{100}$.
+Here, for ensuring the validity of the phonon-hopping interaction, we have supposed that the fundamental parameters of the two mechanical modes are identical, i.e, $\omega_{m,1}=\omega_{m,2}=\omega_{m}$, $\gamma_{m,1}=\gamma_{m,2}=\gamma_{m}$, $\bar{n}_{m,1}=\bar{n}_{m,2}=\bar{n}_{m}$.
+We note that these selected parameters ensure the condition required by the resolved sideband regime and mechanical ground-state cooling, i.e., $\kappa/\omega_{m}<1$, which are feasible for current experimental techniques [66] and convenient for producing optomechanical entanglement and asymmetric EPR steering.
+In terms of the mean thermal phonon excitation number $\bar{n}_{m}$, it corresponds to a bath temperature of $T\backsimeq 73\,\textrm{mK}$, which is achievable with the use of a dilution refrigerator.
+Moreover, since the effective COM coupling strength can be exponentially enhanced by the intracavity squeezing [see Fig.
+1(b)], we have chosen a much smaller value for $G_{j}$ in our simulation, whose typical value can reach up to $G_{j}/2\pi\backsim 3.8\,\textrm{MHz}$ with a laser input power of $P=65\,\mu\textrm{W}$ in the near-field cavity optomechanical system  [66].
+And the asymmetric COM coupling strength for the two mechanical modes can be exmperimentally operated by placing the nanomechanical oscillators at different positions along the field gradient.
 In Fig.
-3, we plot the quantum steerability $\mathcal{G}^{\textit{a}\rightarrow\textit{m}_{\textit{1}}}$, $\mathcal{G}^{\textit{m}_{\textit{1}}\rightarrow\textit{a}}$, $\mathcal{G}^{\textit{m}_{\textit{1}}\rightarrow\textit{m}_{\textit{2}}}$ and $\mathcal{G}^{\textit{m}_{\textit{2}}\rightarrow\textit{m}_{\textit{1}}}$ as a function of the scaled optical detuning $\Delta_{c}/\omega_{m}$ under different squeezing parameter $r_{d}$.
-Figures 3(a) and 3(b) show the EPR steering between the optical mode and the mechanical mode for opposite optical input directions.
-For the CW input case ($r_{d}=r_{e}=0$), since the optomechanical entanglement is absent, EPR steering, as a stronger quantum correlaiton, can also not be produced, where $\mathcal{G}^{\textit{a}\rightarrow\textit{m}_{\textit{1}}}$ and $\mathcal{G}^{\textit{m}_{\textit{1}}\rightarrow\textit{a}}$ always keep zero with the variation of the controlling parameters [see scattered dotted curve in Figs.
+2, we present how to achieve phase-selective generation of quantum entanglement regarding the bipartition of optical and mechanical modes.
+Figures 2(a) and 2(c) show the logarithmic negativity $E_{\mathcal{N}}^{\hat{a}_{s,i}|\hat{q}_{1}}$ and $E_{\mathcal{N}}^{\hat{a}_{s,i}|\hat{q}_{2}}$ for opposite optical input directions as functions of the scaled optical detuning $\Delta_{c}/\omega_{m}$.
+Here $E_{\mathcal{N}}^{\hat{a}_{s,i}|\hat{q}_{j}}$ corresponds to the logarithmic negativity of the bipartition of the squeezed optical mode $i$ and the mechanical mode $j$.
+As discussed in previous section, we have assumed the fulfillment of the phase-matched condition for the CCW input case [i.e., $\delta r_{\circlearrowleft}=\textrm{0}$ and $\delta\theta_{\circlearrowleft}=\pm n\pi(n=1,3,5...)$], whereas such condition is phase-mismatched for the CW input case [i.e., $\delta r_{\circlearrowright}\neq\textrm{0}$ and $\delta\theta_{\circlearrowright}\neq\pm n\pi~{}(n=1,3,5...)$].
+Under this circumstance, it is seen that the profiles of $E_{\mathcal{N}}^{\hat{a}_{s,\circlearrowleft}|\hat{q}_{1}}$ and $E_{\mathcal{N}}^{\hat{a}_{s,\circlearrowleft}|\hat{q}_{2}}$ for the CCW input case are characterized by a sharp peak around the optical detuning at COM resonance $\Delta_{c}/\omega_{m}\sim\textrm{1}$ (purple solid curve).
+In contrast, when reversing the coherent driving field, the values of the logarithmic negativity $E_{\mathcal{N}}^{\hat{a}_{s,\circlearrowright}|\hat{q}_{1}}$ and $E_{\mathcal{N}}^{\hat{a}_{s,\circlearrowright}|\hat{q}_{2}}$ for the CW input case become smaller or even zero within the same considered optical detuning (green dashed curve).
+These results imply that by setting distinct squeezing phase parameters for the counterpropagating optical modes, it could lead to asymmetric response in the entanglement generation between the coherent driving field and the mechanical mode.
+Moreover, it is clearly seen that the response asymmetry of $E_{\mathcal{N}}^{\hat{a}_{s,i}|\hat{q}_{j}}$ is dependent on both $\delta r_{i}$ and $\delta\theta_{i}$, with stronger asymmetry attainable through appropriate optimization of the squeezing phase condition.
+To clearly visualize this behavior, we define the asymmetric response ratio $\mathcal{I}_{j}$ as
+$\mathcal{I}_{j}=\dfrac{E_{\mathcal{N}}^{\hat{a}_{s,\circlearrowleft}|\hat{q}_{j}}-E_{\mathcal{N}}^{\hat{a}_{s,\circlearrowright}|\hat{q}_{j}}}{E_{\mathcal{N}}^{\hat{a}_{s,\circlearrowleft}|\hat{q}_{j}}}.$ (22)
+As shown in Figs.
+2(b) and 2(d), the behavior of $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ are similar with the variation of the squeezing parameters, which show greater sensitivity to $\delta r_{\circlearrowright}$ than to $\delta\theta_{\circlearrowright}$ and these isolation ratios could be more than 0.9 when choosing $\delta r_{\circlearrowright}>\textrm{0.2}$ with proper values of $\delta\theta_{\circlearrowright}$.
+Besides, we have confirmed that the maximum values of $\mathcal{I}_{j}$ could even reach up to $1$ by properly choosing squeezing phase conditions, which means that such optomechanical entanglement can be unidirectionally produced only in one specific optical input direction when reversing the coherent driving field.
+Additionally, to investigate the robustness of such optomechanical entanglement against thermal noises, we further plot $E_{\mathcal{N}}^{\hat{\textit{a}}_{\textit{s},\textit{i}}|\hat{\textit{q}}_{\textit{1}}}$ and $E_{\mathcal{N}}^{\hat{\textit{a}}_{\textit{s},\textit{i}}|\hat{\textit{q}}_{\textit{2}}}$ for opposite optical input directions as functions of the mechanical environment temperature $T$ in Figs.
+2(e) and 2(f).
+It is seen that in comparison with the CW input case, the entanglement with respect to the CCW input case is more robust against thermal noises, which can persist above a temperature of 110mK.
+Figure 4: Phase-selective generation of quantum entanglement and asymmetric EPR steering with respect to the two mechanical modes.
+(a),(b) The logarithmic negativity $E_{\mathcal{N}}^{\hat{q}_{1}|\hat{q}_{2}}$ and the minimum residual contangle $\mathcal{R}^{\textrm{min}}_{\tau}$ for opposite optical input directions are plotted as functions of the scaled optical detuning $\Delta_{c}/\omega_{m}$ with $r_{d}=0.2$ and $\theta_{d}=\pi$.
+The inserts show the dependence of $E_{\mathcal{N}}^{\hat{q}_{1}|\hat{q}_{2}}$ and $\mathcal{R}^{\textrm{min}}_{\tau}$ on the squeezing strength difference $\delta r_{\circlearrowright}$ and the squeezing phase difference $\delta\theta_{\circlearrowright}$ at the optical detuning $\Delta_{c}/\omega_{m}=1.2$ and $1.15$, respectively.
+(c),(d) The measure of quantum steerability $\mathcal{S}^{\hat{q}_{1}\rightarrow\hat{q}_{2}}$ and $\mathcal{S}^{\hat{q}_{2}\rightarrow\hat{q}_{1}}$ for opposite optical input directions are plotted as functions of the scaled optical detuning $\Delta_{c}/\omega_{m}$ under different squeezing strength $r_{d}$ with $\theta_{d}=\pi$.
+With the increase of $r_{d}$, EPR steering for the CCW input case can be stepwise driven from no-way regime, one-way regime to two-way regime, whereas for the CW input case, it can only be driven from no-way regime to one-way regime.
+Subsequently, we further investigate how to regulate the steerability of optomechanical entanglement and achieve one-way EPR steering through tuning the squeezing phase conditions.
+For this purpose, we plot the measure of quantum steerability $\mathcal{S}^{\hat{a}_{s,i}\rightarrow\hat{q}_{1}}$ and $\mathcal{S}^{\hat{q}_{1}\rightarrow\hat{a}_{s,i}}$ as functions of the scaled optical detuning $\Delta_{c}/\omega_{m}$ under different values of $r_{d}$ in Fig.
+3.
+It is seen that in terms of the CCW input case, for small values of $r_{d}$, both $\mathcal{S}^{\hat{a}_{s,\circlearrowleft}\rightarrow\hat{q}_{1}}$ and $\mathcal{S}^{\hat{q}_{1}\rightarrow\hat{a}_{s,\circlearrowleft}}$ are null within the considered optical detuning [see red solid lines in Figs.
+3(a) and 3(b)], implying that the EPR steering are not present for both directions of $\hat{a}_{s,\circlearrowleft}\rightarrow\hat{q}_{1}$ and $\hat{q}_{1}\rightarrow\hat{a}_{s,\circlearrowleft}$.
+When increasing the values of the intracavity squeezing strength $r_{d}$, the quantum steerability $\mathcal{S}^{\hat{a}_{s,\circlearrowleft}\rightarrow\hat{q}_{1}}$ will increase monotonically and become nonzero [see Fig.
+3(a)], resulting in an asymmetric one-way EPR steering with $\mathcal{S}^{\hat{a}_{s,\circlearrowleft}\rightarrow\hat{q}_{1}}>0$ and $\mathcal{S}^{\hat{q}_{1}\rightarrow\hat{a}_{s,\circlearrowleft}}=0$.
+This is because quantum steerability is an increasing function of the effective COM coupling strength that can be enhanced by the intracavity squeezing strength $r_{d}$ [see Fig.
+1(c) for more details].
+However, for the CW input case, since the optomechanical entanglement is absent, EPR steering, as a stronger correlated entangled state, can also not be produced, showing that $\mathcal{S}^{\hat{a}_{s,\circlearrowright}\rightarrow\hat{q}_{1}}$ and $\mathcal{S}^{\hat{q}_{1}\rightarrow\hat{a}_{s,\circlearrowright}}$ always keep zero with the variation of the controlling parameters [see green ball-solid lines in Figs.
 3(a) and 3(b)].
-For the CCW input case, due to the introduction of the squeezed optical mode and the emergence of the optomechanical entanglement, one can create an asymmetric and directional EPR steering by adjusting the squeezing parameters.
-For instance, it is seen that although when choosing $\theta_{d}=\pi$ and some small values of $r_{d}$, such as $r_{d}=0.15$, the EPR steering still does not exist for both $\textit{a}\rightarrow\textit{m}_{\textit{1}}$ and $\textit{m}_{\textit{1}}\rightarrow\textit{a}$, i.e., $\mathcal{G}^{\textit{a}\rightarrow\textit{m}_{\textit{1}}}=\mathcal{G}^{\textit{m}_{\textit{1}}\rightarrow\textit{a}}=0$ [solid curve in the top panels of Figs.
-3(a) and 3(b)].
-However, when the squeezing strength $r_{d}$ is continued to be raised, the value of $\mathcal{G}^{\textit{a}\rightarrow\textit{m}_{\textit{1}}}$ will become nonzero and monotonically increase, resulting in an asymmetric one-way EPR steering with $\mathcal{G}^{\textit{a}\rightarrow\textit{m}_{\textit{1}}}>0$ and $\mathcal{G}^{\textit{m}_{\textit{1}}\rightarrow\textit{a}}=0$.
-Moreover, we also show the EPR steering between the two mechanical modes for opposite optical input directions in Figs.
-3(c) and 3(d).
-For the CW input case, there is no EPR steering for both $\textit{m}_{\textit{1}}\rightarrow\textit{m}_{\textit{2}}$ and $\textit{m}_{\textit{2}}\rightarrow\textit{m}_{\textit{1}}$, i.e., $\mathcal{G}^{\textit{m}_{\textit{1}}\rightarrow\textit{m}_{\textit{2}}}=\mathcal{G}^{\textit{m}_{\textit{2}}\rightarrow\textit{m}_{\textit{1}}}=0$ [see scattered dotted curve in Figs.
-3(c) and 3(d)], which is similar to the case discussed above for the optomechanical bipartition.
-For the CCW input case, accompanied by the occurrence of entanglement, we can get rich properties of EPR steering, where the overall state’s asymmetry is stepwise driven through the no-way regime ($\mathcal{G}^{\textit{m}_{\textit{1}}\rightarrow\textit{m}_{\textit{2}}}=\mathcal{G}^{\textit{m}_{\textit{2}}\rightarrow\textit{m}_{\textit{1}}}=0$), one-way regime ($\mathcal{G}^{\textit{m}_{\textit{1}}\rightarrow\textit{m}_{\textit{2}}}=0$ but $\mathcal{G}^{\textit{m}_{\textit{2}}\rightarrow\textit{m}_{\textit{1}}}>0$), and finally two-way regime ($\mathcal{G}^{\textit{m}_{\textit{1}}\rightarrow\textit{m}_{\textit{2}}}>0$ and $\mathcal{G}^{\textit{m}_{\textit{2}}\rightarrow\textit{m}_{\textit{1}}}>0$) with the increase of the squeezing strength $r_{d}$.
-From the above discussions, it is found that EPR steering can be generated and manipulated only for the CCW input case but not for the CW input case, which implies the achievement of nonreciprocal EPR steering.
+By comparing the results of Figs.
+2 and 3, it can be intuitively found that when optomechanical entanglement is prepared in an asymmetric way, the generation of the corresponding EPR steering also exhibits asymmetric characteristics, with its directionality being asymmetrically controlled via intracavity squeezing parameters.
+Physically, the mechanism behind these phenomena can be understood as follows: As discussed previously in Fig.
+1(c), the optical input noises $N_{s,i}$ and $M_{s,i}$ can either be reduced or enhanced through adjusting the squeezing parameters $r_{e,i}$ and $\theta_{e,i}$, which provides an effective method to regulate the noise flows in CW and CCW modes in a controllable way.
+Here, as a specific example, we have chosen the parameter conditions where optical input noises are reduced for the CCW input case [corresponding to phase-matched condition] but amplified for the CW input case [corresponding to phase-mismatched condition].
+This squeezing-phase-induced asymmetry could lead to differences in the quantum statistical properties of the CW and CCW modes.
+As shown in Fig.
+1(d), without the presence of coherent driving, the steady-state CM $V$ displays a dominant diagonal, which is consistent with a Gaussian distribution of a vacuum state.
+For the CCW input case, except for the dominant diagonal elements, the steady-state CM $V$ also displays some off-diagonal elements, which are not only present in the mechanical bipartition, but also emerges in the optomechanical bipartition.
+For the CW input case, due to the optical input noises $N_{s,\circlearrowright}\neq 0$ and $M_{s,\circlearrowright}\neq 0$, the off-diagonal elements with respect to the optomechanical bipartition vanish in steady-state CM $V$.
+These results indicate that the quantum correlations between the squeezed optical mode and the mechanical modes, enabled by the down-conversion interaction through nonlinear COM coupling, survive only with small thermal photon occupancy.
+Therefore, by asymmetrically choosing phase conditions for the CW and CCW modes, the phase-selective generation of optomechanical entanglement and EPR steering could be achieved in this configuration.
+Finally, we further present how to achieve phase-selective generation of quantum entanglement and EPR steering involving two mechanical modes in Fig.
+4.
+Given that the mechanical entanglement is more fragile to thermal noises, the parameters for numerical simulation are altered as follows: $\kappa/\textrm{2}\pi=\textrm{14.4}\,\textrm{MHz}$, $\omega_{m}/\textrm{2}\pi=\textrm{16}\,\textrm{MHz}$, $Q_{m}\equiv\omega_{m}/\gamma_{m}=\textrm{10}$, $G_{1}/\textrm{2}\pi=\textrm{1.6}\,\textrm{MHz}$, $G_{2}/\textrm{2}\pi=\textrm{2.1}\,\textrm{MHz}$, $\lambda/\textrm{2}\pi=\textrm{4.8}\,\textrm{MHz}$, and $\bar{n}_{m}=\textrm{0.9}$.
+Figures 4(a) and 4(b) show the logarithmic negativity $E_{\mathcal{N}}^{\hat{q}_{1}|\hat{q}_{2}}$ and the minimum residual contangle $\mathcal{R}^{\textrm{min}}_{\tau}$ versus the scaled optical detuning $\Delta_{c}/\omega_{m}$ for opposite optical input directions.
+It is seen that for the CCW input case, by fulfilling the phase-matched condition, the bipartite mechanical entanglement and the tripartite entanglement are present within a finite interval of $\Delta_{c}$ around $\Delta_{c}/\omega_{m}\simeq\textrm{1.2}$.
+In contrast, for the CW input case, while the bipartite mechanical entanglement and the tripartite entanglement can still persist for certain values of mismatched phase (e.g., $\delta r_{\circlearrowright}=\textrm{0.4}$ and $\delta\theta_{\circlearrowright}=\pi$), their maximum values are suppressed compared to those in the CCW input case.
+Moreover, from the inserts of Figs.
+4(a) and 4(b), it is also seen that when setting squeezing phase difference $\delta\theta_{\circlearrowright}=\textrm{0}$, the corresponding bipartite and tripartite entanglement can completely vanish at high values of $\delta r_{\circlearrowright}$.
+These results are consistent with the previous discussions for optomechanical entanglement, implying the achievement of phase-selective asymmetric response in bipartite mechanical entanglement and tripartite entanglement.
+In Figs.
+4(c) and 4(d), to explore the behaviors of EPR steering between the two mechanical modes, we further plot the quantum steerability $\mathcal{S}^{\hat{q}_{1}\rightarrow\hat{q}_{2}}$ and $\mathcal{S}^{\hat{q}_{2}\rightarrow\hat{q}_{1}}$ versus the scaled optical detuning $\Delta_{c}/\omega_{m}$ under different squeezing strength $r_{d}$ with $\theta_{d}=\pi$.
+For the CCW input case, it is seen that accompanied with the occurrence of highly mechanical entangled state, we can get rich properties of EPR steering, where the overall state’s asymmetry can be stepwise driven through the no-way regime ($\mathcal{S}^{\hat{q}_{1}\rightarrow\hat{q}_{2}}=\mathcal{S}^{\hat{q}_{2}\rightarrow\hat{q}_{1}}=0$), one-way regime ($\mathcal{S}^{\hat{q}_{2}\rightarrow\hat{q}_{1}}=0$ but $\mathcal{S}^{\hat{q}_{1}\rightarrow\hat{q}_{2}}>0$), and finally two-way regime ($\mathcal{S}^{\hat{q}_{1}\rightarrow\hat{q}_{2}}>0$ and $\mathcal{S}^{\hat{q}_{2}\rightarrow\hat{q}_{2}}>0$) with the increase of the intracavity squeezing strength $r_{d}$.
+However, for the CW input case, due to the phase-mismatch induced suppression of mechanical entanglement, it is seen that the directionality of EPR steering can only be driven from no-way regime to one-way regime under the same intracavity squeezing strength.
+From the above discussions, it is found that the mechanical EPR steering can be generated and manipulated in an asymmetric way.
 These results can be understood as follows: The proposed three-mode COM system has a bilinear cyclic coupling among the optical mode and the two mechanical modes.
 For an arbitrary bipartition $\mu$ and $\nu$, apart from their direct interaction, there is also an indirect interaction induced by their coupling to the third intermediate mode.
-Both direct and indirect interaction paths can generate EPR steering between $\mu$ and $\nu$, whose superposition may induce quantum interference effect that depends on the relative phase and strength of the two interaction paths.
-As discussed previously, by applying a directional two two-photon parametric driving field, one can manipulate the COM coupling strength in an asymmetric way.
-In this case, the relative strength of the two interaction paths for any bipartition of this system is influenced, based on which the manipulation of the directionality of EPR steering can be achieved by tuning the squeezing strength.
+The superposition of direct and indirect interaction paths can give rise to a quantum interference effect, whose strength is determined by their relative phase and coupling strength.
+Because both quantum entanglement and EPR steering increase monotonically with the coupling strength, their magnitudes can thus be modulated via the interference effect, resulting in the generation of tripartite entanglement.
+In this case, by regulating the two effective COM coupling strengths $\Lambda_{j}$ through tuning the intracavity squeezing $r_{d}$, the entanglement originally arising from nonlinear COM interactions could be transferred to the two mechanical modes via the interference effect.
+Moreover, since we have assumed different phase conditions for the CW and CCW modes, the bipartite and tripartite entanglement will be produced in an asymmetric way when the coherent driving field is applied from opposite directions.
+Also, due to the asymmetric COM coupling induced for the two mechanical modes, the overall asymmetry of EPR steering can be well controlled by the squeezing phase condition.
 IV Conclusion
-In summary, we have presented how to generate and manipulate squeezing-induced nonreciprocal quantum entanglement and asymmetric EPR steering in a three-mode COM system consisting of a bottle WGM resonator, an optical waveguide, and a tapered glass-fiber nanospike.
-The WGM resonator supports two counterpropagating optical modes, which are uncoupled and degenerate.
+In summary, we have presented how to achieve phase-selective generation and manipulation of quantum entanglement and asymmetric EPR steering in a three-mode squeezed COM system.
+Here, as a specific example, we have considered a near-field cavity COM setup to realize our proposal [66, 67], which consists of a microdisk WGM resonator, a string of nanomechanical oscillators, and a tapered optical fiber.
+The WGM resonator supports two counterpropagating optical modes, which are degenerate and stay uncoupled.
 The coherent driving field from the waveguide is coupled into and out of the WGM resonator via evanescent coupling.
-Moreover, when placing the nanospike close to the WGM resonator, it excites two mechanical modes due to the COM interaction induced by the WGM evanescent field, whose vibrating directions are orthogonal or parallel to the resonator surface.
-First, we show that when applying an additional two-photon parametric driving field for one specific input direction of the WGM resonator with a phase-matched squeezed vacuum reservoir, one can introduce a squeezed optical mode and an enhanced COM interaction in this input direction.
-This directional injection of quantum squeezing breaks the time-reversal symmetry of the system, and it allows one to tailor the parameter condition for entanglement generation with respect to its optical input direction, which is essential for achieving quantum nonreciprocity.
-Second, based on this unique feature of the system, we find that when reversing the input direction of the coherent driving field, quantum entanglement and the associated EPR steering is regulated in an asymmetric way, whereby a controllable nonreciprocal generation and manipulation of quantum entanglement and EPR steering is achieved.
-More interestingly, it is also found that by properly tuning the squeezing parameters, the directionality of EPR steering can be well controlled.
-According to this, we further show that with the increase of the squeezing strength, EPR steering can be stepwise driven from no-way regime, one-way regime to two-way regime.
-These results might open up new perspectives for the experimental generation and application of nonreciprocal bipartite and tripartite entanglement and asymmetric EPR steering as a precious resource for a variety of nascent quantum technologies ranging from one-sided device-independent quantum key distribution to no-cloning quantum teleportation Branciard et al.
-(2012); Walk et al.
-(2016); Li et al.
-(2024); He et al.
-(2015b); Chiu et al.
-(2016).
+Moreover, when placing the nanomechanical array close to the WGM resonator, it excites two mechanical modes due to the COM interaction induced by the WGM evanescent field.
+First, we show that when simultaneously applying two-photon pumping fields and broadband squeezed optical fields to CW and CCW modes, one can introduce a pair of squeezed optical modes, with each coupled to an independent squeezed vacuum reservoir.
+This configuration enables a phase-controlled light-reservoir interaction and allows one to tailor asymmetric noise flows for opposite input directions, which plays a key role in breaking the symmetry of the quantum statistical properties of the counterpropagating modes.
+Second, based on this unique feature of the squeezed COM system, it is found that the generation of various types of quantum entanglement and the associated EPR steering become phase-dependent and thus they can be produced in an asymmetric way.
+More interestingly, it is also found that by properly tuning the intracavity squeezing parameters, the directionality of EPR steering can be well controlled.
+According to this, we further show that with the increase of the squeezing strength, the directionality of EPR steering can be stepwise driven from no-way regime, one-way regime to two-way regime.
+Note that our proposal can also be realized with various types of multimode COM system, e.g., microwave electromechanical system [23, 24], cavity magnomechanical system [76] , and nanospike-WGM optomechanical system [70].
+These findings, opening up new perspectives for preparing rich types of quantum entanglement and EPR steering with asymmetric features, might be applicable for a variety of nascent quantum technologies ranging from one-sided device-independent quantum key distribution to no-cloning quantum teleportation [9, 10, 11, 12, 13].
 V Acknowledgment
@@
 11935006, 12421005), the National Key R&\&&D Program of China (Grants No.
-2024YFE0102400), the Hunan Provincial Major Sci-Tech Program (Grant No.
-2023ZJ1010), and the Science and Technology Innovation Program of Hunan Province (Grant No.
-2020RC4047).
+2024YFE0102400), and the Hunan Provincial Major Sci-Tech Program (Grant No.
+2023ZJ1010).
 L.-M.K.
@@
 is supported by the NSFC (Grant No.
-12405029).
+12405029) and the Natural Science Foundation of Henan Province (Grant No.
+252300421221).
 D.-Y.W.
@@
 241100210400).
+Y.-L.Z.
+is supported by the Natural Science Foundation of Hunan Province (Grant No.
+2025JJ60018) and the Scientific Research Foundation of Education Bureau of Hunan Province (Grant No.24B0866).
 Appendix A Derivation of the effective Hamiltonian
-To diagonalize the optical mode of Hamiltonian (1), one can perform a unitary Bogoliubov transformation with $S(\eta_{d})=\exp[(-\eta_{d}\hat{a}_{\circlearrowleft}^{\dagger 2}+\eta_{d}^{\ast}\hat{a}_{\circlearrowleft}^{2})/2]$, where $\eta_{d}=r_{d}e^{-i\theta_{d}}$ is a complex squeezing parameter with $r_{d}$ the squeezing strength and $\theta_{d}$ the squeezing reference angle.
+To diagonalize the optical mode of Hamiltonian (1), one can perform a unitary Bogoliubov transformation with $U(\eta_{d})=\exp[(-\eta_{d}\hat{a}_{\circlearrowleft}^{\dagger 2}+\eta_{d}^{\ast}\hat{a}_{\circlearrowleft}^{2})/2]$, where $\eta_{d}=r_{d}e^{-i\theta_{d}}$ is a complex squeezing parameter with $r_{d}$ the squeezing strength and $\theta_{d}$ the squeezing reference angle.
 By implementing this Bogoliubov transformation, a squeezed optical mode is introduced and the system Hamiltonian (1) becomes
-$\hat{\tilde{H}}_{\circlearrowleft}=$ $S^{\dagger}(\eta_{d})\hat{H}_{\circlearrowleft}S(\eta_{d})$
-$=$ $\dfrac{\omega_{m,1}}{2}(\hat{p}_{1}^{2}+\hat{q}_{1}^{2})+\dfrac{\omega_{m,2}}{2}(\hat{p}_{2}^{2}+\hat{q}_{2}^{2})+\chi\hat{q}_{1}\hat{q}_{2}$
+$\hat{\tilde{H}}_{\circlearrowleft}=$ $U^{\dagger}(\eta_{d})\hat{H}_{\circlearrowleft}U(\eta_{d})$
+$=$ $\dfrac{\omega_{m,1}}{2}(\hat{p}_{1}^{2}+\hat{q}_{1}^{2})+\dfrac{\omega_{m,2}}{2}(\hat{p}_{2}^{2}+\hat{q}_{2}^{2})+\lambda\hat{q}_{1}\hat{q}_{2}$
 $+\left[\Delta_{c}\cosh(2r_{d})-2\Xi_{d}\sinh(2r_{d})\right]\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{a}_{s,\circlearrowleft}$
@@
 $-\zeta_{s,2}\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{a}_{s,\circlearrowleft}\hat{q}_{2}+\dfrac{\zeta_{p,2}}{2}(e^{-i\theta_{d}}\hat{a}_{s,\circlearrowleft}^{\dagger 2}+e^{i\theta_{d}}\hat{a}_{s,\circlearrowleft}^{2})\hat{q}_{2}$
-$\!-\!F_{1}\hat{q}_{1}\!-\!F_{2}\hat{q}_{2}\!+\!\chi\hat{q}_{1}\hat{q}_{2}\!+\!i\varepsilon_{d,\circlearrowleft}\!\cosh(r_{d})(\hat{a}_{s,\circlearrowleft}^{\dagger}\!-\!\hat{a}_{s,\circlearrowleft})$
+$\!-\!F_{1}\hat{q}_{1}\!-\!F_{2}\hat{q}_{2}\!+\!\lambda\hat{q}_{1}\hat{q}_{2}\!+\!i\varepsilon_{d,\circlearrowleft}\!\cosh(r_{d})(\hat{a}_{s,\circlearrowleft}^{\dagger}\!-\!\hat{a}_{s,\circlearrowleft})$
 $+i\varepsilon_{d,\circlearrowleft}\sinh(r_{d})(e^{-i\theta_{d}}\hat{a}_{s,\circlearrowleft}^{\dagger}-e^{i\theta_{d}}\hat{a}_{s,\circlearrowleft})+C,$ (25)
@@
 Appendix B Derivation of the effective master equation under phase-matched condition
-For the CCW input case, to eliminate the optical dissipation induced by the intracavity squeezing, we consider the injection of a broadband squeezed optical field in this direction, which can act as a squeezed-vacuum reservoir.
+To regulate the light-reservoir interaction, we have considered the injection of a broadband squeezed optical field, which can act as a squeezed-vacuum reservoir.
 We assume the squeezed optical field, with squeezing strength $r_{e}$ and reference phase angle $\theta_{e}$, is around the central frequency $\omega_{c}$.
 Besides, the two mechanical modes are assumed to be coupled with two independent thermal reservoirs at the same bath temperature $T$.
-Then, including the dissipations of the optical and mechanical modes, the dynamics of the total system in the original picture is governed by the Born-Markovian master equation Gardiner and Zoller (2000)
-$\dot{\rho}_{\circlearrowleft}=$ $i\big{[}\rho_{\circlearrowleft},\hat{H}_{\circlearrowleft}\big{]}+\frac{\kappa}{2}(N_{e}+1)\mathcal{D}[\hat{a}_{\circlearrowleft}]\rho_{\circlearrowleft}+\frac{\kappa}{2}N_{e}\mathcal{D}[\hat{a}_{\circlearrowleft}^{\dagger}]\rho_{\circlearrowleft}$
-$-\frac{\kappa}{2}M_{e}\mathcal{G}[\hat{a}_{\circlearrowleft}]\rho_{\circlearrowleft}-\frac{\kappa}{2}M_{e}^{\ast}\mathcal{G}[\hat{a}_{\circlearrowleft}^{\dagger}]\rho_{\circlearrowleft}$
+Then, including the dissipations of the optical and mechanical modes, the dynamics of the total system in the original picture is governed by the Born-Markovian master equation [75]
+$\dot{\rho}_{\circlearrowleft}=$ $i\big{[}\rho_{\circlearrowleft},\hat{H}_{\circlearrowleft}\big{]}+\frac{\kappa}{2}(N_{e,\circlearrowleft}+1)\mathcal{D}[\hat{a}_{\circlearrowleft}]\rho_{\circlearrowleft}+\frac{\kappa}{2}N_{e,\circlearrowleft}\mathcal{D}[\hat{a}_{\circlearrowleft}^{\dagger}]\rho_{\circlearrowleft}$
+$-\frac{\kappa}{2}M_{e,\circlearrowleft}\mathcal{G}[\hat{a}_{\circlearrowleft}]\rho_{\circlearrowleft}-\frac{\kappa}{2}M_{e,\circlearrowleft}^{\ast}\mathcal{G}[\hat{a}_{\circlearrowleft}^{\dagger}]\rho_{\circlearrowleft}$
 $-\sum_{j=1,2}\left(i\dfrac{\gamma_{m,j}}{2}[\hat{q}_{j},\{\hat{p}_{j},\rho_{\circlearrowleft}\}]+\gamma_{m,j}\bar{n}_{m,j}[\hat{q}_{j},[\hat{q}_{j},\rho_{\circlearrowleft}]]\right),$ (27)
@@
 and $\kappa~{}(\gamma_{m,j})$ is the optical (mechanical) decay rate.
-$\bar{n}_{m,j}=1/[\exp(\omega_{m,j}/k_{B}T)-1]$ is the thermal phonon number, while $N_{e}=\sinh^{2}(r_{e})$ and $M_{e}=\cosh(r_{e})\sinh(r_{e})e^{i\theta_{e}}$ describe the dissipation and the two-photon correlation of the cavity field caused by the squeezed-vacuum reservoir, respectively.
+$\bar{n}_{m,j}=1/[\exp(\omega_{m,j}/k_{B}T)-1]$ is the thermal phonon number, while $N_{e,\circlearrowleft}=\sinh^{2}(r_{e})$ and $M_{e,\circlearrowleft}=\cosh(r_{e})\sinh(r_{e})e^{i\theta_{e}}$ describe the dissipation and the two-photon correlation of the cavity field caused by the squeezed-vacuum reservoir, respectively.
 By introducing the Bogoliubov transformation in Eq. (4), the Lindblad operators becomes
-$S^{\dagger}(\eta_{d})\mathcal{D}[\hat{a}_{\circlearrowleft}]\rho_{\circlearrowleft}S(\eta_{d})$
+$U^{\dagger}(\eta_{d})\mathcal{D}[\hat{a}_{\circlearrowleft}]\rho_{\circlearrowleft}U(\eta_{d})$
 $=$ $2\cosh^{2}(r_{d})\hat{a}_{s,\circlearrowleft}\tilde{\rho}_{\circlearrowleft}\hat{a}_{s,\circlearrowleft}^{\dagger}+2\sinh^{2}(r_{d})\hat{a}_{s,\circlearrowleft}^{\dagger}\tilde{\rho}_{\circlearrowleft}\hat{a}_{s,\circlearrowleft}$
@@
 $\left.-\sinh(r_{d})\cosh(r_{d})(e^{-i\theta_{d}}\tilde{\rho}_{\circlearrowleft}\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{a}_{s,\circlearrowleft}^{\dagger}+e^{i\theta_{d}}\tilde{\rho}_{\circlearrowleft}\hat{a}_{s,\circlearrowleft}\hat{a}_{s,\circlearrowleft})\right],$
-$S^{\dagger}(\eta_{d})\mathcal{G}[\hat{a}_{\circlearrowleft}]\rho_{\circlearrowleft}S(\eta_{d})$
+$U^{\dagger}(\eta_{d})\mathcal{G}[\hat{a}_{\circlearrowleft}]\rho_{\circlearrowleft}U(\eta_{d})$
 $=$ $2e^{-2i\theta_{d}}\sinh^{2}(r_{d})\hat{a}_{s,\circlearrowleft}^{\dagger}\tilde{\rho}_{\circlearrowleft}\hat{a}_{s,\circlearrowleft}^{\dagger}+2\cosh^{2}(r_{d})\hat{a}_{s,\circlearrowleft}\tilde{\rho}_{\circlearrowleft}\hat{a}_{s,\circlearrowleft}$
@@
 In this case, the effective Born-Markovian master equation could be rewritten as
-$\dot{\tilde{\rho}}_{\circlearrowleft}=$ $S^{\dagger}(\eta_{d})\dot{\rho}_{\circlearrowleft}S(\eta_{d})$
-$=$ $i\big{[}\tilde{\rho}_{\circlearrowleft},\hat{\tilde{H}}_{\circlearrowleft}\big{]}+\frac{\kappa}{2}(N_{s}+1)\mathcal{D}[\hat{a}_{s,\circlearrowleft}]\tilde{\rho}_{\circlearrowleft}+\frac{\kappa}{2}N_{s}\mathcal{D}[\hat{a}_{s,\circlearrowleft}^{\dagger}]\tilde{\rho}_{\circlearrowleft}$
-$-\frac{\kappa}{2}M_{s}\mathcal{G}[\hat{a}_{s,\circlearrowleft}]\tilde{\rho}_{\circlearrowleft}-\frac{\kappa}{2}M_{s}^{\ast}\mathcal{G}[\hat{a}_{s,\circlearrowleft}]\tilde{\rho}_{\circlearrowleft}$
+$\dot{\tilde{\rho}}_{\circlearrowleft}=$ $U^{\dagger}(\eta_{d})\dot{\rho}_{\circlearrowleft}U(\eta_{d})$
+$=$ $i\big{[}\tilde{\rho}_{\circlearrowleft},\hat{\tilde{H}}_{\circlearrowleft}\big{]}+\frac{\kappa}{2}(N_{s,\circlearrowleft}+1)\mathcal{D}[\hat{a}_{s,\circlearrowleft}]\tilde{\rho}_{\circlearrowleft}+\frac{\kappa}{2}N_{s,\circlearrowleft}\mathcal{D}[\hat{a}_{s,\circlearrowleft}^{\dagger}]\tilde{\rho}_{\circlearrowleft}$
+$-\frac{\kappa}{2}M_{s,\circlearrowleft}\mathcal{G}[\hat{a}_{s,\circlearrowleft}]\tilde{\rho}_{\circlearrowleft}-\frac{\kappa}{2}M_{s,\circlearrowleft}^{\ast}\mathcal{G}[\hat{a}_{s,\circlearrowleft}]\tilde{\rho}_{\circlearrowleft}$
 $-\sum_{j=1,2}\left(i\dfrac{\gamma_{m,j}}{2}[\hat{q}_{j},\{\hat{p}_{j},\tilde{\rho}_{\circlearrowleft}\}]+\gamma_{m,j}\bar{n}_{m,j}[\hat{q}_{j},[\hat{q}_{j},\tilde{\rho}_{\circlearrowleft}]]\right),$ (30)
 where
-$N_{s}=$ $\sinh^{2}(r_{d})\cosh^{2}(r_{e})+\cosh^{2}(r_{d})\sinh^{2}(r_{e})$
-$+\dfrac{1}{2}\cos(\theta_{e}-\theta_{d})\sinh(2r_{d})\sinh(2r_{e}),$
-$M_{s}=$ $e^{i\theta_{d}}[\cosh(r_{d})\cosh(r_{e})+e^{-i(\theta_{e}-\theta_{d})}\sinh(r_{d})\sinh(r_{e})]$
-×$\times[\sinh(r_{d})\cosh(r_{e})+e^{i(\theta_{e}-\theta_{d})}\cosh(r_{d})\sinh(r_{e})].$ (30)
-Here $N_{s}$ and $M_{s}$ denote the effective thermal noise and two-photon-correlation strength, respectively, which can be simplified in case of $r_{d}=r_{e}=r$, i.e.,
-$N_{s}=$ $\dfrac{1}{2}\sinh^{2}(2r)[1+\cos(\theta_{e}-\theta_{d})],$
-$M_{s}=$ $\dfrac{1}{2}e^{i\theta_{d}}\sinh(2r)[1+e^{i(\theta_{e}-\theta_{d})}]$
-×$\times[\cosh^{2}(r)+e^{-i(\theta_{e}-\theta_{d})}\sinh^{2}(r)].$ (31)
-Obviously, the thermal noise $N_{s}$ and the two-photon-correlation strength $M_{s}$ can be completely eliminated by choosing $r_{p}=r_{e}$ and $\theta_{e}-\theta_{d}=\pm n\pi(n=1,3,5...)$, i.e., $N_{s}=M_{s}=0$.
-As discussed in Ref.
-, the reservoir of the original optical mode is squeezed along the axis with angle $\dfrac{\theta_{e}}{2}$.
-In the basis of the squeezed optical mode $\hat{a}_{s,\circlearrowleft}$, this effect can be canceled by the squeezing (along axis $\dfrac{\theta_{d}}{2}$) induced by the parametric amplification of $\hat{a}_{\circlearrowleft}$, when $r_{p}=r_{e}$ and $\theta_{e}+\theta_{d}=\pm n\pi(n=1,3,5...)$, yielding the phase-matched condition.
+$N_{s,\circlearrowleft}=$ $\sinh^{2}(r_{d})\cosh^{2}(r_{e,\circlearrowleft})+\cosh^{2}(r_{d})\sinh^{2}(r_{e,\circlearrowleft})$
+$+\dfrac{1}{2}\cos\delta\theta\sinh(2r_{d})\sinh(2r_{e,\circlearrowleft}),$
+$M_{s,\circlearrowleft}\!=$ $e^{i\theta_{d}}\![\sinh(r_{d})\!\cosh(r_{e,\circlearrowleft})\!+\!e^{i\delta\theta_{\circlearrowleft}}\!\cosh(r_{d})\!\sinh(r_{e,\circlearrowleft})\!]$
+×$\times\![\cosh(r_{d})\!\cosh(r_{e,\circlearrowleft})\!+\!e^{-i\delta\theta_{\circlearrowleft}}\!\sinh(r_{d})\!\sinh(r_{e,\circlearrowleft})].$ (31)
+Here $N_{s,\circlearrowleft}$ and $M_{s,\circlearrowleft}$ denote the effective thermal noise and two-photon-correlation strength, respectively, which can be simplified in case of $r_{d}=r_{e,\circlearrowleft}=r$, i.e.,
+$N_{s,\circlearrowleft}=$ $\dfrac{1}{2}\sinh^{2}(2r)[1+\cos(\theta_{e,\circlearrowleft}-\theta_{d})],$
+$M_{s,\circlearrowleft}=$ $\dfrac{1}{2}e^{i\theta_{d}}\sinh(2r)[1+e^{i(\theta_{e,\circlearrowleft}-\theta_{d})}]$
+×$\times[\cosh^{2}(r)+e^{-i(\theta_{e,\circlearrowleft}-\theta_{d})}\sinh^{2}(r)].$ (32)
+Obviously, the thermal noise $N_{s,\circlearrowleft}$ and the two-photon-correlation strength $M_{s,\circlearrowleft}$ can be completely eliminated by choosing $r_{d}=r_{e,\circlearrowleft}$ and $\theta_{e,\circlearrowleft}-\theta_{d}=\pm n\pi(n=1,3,5...)$, i.e., $N_{s,\circlearrowleft}=M_{s,\circlearrowleft}=0$.
+The reservoir of the original optical mode is squeezed along the axis with angle $\dfrac{\theta_{e,\circlearrowleft}}{2}$.
+In the basis of the squeezed optical mode $\hat{a}_{s,\circlearrowleft}$, this effect can be canceled by the squeezing (along axis $\dfrac{\theta_{d}}{2}$) induced by the parametric amplification of $\hat{a}_{\circlearrowleft}$, when $r_{d}=r_{e,\circlearrowleft}$ and $\theta_{e,\circlearrowleft}+\theta_{d}=\pm n\pi(n=1,3,5...)$, yielding the phase-matched condition.
 This means the squeezed-vacuum reservoir of $\hat{a}_{\circlearrowleft}$ corresponds to an effective vacuum reservoir of $\hat{a}_{s,\circlearrowleft}$ under the phase-matched condition.
@@
 By employing the linearized effective master equation (11), the dynamics of the quantum correlation function in the vector $X$ can be derived as
-$\dot{x}_{1}=-\kappa x_{1}+i\Lambda_{1}x_{14}-i\Lambda_{1}^{\ast}x_{13}+i\Lambda_{2}x_{18}-i\Lambda_{2}^{\ast}x_{17},$ (34)
-$\dot{x}_{2}=$ $-\kappa x_{2}+i\Lambda_{1}x_{14}-i\Lambda_{1}^{\ast}x_{13}+i\Lambda_{2}x_{18}-i\Lambda_{2}^{\ast}x_{17}+\kappa,$ (35)
-$\dot{x}_{3}=2\omega_{m,1}x_{9}+2\omega_{m,1}x_{10},$ (36)
-$\dot{x}_{4}=$ $-2\gamma_{m,1}x_{4}-2\omega_{m,1}x_{9}-2\omega_{m,1}x_{10}+2\Lambda_{1}x_{16}$
-$+2\Lambda_{1}^{\ast}x_{15}-2\chi x_{24}+2\gamma_{m,1}\bar{n}_{m,1},$ (37)
-$\dot{x}_{5}=2\omega_{m,2}x_{11}+2\omega_{m,2}x_{12},$ (38)
-$\dot{x}_{6}=$ $-2\gamma_{m,2}x_{6}-2\omega_{m,2}x_{11}-2\omega_{m,2}x_{12}+2\Lambda_{2}x_{20}$
-$+2\Lambda_{2}^{\ast}x_{19}-2\chi x_{23}+2\gamma_{m,2}\bar{n}_{m,2},$ (39)
-$\dot{x}_{7}=-(2i\Delta_{s}+\kappa)x_{7}+2i\Lambda_{1}x_{13}+2i\Lambda_{2}x_{17},$ (40)
-$\dot{x}_{8}=(2i\Delta_{s}-\kappa)x_{8}-2i\Lambda_{1}^{\ast}x_{14}-2i\Lambda_{2}^{\ast}x_{18},$ (41)
+$\dot{x}_{1}=$ $-\kappa x_{1}+i\Lambda_{1}x_{14}-i\Lambda_{1}^{\ast}x_{13}+i\Lambda_{2}x_{18}$
+$-i\Lambda_{2}^{\ast}x_{17}+\kappa N_{s},$ (35)
+$\dot{x}_{2}=$ $-\kappa x_{2}+i\Lambda_{1}x_{14}-i\Lambda_{1}^{\ast}x_{13}+i\Lambda_{2}x_{18}$
+$-i\Lambda_{2}^{\ast}x_{17}+\kappa(N_{s}+1),$ (36)
+$\dot{x}_{3}=\omega_{m,1}x_{9}+\omega_{m,1}x_{10},$ (37)
+$\dot{x}_{4}=$ $-2\gamma_{m,1}x_{4}-\omega_{m,1}x_{9}-\omega_{m,1}x_{10}+2\Lambda_{1}x_{16}$
+$+2\Lambda_{1}^{\ast}x_{15}-2\lambda x_{24}+2\gamma_{m,1}\bar{n}_{m,1},$ (38)
+$\dot{x}_{5}=\omega_{m,2}x_{11}+\omega_{m,2}x_{12},$ (39)
+$\dot{x}_{6}=$ $-2\gamma_{m,2}x_{6}-\omega_{m,2}x_{11}-\omega_{m,2}x_{12}+2\Lambda_{2}x_{20}$
+$+2\Lambda_{2}^{\ast}x_{19}-2\lambda x_{23}+2\gamma_{m,2}\bar{n}_{m,2},$ (40)
+$\dot{x}_{7}=-(2i\Delta_{s}+\kappa)x_{7}+2i\Lambda_{1}x_{13}+2i\Lambda_{2}x_{17}+\kappa M_{s}^{\ast},$ (41)
+$\dot{x}_{8}=(2i\Delta_{s}-\kappa)x_{8}-2i\Lambda_{1}^{\ast}x_{14}-2i\Lambda_{2}^{\ast}x_{18}+\kappa M_{s},$ (42)
 $\dot{x}_{9}=$ $-\dfrac{\gamma_{m,1}}{2}x_{9}-\dfrac{\gamma_{m,1}}{2}x_{10}+\omega_{m,1}x_{4}-\omega_{m,1}x_{3}$
-$-\chi x_{21}+\Lambda_{1}x_{14}+\Lambda_{1}^{\ast}x_{13},$ (42)
+$-\lambda x_{21}+\Lambda_{1}x_{14}+\Lambda_{1}^{\ast}x_{13},$ (43)
 $\dot{x}_{10}=$ $-\dfrac{\gamma_{m,1}}{2}x_{10}-\dfrac{\gamma_{m,1}}{2}x_{9}+\omega_{m,1}x_{4}-\omega_{m,1}x_{3}$
-$-\chi x_{21}+\Lambda_{1}x_{14}+\Lambda_{1}^{\ast}x_{13},$ (43)
+$-\lambda x_{21}+\Lambda_{1}x_{14}+\Lambda_{1}^{\ast}x_{13},$ (44)
 $\dot{x}_{11}=$ $-\dfrac{\gamma_{m,2}}{2}x_{11}-\dfrac{\gamma_{m,2}}{2}x_{12}+\omega_{m,2}x_{6}-\omega_{m,2}x_{5}$
-$-\chi x_{21}+\Lambda_{2}x_{18}+\Lambda_{2}^{\ast}x_{17},$ (44)
+$-\lambda x_{21}+\Lambda_{2}x_{18}+\Lambda_{2}^{\ast}x_{17},$ (45)
 $\dot{x}_{12}=$ $-\dfrac{\gamma_{m,2}}{2}x_{12}-\dfrac{\gamma_{m,2}}{2}x_{11}+\omega_{m,2}x_{6}-\omega_{m,2}x_{5}$
-$-\chi x_{21}+\Lambda_{2}x_{18}+\Lambda_{2}^{\ast}x_{17},$ (45)
-$\dot{x}_{13}=-\left(i\Delta_{s}+\dfrac{\kappa}{2}\right)x_{13}+\omega_{m,1}x_{15}+i\Lambda_{1}x_{3}+i\Lambda_{2}x_{21},$ (46)
-$\dot{x}_{14}=\left(i\Delta_{s}-\dfrac{\kappa}{2}\right)x_{14}+\omega_{m,1}x_{16}-i\Lambda_{1}^{\ast}x_{3}-i\Lambda_{2}^{\ast}x_{21},$ (47)
-$\dot{x}_{15}=$ $-\left(i\Delta_{s}+\gamma_{m,1}+\dfrac{\kappa}{2}\right)x_{15}-\omega_{m,1}x_{13}+i\Lambda_{1}x_{10}$
-$+i\Lambda_{2}x_{24}-\chi x_{17},$ (48)
-$\dot{x}_{16}=$ $\left(i\Delta_{s}-\gamma_{m,1}-\dfrac{\kappa}{2}\right)x_{16}-\omega_{m,1}x_{14}-i\Lambda_{1}^{\ast}x_{10}$
-$-i\Lambda_{2}^{\ast}x_{24}-\chi x_{18},$ (49)
-$\dot{x}_{17}=-\left(i\Delta_{s}+\dfrac{\kappa}{2}\right)x_{17}+\omega_{m,1}x_{19}+i\Lambda_{1}x_{21}+i\Lambda_{2}x_{5},$ (50)
-$\dot{x}_{18}=\left(i\Delta_{s}-\dfrac{\kappa}{2}\right)x_{18}+\omega_{m,1}x_{20}-i\Lambda_{1}^{\ast}x_{21}-i\Lambda_{2}^{\ast}x_{5},$ (51)
+$-\lambda x_{21}+\Lambda_{2}x_{18}+\Lambda_{2}^{\ast}x_{17},$ (46)
+$\dot{x}_{13}=$ $-\left(i\Delta_{s}+\dfrac{\kappa}{2}\right)x_{13}+\omega_{m,1}x_{15}$
+$+i\Lambda_{1}x_{3}+i\Lambda_{2}x_{21},$ (47)
+$\dot{x}_{14}=$ $\left(i\Delta_{s}-\dfrac{\kappa}{2}\right)x_{14}+\omega_{m,1}x_{16}$
+$-i\Lambda_{1}^{\ast}x_{3}-i\Lambda_{2}^{\ast}x_{21},$ (48)
+$\dot{x}_{15}=$ $-\left(i\Delta_{s}+\gamma_{m,1}+\dfrac{\kappa}{2}\right)x_{15}-\omega_{m,1}x_{13}+\Lambda_{1}^{\ast}x_{7}$
+$+\Lambda_{1}x_{2}+i\Lambda_{1}x_{9}+i\Lambda_{2}x_{24}-\lambda x_{17},$ (49)
+$\dot{x}_{16}=$ $\left(i\Delta_{s}-\gamma_{m,1}-\dfrac{\kappa}{2}\right)x_{16}-\omega_{m,1}x_{14}+\Lambda_{1}x_{8}$
+$+\Lambda_{1}^{\ast}x_{1}-i\Lambda_{1}^{\ast}x_{9}-i\Lambda_{2}^{\ast}x_{24}-\lambda x_{18},$ (50)
+$\dot{x}_{17}=$ $-\left(i\Delta_{s}+\dfrac{\kappa}{2}\right)x_{17}+\omega_{m,1}x_{19}$
+$+i\Lambda_{1}x_{21}+i\Lambda_{2}x_{5},$ (51)
+$\dot{x}_{18}=$ $\left(i\Delta_{s}-\dfrac{\kappa}{2}\right)x_{18}+\omega_{m,1}x_{20}$
+$-i\Lambda_{1}^{\ast}x_{21}-i\Lambda_{2}^{\ast}x_{5},$ (52)
 $\dot{x}_{19}=$ $-\left(i\Delta_{s}+\gamma_{m,2}+\dfrac{\kappa}{2}\right)x_{19}-\omega_{m,2}x_{17}+i\Lambda_{1}x_{23}$
-$+i\Lambda_{2}x_{12}-\chi x_{13},$ (52)
+$+\Lambda_{2}x_{2}+\Lambda_{2}^{\ast}x_{7}+i\Lambda_{2}x_{11}-\lambda x_{13},$ (53)
 $\dot{x}_{20}=$ $\left(i\Delta_{s}-\gamma_{m,2}-\dfrac{\kappa}{2}\right)x_{20}-\omega_{m,2}x_{18}-i\Lambda_{1}^{\ast}x_{23}$
-$-i\Lambda_{2}^{\ast}x_{12}-\chi x_{14},$ (53)
+$+\Lambda_{2}^{\ast}x_{1}+\Lambda_{2}x_{8}-i\Lambda_{2}^{\ast}x_{11}-\lambda x_{14},$ (54)
 $\dot{x}_{21}=\omega_{m,1}x_{24}+\omega_{m,2}x_{23},$ (55)
-$\dot{x}_{22}=$ $-(\gamma_{m,1}+\gamma_{m,2})x_{22}-\omega_{m,1}x_{23}-\omega_{m,2}x_{24}-\chi x_{10}$
-$-\chi x_{11}+\Lambda_{1}x_{20}+\Lambda_{1}^{\ast}x_{19}+\Lambda_{2}x_{16}+\Lambda_{2}^{\ast}x_{15},$ (55)
-$\dot{x}_{23}=$ $-\gamma_{m,2}x_{23}+\omega_{m,1}x_{22}-\omega_{m,2}x_{21}-\chi x_{3}$
-$+\Lambda_{2}x_{14}+\Lambda_{2}^{\ast}x_{13},$ (56)
-$\dot{x}_{24}=$ $-\gamma_{m,1}x_{24}-\omega_{m,1}x_{21}+\omega_{m,2}x_{22}-\chi x_{5}$
-$+\Lambda_{1}x_{18}+\Lambda_{1}^{\ast}x_{17}.$ (57)
+$\dot{x}_{22}=$ $-(\gamma_{m,1}+\gamma_{m,2})x_{22}-\omega_{m,1}x_{23}-\omega_{m,2}x_{24}-\lambda x_{10}$
+$-\lambda x_{11}+\Lambda_{1}x_{20}+\Lambda_{1}^{\ast}x_{19}+\Lambda_{2}x_{16}+\Lambda_{2}^{\ast}x_{15},$ (56)
+$\dot{x}_{23}=$ $-\gamma_{m,2}x_{23}+\omega_{m,1}x_{22}-\omega_{m,2}x_{21}$
+$-\lambda x_{3}+\Lambda_{2}x_{14}+\Lambda_{2}^{\ast}x_{13},$ (57)
+$\dot{x}_{24}=$ $-\gamma_{m,1}x_{24}-\omega_{m,1}x_{21}+\omega_{m,2}x_{22}$
+$-\lambda x_{5}+\Lambda_{1}x_{18}+\Lambda_{1}^{\ast}x_{17}.$ (58)
 By rewriting the above equations in a compact form, i.e.,
-$\dot{X}=M\cdot X+N,$ (58)
+$\dot{X}=A\cdot X+b,$ (59)
 and introducing the following expressions
@@
 $\Omega_{2,\pm}$ $=i\Delta_{s}\pm\gamma_{m,1}\pm\dfrac{\kappa}{2},$
-$\Omega_{3,\pm}$ $=i\Delta_{s}\pm\gamma_{m,2}\pm\dfrac{\kappa}{2}.,$ (59)
+$\Omega_{3,\pm}$ $=i\Delta_{s}\pm\gamma_{m,2}\pm\dfrac{\kappa}{2},$
+$\Omega_{4}$ $=\gamma_{m,1}+\gamma_{m,2}.$ (60)
 we have
-$N=$ $\left[0,\kappa,0,2\gamma_{m,1}\bar{n}_{m,1},0,2\gamma_{m,2}\bar{n}_{m,2},0,0,\right.$
-$\left.0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0\right]^{T},$ (60)
+$A=\left(\begin{matrix}A_{11}&A_{12}&A_{13}&A_{14}\\ A_{21}&A_{22}&A_{23}&A_{24}\\ A_{31}&A_{32}&A_{33}&A_{34}\\ A_{41}&A_{42}&A_{43}&A_{44}\\ \end{matrix}\right),$ (61)
+$b=$ $\left[\kappa N_{s},\kappa(N_{s}+1),0,2\gamma_{m,1}\bar{n}_{m,1},0,2\gamma_{m,2}\bar{n}_{m,2},\kappa M_{s}^{\ast},\right.$
+$\left.\kappa M_{s},0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0\right]^{T},$ (62)
 and
-$M=$
-($\begin{smallmatrix}-\kappa&0&0&0&0&0&0&0&0&0&0&0&-i\Lambda_{1}^{\ast}&i\Lambda_{1}&0&0\\ -i\Lambda_{2}^{\ast}&i\Lambda_{2}&0&0&0&0&0&0\\ \\ 0&-\kappa&0&0&0&0&0&0&0&0&0&0&-i\Lambda_{1}^{\ast}&i\Lambda_{1}&0&0\\ -i\Lambda_{2}^{\ast}&i\Lambda_{2}&0&0&0&0&0&0\\ \\ 0&0&0&0&0&0&0&0&2\omega_{m,1}&2\omega_{m,1}&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0\\ \\ 0&0&0&-2\gamma_{m,1}&0&0&0&0&-2\omega_{m,1}&-2\omega_{m,1}&0&0&0&0&2\Lambda_{1}^{\ast}&2\Lambda_{1}\\ 0&0&0&0&0&0&0&-2\chi\\ \\ 0&0&0&0&0&0&0&0&0&0&2\omega_{m,2}&2\omega_{m,2}&0&0&0&0\\ 0&0&0&0&0&0&0&0\\ \\ 0&0&0&0&0&-2\gamma_{m,2}&0&0&0&0&-2\omega_{m,2}&-2\omega_{m,2}&0&0&0&0\\ 0&0&2\Lambda_{2}^{\ast}&2\Lambda_{2}&0&0&-2\chi&0\\ \\ 0&0&0&0&0&0&-2\Omega_{1,+}&0&0&0&0&0&2i\Lambda_{1}&0&0&0\\ 2i\Lambda_{2}&0&0&0&0&0&0&0\\ \\ 0&0&0&0&0&0&0&2\Omega_{1,-}&0&0&0&0&0&-2i\Lambda_{1}^{\ast}&0&0\\ 0&-2i\Lambda_{2}^{\ast}&0&0&0&0&0&0\\ \\ 0&0&-\omega_{m,1}&\omega_{m,1}&0&0&0&0&-\gamma_{m,1}/2&-\gamma_{m,1}/2&0&0&\Lambda_{1}^{\ast}&\Lambda_{1}&0&0\\ 0&0&0&0&-\chi&0&0&0\\ \\ 0&0&-\omega_{m,1}&\omega_{m,1}&0&0&0&0&-\gamma_{m,1}/2&-\gamma_{m,1}/2&0&0&\Lambda_{1}^{\ast}&\Lambda_{1}&0&0\\ 0&0&0&0&-\chi&0&0&0\\ \\ 0&0&0&0&-\omega_{m,2}&\omega_{m,2}&0&0&0&0&-\gamma_{m,2}/2&-\gamma_{m,2}/2&0&0&0&0\\ \Lambda_{2}^{\ast}&\Lambda_{2}&0&0&-\chi&0&0&0\\ \\ 0&0&0&0&-\omega_{m,2}&\omega_{m,2}&0&0&0&0&-\gamma_{m,2}/2&-\gamma_{m,2}/2&0&0&0&0\\ \Lambda_{2}^{\ast}&\Lambda_{2}&0&0&-\chi&0&0&0\\ \\ 0&0&i\Lambda_{1}&0&0&0&0&0&0&0&0&0&-\Omega_{1,+}&0&\omega_{m,1}&0\\ 0&0&0&0&i\Lambda_{2}&0&0&0\\ \\ 0&0&-i\Lambda_{1}^{\ast}&0&0&0&0&0&0&0&0&0&0&\Omega_{1,-}&0&\omega_{m,1}\\ 0&0&0&0&-i\Lambda_{2}^{\ast}&0&0&0\\ \\ 0&0&0&0&0&0&0&0&0&i\Lambda_{1}&0&0&-\omega_{m,1}&0&-\Omega_{2,+}&0\\ -\chi&0&0&0&0&0&0&i\Lambda_{2}\\ \\ 0&0&0&0&0&0&0&0&0&-i\Lambda_{1}^{\ast}&0&0&0&-\omega_{m,1}&0&\Omega_{2,-}\\ 0&-\chi&0&0&0&0&0&-i\Lambda_{2}^{\ast}\\ \\ 0&0&0&0&i\Lambda_{2}&0&0&0&0&0&0&0&0&0&0&0\\ -\Omega_{1,+}&0&\omega_{m,1}&0&i\Lambda_{1}&0&0&0\\ \\ 0&0&0&0&-i\Lambda_{2}^{\ast}&0&0&0&0&0&0&0&0&0&0&0\\ 0&\Omega_{1,-}&0&\omega_{m,1}&-i\Lambda_{1}^{\ast}&0&0&0\\ \\ 0&0&0&0&0&0&0&0&0&0&0&i\Lambda_{2}&-\chi&0&0&0\\ -\omega_{m,2}&0&-\Omega_{3,+}&0&0&0&i\Lambda_{1}&0\\ \\ 0&0&0&0&0&0&0&0&0&0&0&-i\Lambda_{2}^{\ast}&0&-\chi&0&0\\ 0&-\omega_{m,2}&0&\Omega_{3,-}&0&0&-i\Lambda_{1}^{\ast}&0\\ \\ 0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&\omega_{m,2}&\omega_{m,1}\\ 0&0&0&0&0&0&0&0&0&-\chi&-\chi&0&0&0&\Lambda_{2}^{\ast}&\Lambda_{2}\\ 0&0&\Lambda_{1}^{\ast}&\Lambda_{1}&0&-\gamma_{m,1}-\gamma_{m,2}&-\omega_{m,1}&-\omega_{m,2}\\ 0&0&-\chi&0&0&0&0&0&0&0&0&0&\Lambda_{2}^{\ast}&\Lambda_{2}&0&0\\ 0&0&0&0&-\omega_{m,2}&\omega_{m,1}&-\gamma_{m,2}&0\\ 0&0&0&0&-\chi&0&0&0&0&0&0&0&0&0&0&0\\ \Lambda_{1}^{\ast}&\Lambda_{1}&0&0&-\omega_{m,1}&\omega_{m,2}&0&-\gamma_{m,1}\\ \end{smallmatrix}\right).$
+$A_{11}=\left(\begin{matrix}-\kappa&0&0&0&0&0\\ 0&-\kappa&0&0&0&0\\ 0&0&0&0&0&0\\ 0&0&0&-2\gamma_{m,1}&0&0\\ 0&0&0&0&0&0\\ 0&0&0&0&0&-2\gamma_{m,2}\end{matrix}\right),$
+$A_{12}=\left(\begin{matrix}0&0&0&0&0&0\\ 0&0&0&0&0&0\\ 0&0&\omega_{m,1}&\omega_{m,1}&0&0\\ 0&0&-\omega_{m,1}&-\omega_{m,1}&0&0\\ 0&0&0&0&\omega_{m,2}&\omega_{m,2}\\ 0&0&0&0&-\omega_{m,2}&-\omega_{m,2}\end{matrix}\right),$
+$A_{13}=\left(\begin{matrix}-i\Lambda_{1}^{\ast}&i\Lambda_{1}&0&0&-i\Lambda_{2}^{\ast}&i\Lambda_{2}\\ -i\Lambda_{1}^{\ast}&i\Lambda_{1}&0&0&-i\Lambda_{2}^{\ast}&i\Lambda_{2}\\ 0&0&0&0&0&0\\ 0&0&2\Lambda_{1}^{\ast}&2\Lambda_{1}&0&0\\ 0&0&0&0&0&0\\ 0&0&0&0&0&0\end{matrix}\right),$
+$A_{14}=\left(\begin{matrix}0&0&0&0&0&0\\ 0&0&0&0&0&0\\ 0&0&0&0&0&0\\ 0&0&0&0&0&-2\lambda\\ 0&0&0&0&0&0\\ 2\Lambda_{2}^{\ast}&2\Lambda_{2}&0&0&-2\lambda&0\end{matrix}\right),$
+$A_{21}=\left(\begin{matrix}0&0&0&0&0&0\\ 0&0&0&0&0&0\\ 0&0&-\omega_{m,1}&\omega_{m,1}&0&0\\ 0&0&-\omega_{m,1}&\omega_{m,1}&0&0\\ 0&0&0&0&-\omega_{m,2}&\omega_{m,2}\\ 0&0&0&0&-\omega_{m,2}&\omega_{m,2}\end{matrix}\right),$
+$A_{22}=\dfrac{1}{2}\left(\begin{matrix}-4\Omega_{1,+}&0&0&0&0&0\\ 0&4\Omega_{1,-}&0&0&0&0\\ 0&0&-\gamma_{m,1}&-\gamma_{m,1}&0&0\\ 0&0&-\gamma_{m,1}&-\gamma_{m,1}&0&0\\ 0&0&0&0&-\gamma_{m,2}&-\gamma_{m,2}\\ 0&0&0&0&-\gamma_{m,2}&-\gamma_{m,2}\end{matrix}\right),$
+$A_{23}=\left(\begin{matrix}2i\Lambda_{1}&0&0&0&2i\Lambda_{2}&0\\ 0&-2i\Lambda_{1}^{\ast}&0&0&0&-2i\Lambda_{2}^{\ast}\\ \Lambda_{1}^{\ast}&\Lambda_{1}&0&0&0&0\\ \Lambda_{1}^{\ast}&\Lambda_{1}&0&0&0&0\\ 0&0&0&0&\Lambda_{2}^{\ast}&\Lambda_{2}\\ 0&0&0&0&\Lambda_{2}^{\ast}&\Lambda_{2}\end{matrix}\right),$
+$A_{24}=\left(\begin{matrix}0&0&0&0&0&0\\ 0&0&0&0&0&0\\ 0&0&-\lambda&0&0&0\\ 0&0&-\lambda&0&0&0\\ 0&0&-\lambda&0&0&0\\ 0&0&-\lambda&0&0&0\end{matrix}\right),$
+$A_{31}=\left(\begin{matrix}0&0&i\Lambda_{1}&0&0&0\\ 0&0&-i\Lambda_{1}^{\ast}&0&0&0\\ 0&\Lambda_{1}&0&0&0&0\\ \Lambda_{1}^{\ast}&0&0&0&0&0\\ 0&0&0&0&i\Lambda_{2}&0\\ 0&0&0&0&-i\Lambda_{2}^{\ast}&0\end{matrix}\right),$
+$A_{32}=\left(\begin{matrix}0&0&0&0&0&0\\ 0&0&0&0&0&0\\ \Lambda_{1}^{\ast}&0&i\Lambda_{1}&0&0&0\\ 0&\Lambda_{1}&-i\Lambda_{1}^{\ast}&0&0&0\\ 0&0&0&0&0&0\\ 0&0&0&0&0&0\end{matrix}\right),$
+$A_{33}=\left(\begin{matrix}-\Omega_{1,+}&0&\omega_{m,1}&0&0&0\\ 0&\Omega_{1,-}&0&\omega_{m,1}&0&0\\ -\omega_{m,1}&0&-\Omega_{2,+}&0&-\lambda&0\\ 0&-\omega_{m,1}&0&\Omega_{2,-}&0&-\lambda\\ 0&0&0&0&-\Omega_{1,+}&0\\ 0&0&0&0&0&\Omega_{1,-}\end{matrix}\right),$
+$A_{34}=\left(\begin{matrix}0&0&i\Lambda_{2}&0&0&0\\ 0&0&-i\Lambda_{2}^{\ast}&0&0&0\\ 0&0&0&0&0&i\Lambda_{2}\\ 0&0&0&0&0&-i\Lambda_{2}^{\ast}\\ \omega_{m,1}&0&i\Lambda_{1}&0&0&0\\ 0&\omega_{m,1}&-i\Lambda_{1}^{\ast}&0&0&0\end{matrix}\right),$
+$A_{41}=\left(\begin{matrix}0&\Lambda_{2}&0&0&0&0\\ \Lambda_{2}^{\ast}&0&0&0&0&0\\ 0&0&0&0&0&0\\ 0&0&0&0&0&0\\ 0&0&-\lambda&0&0&0\\ 0&0&0&0&-\lambda&0\end{matrix}\right),$
+$A_{42}=\left(\begin{matrix}\Lambda_{2}^{\ast}&0&0&0&i\Lambda_{2}&0\\ 0&\Lambda_{2}&0&0&-i\Lambda_{2}^{\ast}&0\\ 0&0&0&0&0&0\\ 0&0&0&-\lambda&-\lambda&0\\ 0&0&0&0&0&0\\ 0&0&0&0&0&0\end{matrix}\right),$
+$A_{43}=\left(\begin{matrix}-\lambda&0&0&0&-\omega_{m,2}&0\\ 0&-\lambda&0&0&0&-\omega_{m,2}\\ 0&0&0&0&0&0\\ 0&0&\Lambda_{2}^{\ast}&\Lambda_{2}&0&0\\ \Lambda_{2}^{\ast}&\Lambda_{2}&0&0&0&0\\ 0&0&0&0&\Lambda_{1}^{\ast}&\Lambda_{1}\end{matrix}\right),$
+$A_{44}=\left(\begin{matrix}-\Omega_{3,+}&0&0&0&i\Lambda_{1}&0\\ 0&\Omega_{3,-}&0&0&-i\Lambda_{1}^{\ast}&0\\ 0&0&0&0&\omega_{m,2}&\omega_{m,1}\\ \Lambda_{1}^{\ast}&\Lambda_{1}&0&-\Omega_{4}&-\omega_{m,1}&-\omega_{m,2}\\ 0&0&-\omega_{m,2}&\omega_{m,1}&-\gamma_{m,2}&0\\ 0&0&-\omega_{m,1}&\omega_{m,2}&0&-\gamma_{m,1}\end{matrix}\right).$